A Bibliography of Publications about the SAS Statistical Analysis System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 2.45

Title word cross-reference

$109.00 [Sha20]. $118.10 [Sha22]. $139.95 [Sta21]. $144.49 [Sha22]. 2^{k} [Law08]. 2^{k−p} [Law08]. $31.95 [BR09]. $39.95 [McF01]. $59.95 [Ano99]. $69.95 [Bie08, Dob09, Jon07]. $79.95 [Oli20, Rob08]. $89.95 [Bos07, Kel10, Lum08]. $93.22 [Sha22]. $99.95 [Rah24, Woo11]. ^{+} [BN93a, BN93b, BN95a, BN95b]. ^{R} [MJ97]. (R) [Ros15, SB96, Len15]. κ^{2} [MJ97].

//www.wiley.com/WileyCDA/WileyTitle/productCd [Kha15]. //www.wiley.com/WileyCDA/WileyTitle/productCd-1118762614.html [Kha15].

0 [Chi02, Nie08, Rah24]. 0-412-71050-1 [Len97]. 0-8039-4984-7 [Til96]. 0-8039-4985-5 [Til96].

1 [Ano06b, Chi02, Jon07, Lum08, McF01, Sch06b]. 1-2-3 [Fri91b]. 1-4200-6060-0 [Woo11]. 1-58025-357-1 [McF01]. 1-58025-397-0 [Chi02]. 1-58488-469-X [Ano06b]. 1-58488-619-6 [Lum08]. 1-58488-962-4 [Kel10]. 1-59047-080-X [Sch06b]. 1-59047-500-3 [Bos07]. 1-59047-504-6 [Jon07]. 10th [Cou03]. 1118762614.html [Kha15]. 11th [Gia04]. 125.00 [Laz24]. 12th [HS05]. 13th [Yi06]. 14th [ACM96, NF07]. 15th [AV08]. 16th [PS09, SUG91]. 17th [CM11, Hut92]. 18th [Yah11]. 19.99 [Len97]. 1985 [All86]. 1991/SUGI [SUG91]. 1993 [Ogl93]. 1993-94 [Bir94]. 1994 [Bir94, DH94].

1994-95 [Bir94]. 19th [MS12]. 1st [RHWP24, Sau11].

2 [Har93, SAS90o, SAS93v, SAS93w, SAS93k, SAS94i, SAS94-31, SAS94-34, SAS95h, SAS95-29]. 2.0 [KCM08]. 2000 [Sau04b]. 2002 [Sau04a, War05]. 20th [Ano95a]. 22nd [Ano97]. 23.99/$29.99 [Kha20]. 23rd [Ano98]. 294pp [Laz24]. 2nd [Bos07, Gup02a, Gup02b, Hil93, Lou06, Oli20, Rut12].

3 [Bos07, Dob09, Sha22, Zha19]. 3-437-40281-1 [Ano93a]. 3-7785-2157-8 [Ano93b]. 3.1 [Gil96]. 338pp [McF01]. 3D [CS89]. 3rd [Ano08a, Ano09a, Ano09b, Lum08].

49. [Ano93a]. 49.- [Ano93a]. 4th [Cow04a, Cow04b, Hor09b, Sta21, Van97].

5 [Kel10, Kha20]. 5th [CEM82, Lev98, SLMN86].

6 [Göt92, Jon07, Lum08, Sha20, Ano93a]. 6.10-6.11 [LJ95]. 68 [Ano93b]. 6th [CF99].

7 [Sha22, Woo11]. 7-9 [SUG87]. 7th [Pal00].

8 [BR09, Bie08]. ’87 [SUG87]. ’88 [Tie88]. ’89 [Tie89]. 8th [Cou01].

9 [Oli20, Sha22, Sta21]. 9.1.3 [Ano06e]. 90 [SB01]. ’94 [DH94, Le 94]. ’95 [Myc95, LS05]. ’96 [ACM96, CS96]. 978 [BR09, Bie08, Bos07, Dob09, Kel10, Kha20, Nie08, Oli20, Rah24, Rob08, Sha20, Sha22, Sta21, Woo11, Zha19]. 978-0-19-929659-0 [Nie08]. 978-0-19-929660-6 [Nie08]. 978-0-367-35797-9 [Oli20]. 978-0-367-36547-9 [Laz24, Sha22]. 978-0-367-55054-7 [Sha22]. 978-0-367-75538-6 [RHWP24]. 978-1-032-32562-0 [Rah24].

978-1-119-95021-9 [Kha15]. 978-1-138-59152-3 [Sha22]. 978-1-1381-9745-9 [Sta21]. 978-1-4200-6060-7 [Woo11]. 978-1-4822-4936-1 [Oli17]. 978-1-58488-469-X [Rob08]. 978-1-58488-962-5 [Kel10]. 978-1-59047-500-3 [Bos07]. 978-1-59047-886-8 [Bie08]. 978-1-59994-165-3 [Dob09]. 978-1-59994-298-8 [BR09]. 978-1-63526-152-3 [Zha19]. 978-1-64295-806-5 [Kha20]. 978-3-319-388865-6 [Sha20]. 9th [HP02].

= [Mar83].

A. [Ano08a, Ano09c]. A/B [CZZ13]. Aachen [CS96]. Abingdon [RHWP24]. abridged [SAS94o]. Academic [ACM96]. ACCESS [Ano04b, Ano06e, SAS90m, SAS90n, SAS90o, SAS90p, SAS91k, SAS91l, SAS91m, SAS92h, SAS93q, SAS93r, SAS93t, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS94b, SAS94q, SAS94r, SAS94s, SAS94t, SAS95l, SAS95m, SAS95n, SAS96c, Ser93d]. accompany [GI94]. accounting [SAS88r]. ACO [SG80, SH80c, Sea80]. ACT [SUG87]. ADABAS [SAS90m]. Adaptive [Cha08, dSM15, Bra15, Hu09, Kel10, O’B08, Rob09]. addendum [Mil93]. Adding [BS95]. additional [SAS88w]. Additive [BKL05]. ADMB [Liu12]. administrator [SAS88v, SAS91x, SAS93-32, SAS93-44, SAS93-33, SAS95-31]. Advanced [Sch06a, Zel02, Sch06b]. ADX [SAS94-28]. AF [SAS85g, SAS88b, SAS89c, SAS91n, SAS93z, SAS93-27, SAS95o, SAS95p, Spi96]. Agglomerative [Rec10a]. Aided [SH09]. Ákos [All06, Sau04a, Ano03b]. al [Mar83]. Alan [Liu04, Zie98b]. Alex [Bie08, Che08, Chi06, Dur06, Jon07, Sta07c]. Alexander [Puk07]. Algorithm [FTP02, FS13, JMD08]. Algorithms [Che92, Jur15]. Ali [O’B13]. Allen [Cle03a]. Allison [Ano11b, Ano06a, Neu06, Zie97g, Zie00b]. Allocation [FTP02]. allowing [Har81]. Almost [OGR89]. Alonso [Oli17]. Alternative [PG05, EWC94]. Analisis [Mar83]. Analyse [Laf81]. Analyses [Ano09a, Ano09b, DE08, Dha10, Gup02a, Gup02b, Pin10, Wil09, Zie02, DE02, ED96, FMT02, Len97, SH80a, SH80b]. Analysing [Kar08]. Analysis [All95, AV08, AR91, Ami87, Ano06d, Ano06b, Ano09d, Ano11b, Cle03a, Cle03b, CM11, Dam09, Den12, DE05, Die20, Dra11, FZZZ09, FCY20, Gie11, Grö10, Har08, Hec96, Hec11, HH15, HP02, Ke11a, Kha18, KS10, KH10, KGHS84, KW85, Law08, Law10, Laz24, Len09, Lew07, Liu04, Liu17, LM86, LV16, Lu19, Mar83, Mil93, MS12, Mwi11, NZ88, NO06, NF07, OH13, PS09, Pel95, Pri05, Ros15, Sch06c, SC07, Sho18, Spa06, Ste11, Ste06, Utl05, Ven01, Wan19, Wie06, Yah11, Yi06, Zie89, Zie96b, Zie97g, Zie97c, Zie97b, Zie97f, Zie98b, Zie99b, Zie01a, Zim07, ZK12, Agr13, Ano94a, Bir94, Che88, CF99, CS96, Cou01, Cou03, DH96, Dri95, Dru79, FLW11, GG82, Gia04, Gil12, GI94, Gri93, HS05, Hat94, HH04]. analysis [Ing81, JXW^{+}11, KHSG83, KK96, LaP81, LJ95, Law03, Le 94, Lev98, LV17, MK08, Mer83, MPV12, Myc95, NTGD18, Pal00, Rob08, SAS93c, SAS94-35, SAS95-32, SL87, SWH^{+}11, Sha22, Sta86, SDK95, THT12, Tur12, Van97, WWM^{+}24, Wei82, Whi94, Wit92, Woo11, Hei90, Zie97f, Zie99c, Nie11, Sha20, Ano08a, Chi06, Dur06, Jon07, Lum08, Oli10, Sta21]. Analytics [Nas08, Rah24, Sta07a, Wie08]. Analyzing [BKL05, Hal08, KAK05, Len10, BR09, Gho10]. and/or [KM16]. Angeles [PS09]. Ann [Ano06c]. Annotated [SG80, SH80c, SH80a, SH80b, Sea80, Zie89, Zie89]. Annotations [Bar83]. Annual [ACM96, Ano95a, Ano96a, Ano96b, Ano97, Ano98, Gra96, Hut92, Ogl93, Ona90, SUG82, SUG85, SUG89, SUG91, DH94, SUG83, SUG84, SAS88y]. ANOVA [AC04, Gun03, FL81, HP92, MF02, MF03, PKC^{+}20, Zie89]. ANOVA robust [PKC^{+}20]. Anthony [Puk07]. AOS [SAS90a]. AOS/VS [SAS90a]. APB [GL76]. APL [Kar87a, Kar87b]. AppDev [Ano08b, Pra08, BI07]. Application [Hud89, CL90, GL76, GL79, SAS91e, SAS95o]. Applications [Ahm08, LaL92, Ven01, Cou80, DiI91, FMT02, Fer03, How88, KB95, Laf81, Law03, Len15, Mwi12, Rai95, SAS85a, SAS91b, SAS93-35, SAS94j, SAS95d, Sal78a, Sal78b, Sal81a, Spi96, SM82, UM14, Wal90, GI94, Ano11a, Cab04, Ste06, War05, Zha19]. Applied [CS85, CS87, CS91, CS97, FLW11, Fra14, KN95, KN99, Par06, THT12, Ano00a, And86, Ano99, Ano06d, JS13, McF01, O’B13, Oli17, Ric13, Zie96a, Zie98a, Zie00a]. applying [And89]. Approach [AC04, Ano06c, Gun03, Hec96, Lu19, OH13, Oli10, Ros15, Ste99, Zie95a, Zim07, Hat94, HS94, LW95, MF02, MF03, NTGD18, Par06, Lac99]. Approaches [Len09]. April [Ano95a, CM82, Hut92, Ona90, SUG89, Tie88]. Archer [Zie99e]. architecture [EWC94]. Ariel [Oli17]. Arthur [Cle03a]. AS/400 [SAS92h, SAS93q, SAS93r, SAS93s]. ASF [Bir94]. Asia [SUG87]. Asia-Pacific [SUG87]. Aspekte [Ric08]. Assembly [Nor89]. ASSIST [SAS90q, SAS91c, SAS92a, SAS94u, SAS95q, SAS96b, Spi96]. Asymmetry [LS02]. Atkinson [Hor09a, Lu08, Nie08, Puk07]. Attributable [LVK^{+}11]. Audit [Yor94]. auditor [CL90, GL79]. August [Gia04, NF07, PS09, Yi06]. Australia [SUG87]. Auswertungen [KW85]. automatic [RMM18]. autoregressive [QDG17]. available [Gre81]. AXP [SAS94-34].

B [Ano06b, Ano09b, Kar09, LaL09, Len97, Lew07, Liu04, Liu12, Pin10, Sin08, Sta07d, Zie97b, CZZ13]. backgrounds [RHWP24]. Bailer [Oli20, Ng12]. banking [SAS93p]. Barbara [Pal00]. Barry [Jen08]. BASCOM [Ano00b]. Base [Wei12, Fan94, SAS91a, SAS93-32, SAS93-33, Tru88]. Based [DS98b, FS13, LG11, NO06, SS98, Zie99b, ZK12, Ano00b, SHD88]. Basic [ACL04, Ano06c, LF83, LF87, SAS94-36, Van04, Zha19, Mir93, Bas84, Kar09]. Basics [Ann05, All82a, Ray82, SAS85e, Ste83, Ste86]. Bayes [Woo05]. Bayesian [BKL05, WWM^{+}24]. BayesX [BKL05]. Bayo [Ste06]. BDM [MH05]. Beach [SUG84]. beginner [Gil96]. behavior [How88]. Beispielen [Wei94]. Belgium [Le 94]. Bennett [Cle03a]. Berestizhevsky [Zie99b]. Bethel [AC04, Gun03]. Between [WJ19, Fra79, SAS89i]. Beyond [Ann05, Sta94]. Bigirumurame [Oli17]. Bin [Laz24, Sha22]. binary [MPT06]. biologiques [Laf81]. Biometrie [Wei94]. Biostatics [Oli17]. Bivariate [MH05]. Bleicher [Ano93b]. Blind [GL14]. block [Spi96]. Blocking [LSC09]. BMD [Laf81]. BMDP [Bar83, DCT89a, DCT89b, EN88, Fra79, HP92, KW85, Laf81, Pel95, Wit92]. BMDP-77 [Fra79]. Boca [Ano06b, Kel10, Liu17, Lum08, Oli20, RHWP24, Rah24, Rob08, Sha22, Sta21, Woo11]. Bocklebank [Dow88]. Bohnenkamp [Ano08b]. Book [Ahm08, Ale91, All06, And10, AC04, ACL04, AC08, And86, Ano93b, Ano93a, Ano99, Ano03b, Ano03a, Ano04a, Ano06a, Ano06c, Ano06d, Ano06b, Ano08a, Ano08b, Ano09a, Ano09b, Ano09c, Ano09d, Ano11a, Ano11b, Ano12, BR09, Bas03, Bie08, Bos07, Bra15, Bun07, Bzi11, Cab04, Cha04a, Che08, Chi06, Chi02, Cla03, Cle03a, Cow04a, Cow04b, DS98a, Den12, Der10, Die20, Dix07, Dob09, DL04, Dow88, Dra11, Dur06, Gen11, Gho10, Gie11, Grö10, Gun03, Gup02a, Gup02b, Hac84, Har08, Hec90, Hec96, Hec11, Hei90, Hol08, Hor09a, Hor09b, Hu09, JS13, Jen08, Jen09, Jon07, Kar08, Kar09, Kar07, Kas98a, Ke11a, Kel10, Kha15, Kha18, Kha20, Khu13, LaL92, LaL09, Lac99, Laz24, Len09, Len97, Len10, Len15, Lew07]. Book [Liu04, Liu12, Liu17, Lou06, Lu08, Lu19, Lum08, Mai12, Mar88, Mas88, McF01, MT14, Moo03, Mor02, Muk15, Nas08, Neu06, Ng12, Nie08, Nie11, O’B08, O’B13, Oli10, Oli20, Oli17, Oom10, Pin10, Puk07, RHWP24, Rah24, Ree09, Ric13, Ric08, Rob07, Rob08, Rob09, Ros15, Rut12, San03, Sau11, Sch06b, Sch06c, Sea11, Sea12, Sha06, Sha15, Sha20, Sha22, Sho90, Sin08, Spa06, Sta21, Sta07d, Sta07a, Sta07c, Sta07b, Ste11, Ste06, Ste99, Tho92, Til96, Tur12, Utl05, Van04, Ven01, Voe09, Wan19, War05, Web12, Wil09, Woo11, Zha19, Zie89, Zie92a, Zie92b, Zie93, Zie95b, Zie95c, Zie95a, Zie96a, Zie96g, Zie96c, Zie96b, Zie96d, Zie96e, Zie96f, Zie97e, Zie97d, Zie97g, Zie97a, Zie97c, Zie97b, Zie97f, Zie98a, Zie98b]. Book [Zie99b, Zie99d, Zie99c, Zie99a, Zie99e, Zie00c, Zie00b, Zie00a, Zie01b, Zie01a, Zie01c, Zie02, DS95, Mir95]. Books [Sha06, Sha09, Sha15, Sha22]. Bootstrapping [QDG17]. Bootstraps [Opd13]. both [WWM^{+}24]. Brian [Ano06d, Ano09a, Ric13, Rob08, Zie02]. Brocklebank [Ano04a]. Browser [Sun98]. Buch [Ano93b]. BUGS [Lu19]. BugsXLA [Woo05]. Building [ACM96, SAS95a, SMHBR06]. Burch [Zie95c]. Burlew [Bas03, Zie00c]. Burt [Utl05]. Burzykowski [Oli17]. Business [Jen08, Zie97a, Ano84a, Ano94a, CL90, LaP81, Par06, SAS96a, Wan19]. Butte [Tru88]. Buyse [Oli17].

C [All06, Ano03b, Ano04a, Bos07, Bun07, Dow88, Hei90, Hor09a, LaL92, Lou06, Lu08, Mas88, Nie08, Rob07, Sau04a, Sch06c, Zie96b, Zie96e, Zie99c, Zie01c, SAS88c, SAS88d, SAS88e, SAS88f, SAS88g, SAS89d, SAS89e, SAS91p, SAS91o, SAS93b, SAS94v].

C404 [PYC93]. CA [Cou03, Pal00, PS09, Ano97, SAS90n, SAS95l, SAS96c]. CA-DATACOM [SAS90n]. CA-DATACOM/DB [SAS90n]. CA-IDMS [SAS95l, SAS96c]. CALC [SAS91b, SAS91q]. Calculation [LVK^{+}11]. calculations [GL76]. Caldera [Ano00b]. California [SUG82, SUG89]. Canberra [SUG87]. Cantor [Liu04, Zie98b]. capabilities [Fan94, SAS94a]. Carcinogenesis [KGHS84, KHSG83]. Carlo [All06, Ano03b, Sau04a, JXW^{+}11]. Carolina [DH94, Nor89, Sch06b, CHM95]. Carpenter [Zie96e]. carte [CP94]. Cary [Ano99, BR09, Bie08, Bos07, Chi02, Dob09, Jon07, McF01, Sch06b, Hei90]. case [BVFB19, SWH^{+}11]. categorial [THT12]. Categorical [Agr13, Law03, Sta86, SDK95, Sau04b, Ste06, Sch06c, Ste06, Zie96b, Zie01a]. categories [Wei82]. Causal [LM86]. CD [Dob09, War05]. CD-ROM [War05]. CEM [BBV17]. Censored [Lu19]. Center [Ing81]. Central [Ano00b]. Cerrito [Sin08]. certain [GL76]. chair [SUG91]. Chan [Sha20]. Chang [Hu09, Kel10, Sha20, O’B08]. Changes [SAS91a, SAS93b, SAS93a, SAS80, SAS81a, SAS88t, SAS88u, SAS91m, SAS91r, SAS91v, SAS91s, SAS91n, SAS92j, SAS92g, SAS93-28, SAS93-44, SAS93n, SAS93o, SAS94h, SAS94i, SAS94r, SAS94s, SAS94-27, SAS94p, SAS94-33, SAS95g, SAS95h, SAS95q, SAS95t, SAS95w, SAS95k, SAS96d]. Chapman [Kel10, Len97, Lum08, Oli20, Oli17, Rah24, Rob08, War05, Woo11, Zha19]. character [SAS93f]. Characteristic [Gho10, Hal08, Kar08, Len10, BR09]. Charles [Ste06, Zie01a]. Charleston [DH94]. chart [SAS95s]. Chaudhary [Ano08a, Lum08]. CHD [GSD04]. Chen [Rah24]. Cheng [Sha20]. Cheng-Few [Sha20]. chi [Har81]. chi-squared [Har81]. Chicago [Ano96b, Gra96]. Chichester [Kha15]. Christy [Bie08, Che08, Chi06, Dur06, Jon07, Sta07c]. Chronic [KHSG83, KGHS84]. Chuang [Bie08, Bun07, Che08, Chi06, Dur06, Jon07, Sta07c]. Chuang-Stein [Bie08, Bun07, Che08, Chi06, Dur06, Jon07, Sta07c]. CICS [Ano84d, SAS91o]. Cinder [Tru88]. CITIBASE [Sal80]. Claassen [Zha19]. class [SAS95p]. Classical [Len09]. Classification [DS98b, SS98]. click [LW95]. Client [SAS95b, SAS95a]. Client/server [SAS95b, SAS95a]. Clinical [Ano03a, Ano12, Chi06, DL04, Dur06, SAS93-44, Wal02, WWG09, Zie99a, Jon07, TPAM07, WWM^{+}24]. close [JXW^{+}11]. cloth [Til96]. Cluster [Ano00b, Gre17]. Cluster-Randomized [Gre17]. Clustering [Rec10a]. CMS [SAS83c, SAS84b, SAS86, SAS88u, SAS90b, SAS93-45, Saf83, Sal81b, Sal82]. co [Kar87a, Kar87b]. co-operation [Kar87a, Kar87b]. coarsened [BBV17]. Code [Dha10, Gos14, Gre17, Jur15, KM16, LSC09, LG11, LV16, OC15, Pan09, PKC^{+}20, Rec10a, Rec10b, TR14, VDT14, WWG09, Wei12, BVFB19, RMM18, UM14, RMM18]. Coding [GB82, SB96]. Cody [And86, Ano06d, Dob09, Zie96f, Zie98a, Cla03, Hol08, Khu13, San03, Sta07b]. Coefficient [Che92]. Coefficients [TR14]. Cohort [LVK^{+}11]. collaboration [SWH^{+}11]. Cologne [Tie89]. Combinations [JMD08]. combinator [SHD88]. combinator-based [SHD88]. Combining [BMGT15, SAS95c]. comments [Bir94]. Commission [Nor89]. Common [Wal02, Woo05, Ano03a, Ano12, DL04, Zie99a]. CommuniGate [Ano00b]. compact [Göt92]. Companion [Gen11, Kar07, Per10, Sea11, SAS84a, SAS86, SAS90a, SAS90b, SAS90c, SAS90d, SAS92d, SAS93h, SAS93j, SAS93i, SAS93k, SAS93l, SAS93m, SAS95-33, Saf83, Sal81b, Sal82]. comparative [Fra81, GL79]. Comparing [NMvdH14]. Comparison [Gre81, Mon97, EN88, Fan94, HP92, Pel95, Sil96]. Comparisons [Chi02, NZ88, MJ97]. Compensating [CBDM07]. competing [BVFB19]. compilation [Uni94a, Uni94b]. Compiler [SAS88c, SAS88d, SAS89d, SAS91p, SAS93b, SAS94v]. complex [Tru88, Kha18]. Complexity [BN93a, BN93b, BN95a, BN95b]. components [Hoh85, SG80, SH80c, Sea80]. comprehensive [Her90]. COMPSTAT [CEM82]. Computation [And10, Der10, SB01, Web12, Wil89]. Computational [Din06, Oli10, PG05]. Compute [Jur15, Har81]. Computer [All86, Ano00b, Bar83, EM09, How88, Ing81, Zie89, Ell95, Ell00, EM10, EWC94, Hil93, LaP81, Mer83, Mer84, SG80, SH80c, SH80a, SH80b, Sea80, Tru88, Wal90, WHG88, All86]. Computers [SAS88n, AK85, Ano88, CA85, Eas87, LS87, SAS85j, SAS85b, SAS85c, SAS88i]. Computing [JMS^{+}09, Pan09, SAS95b, Sha09]. Conception [Hud89]. Concepts [BVFB19, DS98b, SS98, GI94, SAS95o]. Conceptual [Oli10]. Concise [Fri91b]. conditional [UM14]. conduct [NB18]. Conducting [Cle03b, Cle03a]. Cone [Tru88]. Conference [ACM92, ACM96, Ano96a, Ano96b, Bjø01, CM82, Gra96, Hut92, Ogl93, Ona90, SUG82, SUG83, SUG84, SUG85, SAS88a, SAS88y, SUG89, SUG91, TD85, Tie88, Tie89, Ano95a, Ano97, Ano98, DH94]. Conferences [SUG87]. Confidence [SB01, GSD04]. Confounding [Laz24, Sha22]. Confusion [SB96]. Conjoint [SAS93c]. CONNECT [SAS90r, SAS93-28, SAS94w]. Constable [Rut12]. Constrained [Gre17, Rec10a]. Constructing [GGS23]. Construction [WWG09]. consultant [SAS94k]. Consumer [Ano00b, How88]. context [BVFB19]. Contiguity [Rec10a]. Contiguity-Constrained [Rec10a].

Contingency [LS02]. Continuous

[VGDM^{+}18, TPAM07]. Control [Kas98a, Kas98b, SAS91j, SAS92f, SAS94n, Sta94, Zie97e, SAS90l, SAS90x, SAS91n, SAS95s]. conversion [SHO^{+}96]. converting [SAS88s]. cooling [Sou94]. Copenhagen [Tie88]. copyrights [Sam12]. Corporate [ACM96, SBJ85]. Corporation [Ano00b]. Correlated [AS04, Ano08a, KM16, KS10, Sho18, Len15, Lum08, SC07, Sta21]. Correlation [Gos14, Hoh85]. cotton [Bir94]. Count [Hil14, KM16, THT12]. Course [Cha04b, HH15, Pri05, Utl05, Car93, EWC94, Har93, HH04, LJ95, LW95, SAS94-34, Spe04, Cha04a]. Covariance [ZK12]. Covariate [Gre17]. Covariate-Constrained [Gre17]. Cox [TR14]. CPE [SAS93-29, SAS94x]. CRC [Ano06b, Kel10, Laz24, Liu17, Lum08, Oli20, Oli17, RHWP24, Rah24, Rob08, Sha22, Sta21, War05, Woo11, Zha19]. Creating [Lan82, MYM21, SWH^{+}11, SAS83b]. Criteria [Gos14, MJ97]. Critical [SB01]. Cross [MYM21, WWG09, Whi94, Dru79, SAS94v]. Cross-Over [WWG09]. cross-platform [SAS94v]. cross-section [Dru79]. Cross-Tabulations [MYM21]. Curves [Ami87, Gho10, Hal08, Kar08, Len10, NZ88, BR09]. customer [SAS91d, SAS91e].

D [Ano06a, Bos07, Chi02, Hei90, Hor09a, Hor09b, Lou06, Lu08, McF01, Neu06, Nie08, Rob07, Sch06b, Zha19, Zie96a, Zie97g, Zie99c, Zie00b]. D. [Ano11b]. D.Tobias [Chi02]. D’Agostino [Bie08, Che08, Bun07, Sta07c]. Dale [MH05]. DATA [MT14, AR91, Ami87, Ano06a, Ano06d, Ano06b, Ano08a, Ano09d, Ano11a, BS95, BH14, CBDM07, Cla03, Dam09, Den12, DE05, Die20, FZZZ09, Fer03, Fri06, GHR18, Har08, Hec11, HH15, Hil14, Jen08, KM16, Ke11a, Kha18, Kha20, KS10, KH10, Lew07, Liu17, Lu19, MH05, Muk15, Mwi11, Neu06, Pri05, Rah24, San03, Sch06c, Sho18, Sin08, Spa06, Ste11, Ste06, SA87, Tur12, Utl05, VGDM^{+}18, Wie06, Wie08, ZCI^{+}16, Zie89, Zie96b, Zie97c, Zie99d, Zie01b, Zie01a, Zim07, Agr13, Bir94, CHM95, DH96, Dru79, Fan94, GG82, GGS23, Gre81, HH04, JXW^{+}11, Lan81, Lan82, LJ95, Law03, Len15, Lum08, MK08, MPT06, Mer83, Mir93, Mwi12, NTGD18, Nie11, NMvdH14, Pel95, Rob08, Ros83, SAS91s, SAS92h, SAS93q, SAS93r, SAS93s, SAS95c, Sal79b]. data [SH80a, SH80b, SC07, Spi96, Sta21, Sta86, SDK95, THT12, Tru88, Uni94a, Uni94b, WWM^{+}24, WB85, SAS96c, Bur92, KW85, Mon97, Sau04b, Cab04, Jen09, Nas08, RHWP24, Sta07a, War05]. database [JXW^{+}11, SAS90o, SAS93v, SAS93w]. databases [SAS94t]. DATACOM/DB [SAS90n]. Datenanalyse [Hac84, Ric08]. Datenanalysessyteme [KW85]. Datenverarbeitung [Hac84]. David [Ano04a, Dow88]. Davis [Sch06c, Ste06, Zie96b, Zie01a, LaL09, Sta07d]. Day [Cha04a, Cha04b, Spe04]. Dayanand [Ano99, Zie00a, Zie01b]. DB [SAS90n]. DB2 [Bro89, SAS94u, Ser93d]. DC [ACM92]. Deauville [MS12]. Debugger [Ano00b, SAS88g, SAS93b]. Debugging [Bas03]. December [Bir94]. Decision [SBJ85, SAS96a, Jen08]. decomposition [Che88, CS89]. degrees [Har81]. Delivery [Ano08b, Pra08, BI07, SAS93p]. Delwiche [Hor09b]. Denmark [NF07]. Dennis [Zie97e]. Density [VDT14]. departures [CL90]. depositions [GB82]. derivatives [SAS88z, SAS90d, SAS91d, SAS91x]. Description [SMHBR06, TD85, Tru88]. Descriptive [Dha10, Can93]. Design [Bar83, Bra15, Gil12, Hu09, LVK^{+}11, Law10, O’B08, WWG09, Cha08, Kel10, Rob09, SAS94-28, Dra11, Gie11, Grö10, Len09, Woo11]. Designs [AC08, FS13, Gre17, Hor09a, Law08, Lu08, Oli10, Puk07, Nie08, Wit92]. Deskriptive [Wei94]. Developing [LS05, SAS95d]. development [KB95, SAS95o]. developments [Fra79]. devices [SAS93-38]. diagnostics [HP06, PG05]. Diary [RMM18]. Dickey [Ano04a, Dow88]. dictionary [SAS95p]. Diego [Ano97, Cou03]. Differences [PKC^{+}20]. different [GSD04, NTGD18]. digitizer [Lan82]. DiIorio [LaL92]. Ding [Rah24]. Ding-Geng [Rah24]. Dirk [Kha15]. disclosure [GL76]. Discovering [FMF12, BK94]. Discrepancies [WJ19]. Discrete [SA87]. discriminant [Whi94]. Discrimination [Zie01b]. DISPLA1 [OK80]. Display [HH15, Pri05, Utl05, HH04, Ser93b]. Displays [Fri06]. Distance [SH09]. Distance-Aided [SH09]. distributed [HP92]. Distributions [Pan09, Har81]. DKK [Chi02]. DL/I [SAS91k, Sto82, SAS95m]. DM [Ano93b, Ano93a]. Dmitrienko [Bie08, Bun07, Che08, Chi06, Dur06, Jon07, Sta07c]. Do [Sam12]. Document [Pop97]. documentation [Bir94, RMM18, SAS87, SAS91g, SAS92b]. documenting [SAS92c, Uni92, Uni93]. Doing [SAS92a]. domain [Che88, CS89]. Donev [Hor09a, Lu08, Nie08, Puk07]. Donors [SH09]. Dose [AR91, Ami87]. Dose-Response [AR91, Ami87]. Dothang [RHWP24]. download [OGR89]. Driven [Zim07, KB95]. driver [Ano94b, Ano95c]. Dror [Chi02]. DS [SAS90p]. dwarf [Ano00b]. dyadic [MPT06]. Dynamic [Ano00b]. dynamics [CS93].

e-book [Kha20, Zha19]. easy [Spi96, Zie00c]. econometric [Dri95, SAS93-35, Sal79a]. econometrics [Hil93, WHG88]. Economics [Zie97a]. ed [Ano09a, Cow04a, Cow04b, Hor09b, Oli20, RHWP24, Rut12, Sau11, Sta21, Hil93, Gup02b, Gup02a]. Edited [CM82, Bun07, Lou06]. Edition [Mai12, McF01, SAS85f, Wil09, AK85, Ano99, Ano08a, Ano09b, CH81, CA85, Eas87, Joy85, Kal86, Lou06, LS87, SAS84a, SAS84c, SAS85j, SAS85b, SAS85i, SAS85c, SAS87, SAS88i, SAS88k, SAS88j, SAS88m, SAS88n, SAS88o, SAS88p, SAS88l, SAS88q, SAS88b, SAS88h, SAS89g, SAS94c, Sal81b, SLMN86, WHG88, Bra15, Ros15]. editor [SAS91t]. edn [Bos07, Lum08]. eds [Bie08]. Edward [Ano06c, Len15]. Effective [Nat93, SM82]. Effects [Ano06a, LG11, Neu06, VDT14]. Efficient [Nov03, SAS90k]. Eighteenth [Ogl93]. eighth [SUG83]. Einführung [KW85, Wei94, MB92, Hac84, Ano93b]. EIS [SAS93d, SAS93-30, SAS93-31, SAS93-50, SAS94a, SAS95r, SAS95s]. elasticity [CS89]. Elementary [Hei90, Zie99c, DCT89a, Dud86, DCT89b, SL87]. Elimination [JMD08]. Elizabeth [Zha19]. Ellipses [Fri06]. elrm [ZMG07]. Elst [Oli17]. Embedded [Ano00b]. Embedix [Ano00b]. emergency [Sou94]. EMJ [Ano00b]. Emrouznejad [O’B13]. Enables [Wei12]. End [Ano00b, Kha20]. End-to-End [Ano00b, Kha20]. endpoint [Oli17]. endpoints [TPAM07]. engine [Ros83]. Engineers [Bzi11]. ENGLISH [SAS93-32, SAS93-34, SAS93-33]. enhanced [SAS96c]. Enhancements [SAS83a, Ano84c, SAS80, SAS81a, SAS82a, SAS82d, SAS82e, SAS88t, SAS88u, SAS91a, SAS91m, SAS91r, SAS91v, SAS91s, SAS91n, SAS92j, SAS92g, SAS93b, SAS93a, SAS93-28, SAS93-44, SAS93n, SAS93o, SAS94h, SAS94i, SAS94r, SAS94s, SAS94-27, SAS94p, SAS94-33, SAS95g, SAS95h, SAS95q, SAS95t, SAS95w, SAS95k, SAS96d, Sto82]. Enough [Bzi11]. enquête [CP94]. ENTERPRISE [Mwi11, Dix07, Jen08, Jen09, Kar09, LaL09, RHWP24, Rut12, Sin08, Spe08, Sta07d]. enterprises [CL90]. entry [SAS93z, SAS93-27]. Environment [Wei12, Rai95, SAS90a, SAS90b, SAS90c, SAS90d, SAS91d, SAS91x, SAS92d, SAS93a, SAS93h, SAS93j, SAS93i, SAS93k, SAS93-38, SAS93-45, SAS93-46, SAS93-51, SAS94h, SAS94i, SAS94x, SAS94-29, SAS94-30, SAS94-31, SAS95e, SAS95g, SAS95h, SAS95-27, SAS95-28, SAS95-29]. Environments [Har93, SAS94-34, SAS93f, SAS93l, SAS93m, SAS94-32, SAS95-30]. EPA [Tru88]. Epidemiologie [Wei94]. Equation [Hec96, OH13, Ros15, Hat94]. equations [PYC93]. ERA [LV16]. Error [SAS88k]. eServer [Ano00b]. Essentials [Sha20]. Estadistico [Mar83]. Estimation [Liu12, OC15, TR14, XM14, GSD04, Sil96]. Estimators [Jur15]. Etnus [Ano00b]. ETS [All82b, SAS84c, SAS85h, SAS88x, SAS93-35, SAS93-36, SAS94y, SAS95t, SAS95u]. European [Tie88, Tie89]. evaluating [XB10]. Evaluation [NTGD18, BvBGOG03, Mer83, Mer84, Oli17, SAS94-35, TPAM07]. Evaluations [KW85]. Event [GHR18]. Everitt [Ano06b, Ano09b, Len97, Rob08, Zie97b, Ano06d, Ano09a, Kar09, Lew07, Pin10, Ric13, Zie02]. Evolution [Hud89]. Exact [ZMG07, BBV17]. Exact-Like [ZMG07]. Examination [WJ19]. Example [Gur08, Hol08, Ke11b, Sta07b, CPI95, Dob09, Khu13, Zie96f]. Examples [Ano08b, Dob06, Fri06, HH15, Liu12, Lu19, Pra08, Pri05, Utl05, Wal02, Zie96d, Zie97a, BI07, HH04, SAS89h, SAS93c, SAS93-40, SAS94l, SAS95c, SAS95f, DL04, Ano03a, Ano12, Zie99a]. Excel [Sha20]. excellent [Sha22]. Exclusion [FTP02]. Execsolutions [SAS96a]. Execute [NZ88]. Execution [SR07]. Exercises [ACL04]. Expanded [Mer84, Ano88]. experiences [Wit92]. Experimental [AC08, Hor09a, Lu08, Puk07, Nie08]. Experiments [Dra11, Gie11, Grö10, Law10, Len09, Gil12, SAS94-28, Woo11]. experts [SAS94j]. Extended [LV16]. Extending [SAS94a, Zie98b]. extension [MJ97]. extensions [SAS94d].

F [Len15, Zie97c, Har81]. facilitate [SM82]. facility [SAS83a, SAS91d, SAS91e, SAS93t, SAS94q, SAS94s, SAS94m]. Factor [Hec96, LS02, OH13, PKC^{+}20, Ros15, Hat94]. Factorial [FS13]. Factorials [LSC09]. Fan [All06, Ano03b, Sau04a]. FAQ [Bee05]. faster [Opd13]. features [Fra79, SAS80]. February [SUG82, SUG91]. Felsovályi [All06, Ano03b]. Felsvlyi [Sau04a]. FEMCAD [SAS88a]. FEMCAD-88 [SAS88a]. Fernandes [War05]. Fernandez [Ano11a, Cab04]. Fetterman [Gun03, AC04]. Few [Sha20]. field [SAS94j, Tru88]. Fifteenth [Ona90]. file [SAS83b, SAS93x, SAS94r, SAS95n]. files [Ast94a, CHM95, SAS89i, Uni94a, Uni94b]. financial [CL90, GL79, SAS93p, Sha20]. First [Ano96a, Ano96b, Gra96, Le 94, SAS94c, SAS89g]. Fischer [Ano93a]. Fit [GHR18, XB10]. Fitting [MH05, Pat02, CS93]. Fixed [Ano06a, Neu06]. FL [Oli20, Rah24, Rob08, Sha22, Sta21, Ano95a]. Flexible [TPAM07, VDT14]. Florence [CM82]. Florida [Ano06b, SUG84, SAS88y]. fmixed [LG11]. Forces [ACM96]. Forecasting [Ano04a, Dow88, Ven01, Bir94, BD86, BD03, SAS93-35, SAS95u, Zie97a]. forma [GL76]. formation [Ser93b, Ser93a, Ser93c, Ser93d]. formats [SAS93x, SAS94r, SAS95n]. Formatted [Sun98]. formulae [PG05]. FORTRAN [SB01, LS05]. Foundations [FMT02, Ast94b]. Fourteenth [SUG89]. Fourth [SAS88a]. Fraction [LVK^{+}11]. Fractional [FS13, Har81, SMHBR06]. FRAME [SAS93z, SAS93-27, SAS95o, SAS95p]. Framework [Ano95b]. France [Cou01, CM11, MS12, Van97]. Francis [Ano06b, Oli20, RHWP24]. Francis/CRC [Ano06b]. Francisco [SUG82, SUG89]. Frank [Ano93b, LaL92]. free [OGR89]. freedom [Har81]. frequentist [WWM^{+}24]. Freund [Mar88, Mas88, Zie92a, Zie92b, Zie01c, Ale91, Moo03]. Friendly [Sau04b, Zie93]. FSCALC [Ano84d]. FSP [SAS82c, SAS88h, SAS89f, SAS91n]. full [SAS82c]. full-screen [SAS82c]. Functional [Hud89, LG11, SHD88]. Functionalities [BMGT15]. functions [Har81]. fundamentals [Har93, LW95, SAS94-34].

G [Ano06b, Ano09b, Ano09d, Ano12, Kar09, Lac99, Len97, Lew07, Lou06, Oli10, Pin10, Rob07, Sch06c, Ste06, Zie96b, Zie97d, Zie97b, Zie01a]. G.lite [Ano00b]. Gamst [Oli10]. Gary [Ste06, Zie01a]. GAUSS [Dri95]. GAUSS-386 [Dri95]. GBP [Oli17]. Gearheart [Kha20]. GEE [HP06, NMvdH14]. geepack [NMvdH14]. Geert [Chi06, Dur06, Jon07, Oli17, Ste99]. General [Ahm08, Gre17, Nor89]. Generalized [GHR18, SA87, VDT14, LV17, PG05, Len15]. Generating [Can93, KM16, LS02]. Generation [FS13]. Generic [MYM21]. Genetic [Jur15]. Geng [Rah24]. GENLIN [NMvdH14]. GENMOD [NMvdH14, SAS93-48]. Gentle [LaL92, DiI91]. Geoff [Ano06d, Ano09a, Ric13, Rob08, Zie02]. George [Ano11a, Bos07, Cab04, War05, Zha19]. Gerhard [Nas08, Sta07a]. German [Göt92, KW85, MB92]. Germany [CS96]. Getting [SAS91c, SAS91b, SAS93d, SAS94b, SAS95e, SAS96b]. GIS [SAS95v]. glance [And89]. Glasgow [Myc95]. Glenn [Ano03a, DL04, Zie99a]. GLM [FL81, SH80a, Zie89]. GLMM [XM14]. Global [Ano00b]. GLUMIP [KCM08]. Gönen [BR09, Kar08, Len10]. Gonena [Gho10]. Goods [JLC91]. Göttsche [Ano93a]. Gra [LV17]. graf [SAS76]. grafischen [Ric08]. Graph [CH81, Joy85, Zie96g, Zie96e, SAS81d, And89, Ano84c, CSS95, Eas87, Lan81, Lan82, Mir95, SAS82d, SAS82e, SAS83c, SAS88i, SAS88j, SAS88w, SAS90t, SAS90s, SAS91u, SAS91r, SAS91s, SAS91t, SAS93-37, SAS93-38, SAS94d, SAS94z, dV84]. Graphics [Ano08b, BI07, Fri91a, KH10, Pra08, OK80, SAS91t, SAS93-38, SAS94g, Liu17, Den12, Hec11, Ste11, Zie93, Tur12]. Gravely [Zie99e]. Griffiths [WHG88, Hil93]. Griffiths/Hill/Judge [Hil93]. Group [Ano96a, Ano96b, Ano97, Ano98, DH94, DS98b, Gra96, Hut92, Ogl93, Oli20, Ona90, RHWP24, SUG82, SUG83, SUG84, SUG85, SUG87, SAS88y, SUG89, SUG91, SS98, Tie88, Tie89]. Growth [Ami87]. Guarino [Oli10]. Guide [All06, All95, Bra08, Bro89, Bzi11, CH81, Dix07, DCSD07, Dob06, Fri91b, Gur08, Joy85, Kar09, SAS88g, SAS88k, SAS88j, SAS88n, SAS88o, SAS88p, SAS88l, SAS88q, SAS88b, SAS88h, SAS91d, SAS91e, Spe08, All82b, All82a, AK85, Ano88, Ano94b, Ano95c, AA90, Bar76, Bar79, BK94, Cou80, CA85, CB85, EK86, Eas87, Fow93, FL81, Har93, Hel77, Hel78, HR79, Hel83, Hel85, Hil93, Kal86, Laf02, LS87, MPT06, Mer83, Mer84, Rad94, Ray82, Rei80, SAS81c, SAS81d, SAS82c, SAS82b, SAS84c, SAS85a, SAS85g, SAS85h, SAS85j, SAS85b, SAS85i, SAS85c, SAS85d, SAS85e, SAS85f, SAS88x, SAS88m, SAS88v, SAS89d, SAS89a, SAS89g, SAS90e, SAS90f, SAS90w, SAS90j, SAS90k, SAS90v, SAS91x, SAS91p, SAS91o, SAS92e, SAS92i, SAS93-32, SAS93-34, SAS93-35, SAS93-36, SAS93-39, SAS93-42, SAS93-33, SAS93-43, SAS94f, SAS94g, SAS94k, SAS94l, SAS94-34, SAS95i, SAS95j, SAS95x, SAS95-31]. guide [Sau04a, Shi96, WB85, Yor94, Mwi11, ACL04, Ano11b, Bun07, Che08, Chi06, Cla03, Dur06, Hol08, Kha20, LaL09, San03, Sta07c, Sta07b, Van04, Bie08, Dob09, Jon07]. guidebook [CHM95]. Guide(R)

[Rut12, Sta07d]. Gustav [Ano93a].

H [Cha04a, Chi02, Hac84, Kha15, Rah24]. H.-M [Hac84]. HADOOP [CZZ13]. Hall [Kel10, Lum08, Oli20, Oli17, Rah24, Rob08, War05, Woo11, Zha19, Len97]. Hall/CRC [Kel10, Lum08, Oli20, Oli17, Rah24, Rob08, War05, Woo11, Zha19]. Handbook [Bas03, DE08, Gil99, Gup02a, Gup02b, MT14, Pin10, Wil09, Zie02, DE02, ED96, Gil96, Hil93, Len97, WHG88, Ano09a, Ano09b, Zie97b]. Hands [Kha20]. Hands-On [Kha20]. Hardback [Oli17, Nie08, Sta21]. hardcover [Sha22]. hardware [SAS88i, SAS88w]. Hardy [Zie97c]. HARVEY [SH80b, Sea80]. Hatcher [ACL04, Ano06c, Zie95a, Hec96, Van04, ACL04]. Hawaii [Hut92]. hbk [RHWP24]. Health [Dam09, Rah24]. Heiberger [Utl05]. held [CEM82]. help [Rad94]. Herzberg [Tho92]. Hidelberg [Ano93b]. Hierarchical [Rec10a, Rec10b]. Higgins [Kar07]. High [XM14]. High-Performance [XM14]. Highlighting [SAS94c]. Highway [CHM95]. Hill [WHG88]. Hochberg [Chi02]. Hogback [Tru88]. Holland [Utl05]. Hollywood [SUG84]. Homogeneity [Ami87]. Honolulu [Hut92]. Horton [Liu17, Ste11, Den12, Hec11]. host [SAS89i]. Hotel [All86, Bjø01, DH94]. Household [JLC91]. How-To [Zie96g, Mir95]. Howard [Ke11b]. HPGLIMMIX [XM14]. HTML [BMDB98, LS98, Pop97, Spi96, Sun98]. http [Kha15]. Huffcutt [Cle03a]. Hungary [TD85]. Hüthig [Ano93b]. Hyperlink [LS98]. Hypothesis [Kha15, Tae14].

IBM [SAS83c]. ID [Tru88]. identifying [PG05]. IDMS [SAS95l, SAS96c]. IFIP [TD85]. ihre [Ric08]. II [Fra79]. Illinois [Gra96]. Illustrations [CL90, GL76, GL79]. Image [SAS94d]. IML [Gen11, Sea11, Sea12, AS02, CA85, CB85, HP06, JMS^{+}09, KCM08, NZ88, NO06, Pan09, Per10, SAS85i, SAS88l, SAS90u, SAS91v, SAS95w, Sal88, SB01, ZK12]. Implement [BMGT15, VDT14]. Implementation [Bra15, Hu09, KM16, O’B08, Cha08, Kel10, Mil93, Rob09, Wei82]. Implementing [oCPA96, JMD08, SHD88, ZMG07]. improvement [SAS95y]. Imputation [BH14, Muk15, SH09, WJ19, Yua11, BvBGOG03]. IMS [SAS91k, SAS95m]. IMS-DL [SAS91k, SAS95m, Sto82]. IMS-DL/I [SAS91k, SAS95m]. Inc. [Ano00b]. Inc./John [Ano99]. included [Dob09]. including [MJ97]. income [Sou94]. incomplete [Wit92]. independence [LS02]. index [SAS87, SAS91g, SAS91h, SAS92b]. individual [WB85]. industry [RHWP24]. Inference [Liu12, VGDM^{+}18, ZMG07]. influential [PG05]. Informatics [CM82]. Information [JLC91, CHM95, SAS93p]. Informix [SAS94q]. INGRES [SAS91l]. Input [Bar83]. INQUIRE [Ros83]. INSIGHT [LJ95, SAS93-39, SAS94-27, SAS95x]. Installation [SAS93e, Ast94b]. Institute [Ano94a, Ano99, BR09, Bos07, Chi02, Dob09, SAS94f, SAS94g, Sau04a, Sau04b, Sch06b, Zie96d]. instructions [SAS93e]. Integrated [AC04, Gun03, GG82, MF02, MF03]. Intelligence [Jen08]. Interactive [LJ95, Sal88]. interest [Fra79]. Interface [All86, Fra79, JLC91, SAS76, SAS90m, SAS90n, SAS90o, SAS90p, SAS90q, SAS91k, SAS91l, SAS92h, SAS93q, SAS93r, SAS93t, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS94q, SAS94r, SAS95l, SAS95m, SAS96c]. interfaces [Ast92, SAS88i, SAS88w, SAS93m]. Intermediate [Ano09d, Har08, HH15, Ke11a, Pri05, Utl05, HH04, MK08]. intermittent [UM14]. Internal [Ano00b, KCM08]. International [ACM92, ACM96, AV08, Ano96a, Ano96b, Ano00b, CM82, CF99, CS96, Cou01, Cou03, CM11, Gra96, HP02, Hut92, Le 94, Lev98, MS12, Myc95, NF07, Ogl93, Ona90, Pal00, PS09, SUG82, SUG83, SUG84, SUG85, SAS88y, SUG89, SUG91, SAS93f, Tie88, Tie89, Van97, Yah11, Yi06, Ano97, Ano98, Gia04, HS05]. Internet [Ano00b, Sha06]. Interval [Lu19, GSD04]. Interval-Censored [Lu19]. Intervals [SB01]. IntraLinux [Ano00b]. Introduccion [Mar83]. Introducing [SAS91f]. Introduction [Fri91b, Hil93, KW85, MPV12, Ros10, Ros14, SAS90s, SAS94e, Ser93a, Wan19, WHG88, Zha19, DiI91, Her90, LaP81, MB92, SAS90g, LaL92, Sha22, Sin08, Ven01, Zie96c]. introductory [Hel78, Hel83, Hel85, SAS85b, SAS88m, SAS90w, SAS94f]. iNUX [Ano00b]. inverses [Har81]. Iorio [Zie97c]. IQ2 [Ano00b]. ISBN [Ano93b, Ano93a, Ano06b, BR09, Bie08, Bos07, Chi02, Dob09, Jon07, Kel10, Kha15, Kha20, Laz24, Len97, Lum08, McF01, Nie08, Oli20, Oli17, RHWP24, Rah24, Rob08, Sch06b, Sha20, Sha22, Sta21, Til96, Woo11, Zha19]. ISPF [SAS95d]. Italy [CM82, CF99, Gia04, Lev98, Yah11]. item [Wei82]. Iverson [Ano08b]. IX [RHWP24, Bie08, Sha22].

J [And86, Ano06c, Den12, Gen11, Hec11, Hor09b, Kar07, Liu17, Mas88, Oli10, Oli20, SUG91, Sea11, Sho90, Zie01c]. J. [Ano09d, Ano12, Gen11, Gie11]. Jackie [Ano08b]. Jaffe [Hec90, Sho90, Zie95b]. James [Kar07, Kha20, LaL09, Sta07d]. Jamis [Sea11]. January [SUG83]. Japanese [Ano00b]. Java [Ano08b, BI07, Pra08]. JAWS [Ano00b]. Jay [Hec90, Zie95b]. Jeffrey [Ano06d, Rah24, Zie98a]. JetForm [Ano00b]. Jin [Sha22]. JM [GHR18]. JMFit [ZCI^{+}16]. JMP [SAS94f, SAS94g]. JMP(R) [Ano06c]. John [Ano99, Ano04a, Dow88, Dra11, Kha15, Ng12, Sha20, Woo11, Gil12]. John-Ran [Sha20]. Joint [ZCI^{+}16]. Jointly [GHR18]. Joy [Oli17]. Jr [Cle03a]. Judge [Hil93, WHG88, Hil93]. Judge-Hill-Griffiths-Lutkepohl-Lee [WHG88]. Judge/Hill/Lutkepohl/Lee [Hil93]. July [ACM92, AV08, Cou01, Pal00]. June [Cou03, Pal00]. Just [Bzi11]. JWAVE [Ano00b].

Karl [Rah24]. Kecskemet [TD85]. Keenan [Sau04a, All06, Ano03b]. Keith [AC04, Gun03]. Ken [Hec11, Liu17]. Kennedy [Ano09d, Ke11a]. Kentucky [All86]. Khattree [Ano99, McF01, Zie96a, Zie00a, Zie01b]. Kim [Ahm08]. King [Zie95c, Zie97e]. Kleinman [Den12, Hec11, Ste11, Liu17]. knowledge [SAS93-32, SAS93-33]. Koch [Sch06c, Ste06, Zie96b, Zie01a]. Kolosova [Zie99b]. kompakt [Ano93a, Göt92]. Kongens [NF07]. Korea [Yi06]. Korosteleva [Sha22]. Krämer [Ric08]. Kruskal [MJ97]. Kuhnt [Kha15].

L [Oli10, Zie95c, Zie95a]. Lab [EM09, Ell95, Ell00, EM10, SAS92i]. laboratory [GI94]. Language [SAS88n, SAS88o, SAS91j, SAS92f, SAS94n, Sal88, AK85, Bas84, Car93, CS85, CS87, CS91, CS97, SAS89b, SAS90g, SAS90h, SAS90i, SAS90l, SAS90x, SAS91i, SAS91n, SAS93l, Sha22, SLMN86, Sta94, And86, Ano06d, Zie98a]. Languages [Hud89, Hsu86, SHD88]. Lantronix [Ano00b]. Large [Pop97, Sal79b]. Larry [ACL04, Ano06c, Hec96, Van04]. Lars [Ric08]. Lava [Ano00b]. Law [CM82]. Lawal [Ste06]. Lawson [Gie11, Woo11, Dra11, Gil12]. Leacky [Oli17]. Learning [Ell95, Ell00, EM09, EM10, Gur08, Hil93, RHWP24, Ale91, Dob09, Hol08, Sta07b]. Least [KS10]. Lecture [Ste99]. lecturers [RHWP24]. Lee [Hil93, Sha20, WHG88]. Legally [Sam12]. Lehman [Ano06c]. Level [FS13, JMD08, LSC09, RHWP24, Rec10b, SAS88g]. Lexington [All86]. Li [Laz24, Sha22]. Library [SAS88d, SAS88e, SAS88f, SAS91d, SAS91e, GG82, Hel77, Rei80, SAS82a, SAS89d, SAS89e, SAS91p, SAS93b, Sal79a]. Lifetime [Wan19]. LIHEAP [Sou94]. Like [ZMG07]. Likelihood [Jur15, Liu12, OC15]. Linear [Ahm08, Che92, Cow04a, Cow04b, Fri06, Gen11, JMS^{+}09, Mar88, Moo03, Pat02, Per10, Rec10b, Sch06a, Sch06b, Sea11, SA87, VDT14, Zel02, Zie92a, FL81, FLS86, Len15, LFS91, LSF02, MPV12, PYC93, PG05, WWM^{+}24, Lac99, Ste99]. Lineo [Ano00b]. link [WB85]. Linux [Ano00b]. Linux-based [Ano00b]. listing [Bir94]. Literate [LH07]. Littell [Bos07, Hei90, Lou06, Mar88, Moo03, Rob07, Zie92a, Zie97d, Mas88, Zie92b, Zie99c, Zie01c]. Little [DS98a, Dix07, Hor09b, DS95]. Log [Sch06a, Sch06b, Zel02]. Logic [CM82]. logiciel [Ser93a]. logiciels [Laf81]. Logistic [Dha10, SAS95f, SED14, ZMG07, XB10, SB96, Zie96d, Zie00b]. London [HS05, Len97]. Longitudinal [Ano06a, CBDM07, GHR18, Neu06, OC15, UM14, ZCI^{+}16, FLW11, NMvdH14, WWM^{+}24, Cla03, San03]. look [Ast94a]. Lora [Hor09b]. Lotus [Fri91b]. Louisiana [SUG83, SUG91]. Low [Sou94]. Low-income [Sou94]. lrasch mml [OC15]. Ltd [Ano00b]. lunch [OGR89]. Lutkepohl [Hil93, WHG88]. Lyngby [NF07].

M [And10, Ano08a, Ano09d, Bas03, Hac84, Kar08, Lum08, Sch06c, Sta21, Utl05, Ven01, Web12, Zie93, Zie96b, Zie00c]. M. [Hu09]. Machine [RHWP24, SHD88]. Macintosh [SAS94f, SAS94g, SAS95-33]. Macro [Dha10, GHR18, Gre17, Jur15, KS10, LG11, LV16, MH05, MYM21, OC15, Pan09, PKC^{+}20, SH09, SR07, VDT14, VGDM^{+}18, Wei12, XM14, ZCI^{+}16, Zie00c, ZK12, dSM15, BBV17, Car93, Kal86, LV17, MJ97, SAS83a, SAS90e, SAS94m, SLMN86]. Macros [FCY20, LVK^{+}11, Law08, NO06, NB18, GSD04, KK96, SWH^{+}11, Tam94]. Made [Zie00c]. Madrid [HP02]. magazine [SAS96a]. magnitude [Opd13]. Mainframe [Ami87]. major [Pel95]. makers [SAS96a]. Making [Mon97, Bur92]. Man [Woo05]. Management [Den12, Hec11, KH10, Liu17, Ste11, And89, Ano84a, Fan94, Gre81, SAS93g, SAS93-42, SAS93-40, SAS93-43, Tur12]. management/manipulation [Fan94]. manager [SAS93v, SAS93w, Ser93b]. manipulation [Fan94]. manual [Dud86, GI94]. map [Lan81, Lan82, SAS91s]. mapas [dV84]. Marasinghe [Ano09d, Ke11a]. Marc [Oli17]. March [All86, Ano96b, Ano97, Ano98, Gra96, SUG84, SUG85, SAS88y]. Marginal [KAK05, OC15]. Marginalized [VGDM^{+}18]. Mark [Kel10, O’B08]. marker [TPAM07]. Market [Zie96c, Zie97f, SAS94e, SAS94-35, SAS95-32]. marketing [How88, SBJ85]. Marshaling [ACM96]. Martin [SUG91]. Master [SAS87, SAS91g, SAS91h, SAS92b]. Mastering [Jaf89, Jaf94, Jaf96, Hec90, Sho90, Zie95b]. matching [BBV17]. Mathematics [Dud86]. Matignon [Jen09]. MATLAB [Bjø01]. Matrices [Che92]. Matrix [Sal88]. Maura [Ste06, Zie01a]. Maximum [Jur15, OC15, Liu12]. May [Ogl93, TD85, Tie89]. McGee [Ven01]. Mean [PKC^{+}20]. Measures [KAK05, Rec10b]. Mediation [FCY20, Laz24, Sha22]. Medical [Ano06d, Ano06b, DE05, JS13, Lew07, Liu04, Ric13, Spa06, Wie06, Zie98b, Fra14, Rob08]. medizinischen [Wei94]. Menu [SAS95y, SAS93-41, SAS94-28]. menus [Spi96]. merge [Sal78a]. merged [Pie87a, Pie87b]. Mervyn [Ke11a]. Messages [SAS88k]. Meta [Cle03a, Cle03b, KK96, SWH^{+}11]. Meta-analysis [KK96, SWH^{+}11]. method [Che88, CS89]. methodologies [TD85]. Methods [And10, Ano03a, Ano06a, Ano12, Der10, DL04, KAK05, Laz24, Neu06, Wal02, Web12, Zie99a, BvBGOG03, MK08, NTGD18, Oli17, SAS93-50, Sha22, Tam94, WWM^{+}24, Ano09d, Ke11a, Har08]. Mexicana [dV84]. Meyers [Oli10]. Michael [Ano93b, Sau04b]. Michele [Bas03, Zie00c]. Microcomputer [Hsu86]. Microsoft [SAS93e, SAS93h, SAS93j, SAS93i, SAS94h, SAS94-30, SAS94-34, SAS95g, SAS95-28]. MIDAS [SH09]. Migration [Ser93c]. Millar [Liu12]. Milliken [Bos07, Lou06, Rob07, Zha19, Zie97d]. Miner [RHWP24, Sin08]. Miner^{T M} [Jen08, Jen09]. minicomputers [Ano84b]. Minimum [FS13]. Minimum-Size [FS13]. Mining [Ano11a, Cab04, Jen08, Jen09, Sin08, War05, Fer03, Mwi12]. Minitab [GL14, Sha20, DCT89a, DCT89b, HP92, Fri91b]. Miron [Zie96g]. Missing [BH14, CBDM07, Muk15]. missingness [UM14]. Mithat [BR09, Gho10, Len10]. Mixed [FS13, GHR18, Lac99, LSC09, LG11, Lou06, Rob07, Ste99, SMCW18, VDT14, Wul07, Zha19, Zie97d, Bos07, Lit06]. Mixed-Level [FS13, LSC09]. Mixture [BMGT15]. MLGA [Jur15]. MOD [JLC91]. MODECLUS [SAS93-49]. Model [MH05, CS93, PG05, QDG17, SMHBR06, XB10]. Modeling [And10, Der10, Hil14, OH13, Sel11, Web12, Bir94, Hat94, MPT06, Par06, SAS93-35, Sha22, Hec96]. Modelling [Ros15]. Models [Ahm08, BKL05, BMGT15, Cow04b, Fri06, GHR18, JMS^{+}09, KAK05, Lac99, LS02, LG11, OC15, Per10, PKC^{+}20, Rec10b, Sch06a, Sch06b, Ste99, SMCW18, TR14, VDT14, Wul07, ZMG07, Zel02, ZCI^{+}16, Zha19, Bir94, FL81, FLS86, GSD04, HP92, Len15, LFS91, LSF02, Lit06, Ros10, Ros14, UM14, WWM^{+}24, Cow04a, Gen11, Lou06, Mar88, Moo03, Rob07, Sea11, Zie92a, Zie97d, Bos07]. Modem [Ano00b]. Moderation [Laz24, Sha22]. Modern [DCT89a, Dud86, DCT89b]. Modernization [JLC91]. modifying [SAS95c]. Mohamed [Lum08, Sta21]. Mohammad [Lum08]. Molenberghs [Chi06, Dur06, Jon07, Oli17, Lac99, Ste99]. Monte [All06, Ano03b, Sau04a, JXW^{+}11]. Most [Bur92]. MTPmle [VGDM^{+}18]. Muchene [Oli17]. Muenchen [Ano09c, Mai12, Ree09, Sau11, Voe09]. Muller [AC04, Gun03]. multgee [NMvdH14]. Multicluster [Che92]. multilevel [MPT06]. multimodality [NB18]. Multiple [BH14, Chi02, FCY20, HP92, NZ88, SH09, SB01, WJ19, Yua11, BvBGOG03, MJ97, Chi02, Muk15]. Multivariable [AS02, SMHBR06]. Multivariate [Ahm08, AS04, Ano00a, Ano06c, Fri06, JMS^{+}09, NO06, Zie95a, Zie96a, Zie00a, Ano99, HS94, KN95, KN99, MJ97, McF01, Zie01b]. multivariée [Laf81]. Multrus [Ano93b]. Muncher [Mon97]. MVS [Rai95, SAS93-46, SAS94x, SAS95e]. MYSQL [Ano06e, Ano04b].

N [Ahm08, Den12, Hor09a, Lu08, McF01, Nie08, Zie00a, Zie01b]. N. [Ste11]. Naik [Ano99, McF01, Zie96a, Zie00a, Zie01b]. Namur [Le 94]. Nashville [Ano98, Ona90]. National [Ing81]. Native [Wei12]. NC [Ano99, BR09, Bie08, Bos07, Dob09, McF01, ACM96, Nor89]. NC/New [Ano99]. nearest [Whi94]. neighbour [Whi94]. Neil [Cha04a, Rut12]. NESTED [SH80c]. NetCONNECT [Ano00b]. Network [Ter81, WB85]. Nevada [SUG85]. next [Hoy96]. Nicholas [Hec11, Liu17]. ninth [SUG84]. NLMIXED [MPT06, Pat02]. NM [Tru88]. No [BR09, Bie08, Bos07, Chi02, Dob09, Jon07, Kel10, Lum08, Nie08, Rob08, Woo11, oCPA96, CL90, GL76, GL79]. Nolen [Oli17]. Non [LS02, Pat02, RHWP24, PYC93, WWM^{+}24]. Non-independence [LS02]. Non-Linear [Pat02, PYC93, WWM^{+}24]. Non-Programmers [RHWP24]. noncentrality [Har81]. Nonlinear [VDT14, Len15]. Nonparametric [Ami87, DS98b, Kar07, SS98, ZK12]. nonwindowing [SAS93-51]. Nordic [Bjø01]. Norm [Ano06c]. normally [HP92]. North [Nor89, Sch06b, CHM95]. nos [GL76]. Notes [SAS91d, SAS91e, Ste99, Car93, LJ95, LW95, Mil92]. Novel [WWM^{+}24]. November [Bir94]. novice [Rad94, WB85]. NParCov3 [ZK12]. NT [Har93, SAS93j, SAS94-34]. Numbers [AS04]. Numerics [Ano00b]. NV [Tru88]. NWA [Fri91b].

O [Rob07]. O. [Lou06]. Object

[Zie99b, SAS95s]. Object-Based [Zie99b]. objects [SAS93-50]. observations [PG05]. obvious [Sta94]. October [ACM96, Bjø01, SUG87, SAS88a]. ODBC [Ano94b, Ano95c, SAS93t]. odds [XB10]. Offen [Jon07, Chi06, Dur06]. Olaf [Ric08]. Olga [Sha22]. Oliver [Bos07]. Omni [DH94]. One [Cha04a, Cha04b, PKC^{+}20, Spe04]. One-Day [Cha04a, Cha04b, Spe04]. One-Factor [PKC^{+}20]. Online [Din06, SAS91d, SAS91e]. open [SAS93-29]. OpenLinux [Ano00b]. OpenVMS [Har93, SAS93a, SAS94-29, SAS94-34]. Operating [Gho10, Hal08, Kar08, Len10, BR09, SAS84a, SAS86, SAS88t, SAS88z, Saf83]. operation [Ast94b, Kar87a, Kar87b]. Operational [O’B13]. opinion [GL76]. Optimization [CZZ13]. Optimum [AC08, Hor09a, Lu08, Nie08, Puk07]. ORACLE [SAS93u]. orders [Opd13]. Ordinal [MH05, NMvdH14]. organizations [oCPA96]. Organize [LS98]. Oriented [ACM96, Lac99, Ste99]. Orlando [Ano95a, SAS88y]. Orleans [SUG83, SUG91]. O’Rourke [Ano06c]. Orthogonal [FS13]. OS/2 [Har93, SAS90o, SAS93v, SAS93w, SAS93k, SAS94i, SAS94-31, SAS94-34, SAS95h, SAS95-29]. Oslo [Bjø01]. outcomes [MPT06]. Outlier [Law08]. Output [Ano08b, Bar83, LS98, Sun98, Zie89, SG80, SH80c, SH80a, SH80b, Sea80, BI07, Pra08]. Outsourcing [Sha06]. Overdispersed [KM16]. own [Lan82]. Oxford [Nie08].

P [And86, Ano06d, Ano11b, Mar88, Mor02, Zie89, Zie92a, Zie98a, Til96, SAS88r, SAS88s, SAS88t, SAS88u, SAS88v, SAS88w, SAS96c]. P-163 [SAS88r]. P-173 [SAS88s]. P-174 [SAS88t]. P-175 [SAS88u]. P-177 [SAS88v]. P-178 [SAS88w]. P-269 [SAS96c]. Pacific [SUG87]. Package [DS98b, SED14, SS98, XB10]. Packages [Bar83, Fri91b, Gre81, Mil92, Pel95]. PAD [GG82]. pages [Ano99, BR09, Bie08, Bos07, Chi02, Dob09, Jon07, Kel10, Kha20, Lum08, Nie08, RHWP24, Rob08, Woo11, Zha19]. Painless [Gil96, Gil99]. Panik [And10, Web12]. paperback [Nie08, Sha20, Til96]. papers [CM82]. para [Mar83, dV84]. Parallel [Ano00b, Che92]. Parametric [PKC^{+}20]. Paris [Cou01, SAS88a, Van97]. Park [ACM96]. parser [RMM18]. Part [Fra79]. pass [SAS93t, SAS94q, SAS94s, Zie96f]. pass-through [SAS93t, SAS94q, SAS94s]. Patricia [Sin08]. Pattern [BMGT15]. Pattern-Mixture [BMGT15]. Paul [Ano06a, Neu06, Tho92, Til96, Zie97g, Zie00b]. PC [Dri95, EN88, Fri91b, Gri93, SAS93x, SAS94r, SAS95n]. PCR [NTGD18]. PCs [Mil92]. Peace [Rah24]. Pearson [Pan09]. Percentage [Pan09]. perform [BBV17]. Performance [XM14, Mer83, Mer84, Ter81]. permutation [Opd13]. Perpignan [CM11]. Perret [Sea11]. Perrett [Gen11]. Personal [Ano00b, SAS88n, AK85, Ano88, CA85, Eas87, LS87, SAS85j, SAS85b, SAS85c, SAS88i, WB85]. Perualila [Oli17]. Peter [Chi02, Zie99d]. PH-Clinical [SAS93-44]. Pharmaceutical [Bra08, DCSD07, Bie08, Bun07, Che08, Sta07c]. Phase [SED14]. Pilots [KCM08]. Pisa [Lev98]. Place [DH94]. Planning [BN93a, BN93b, BN95b, KCM08, SAS92c, Uni92, Uni93, Ano84a, BN95a, SBJ85, SM82]. Plant [Tru88]. platform [SAS94v]. platforms [Ano94a]. Plaza [All86]. plot [GGS23]. Plots [Fri06, Bir94]. PLUS [Utl05, Pri05, HH04]. pocket [Ast93]. point [LW95]. point-and-click [LW95]. Points [Pan09]. Policy [Dam09, Dri95]. polynomials [SMHBR06]. Polytomous [OC15]. Population [LVK^{+}11, CS93]. portfolio [SAS94-35]. Possible [Law08]. pour [Laf81, Ser93d]. Power [JMS^{+}09, Laf02, Tru88, SAS88r]. POWERLIB [JMS^{+}09]. pp [Ano93b, Ano93a, Ano06b, Kha15, Len97, Oli20, Rah24, Sch06b, Sha20, Sha22, Sta21]. Practical [All95, Ano11b, Bra08, Bun07, Che08, Chi06, DCSD07, Dur06, Lu19, Sta07c, Bie08, Jon07, MB92, Yor94]. Practice [Lac99, Ste99, Hil93, WHG88]. practicing [Hil93]. Practitioners [GL14]. praxisbezogene [Ano93b, MB92]. Precedence [FTP02]. prediction [GSD04]. predictive [Sha22]. preliminary [Sal79a]. Preparation [Nas08, Sta07a, Wie08, Yor94]. preparing [SAS92c, Uni92, Uni93]. present [SAS94j]. Press [Ano06b, Bie08, Jon07, Kha20, Laz24, Nie08, RHWP24, Rah24, Sha20, Sha22, Sta21, Liu17]. Prevalence [Ami87]. Price [BR09, Bie08, Bos07, Chi02, Dob09, Jon07, Kel10, Lum08, Nie08, Rob08, Woo11]. primer [DS95, DS98a, Hor09b, Kar09]. PRIMOS [SAS90c]. principal [Hoh85].

Principles [Zie99d]. Pro [Ano00b, GL76]. Probabilities [Wan19]. probability [DCT89a, Dud86, DCT89b, Ros10, Ros14]. Problem [SAS88k]. problemorientierte [Hac84]. Problems [UE90, CS89]. PROC [MPT06, Ann05, Ke11b, Mon97, SB96]. PROC R [Wei12]. Procedure [KGHS84, Pat02, Dru79, KHSG83, OK80, OGR89, SAS89a, SAS90f, SAS93t, SAS93-48, SAS93-49, SAS93-51, SAS94q, SAS94s, SAS95i, Sal80]. Procedures [SAS88p, UE90, WJ19, FL81, HP92, SAS81b, SAS85c, SAS89b, SAS90g, SAS90h, SAS90j, SAS91i, SM82]. Proceedings [ACM96, All86, Ano96a, Ano96b, DH94, Ona90, SUG82, SUG83, SUG84, SUG87, SUG89, SUG91, ACM92, Bjø01, Gra96, Hut92, Ogl93, SUG85, SAS88a, SAS88y, SAS91h, TD85, Tie88, Tie89, AV08, CF99, Cou01, Cou03, CM11, HP02, Lev98, MS12, NF07, Pal00, PS09, Yah11, Yi06, CS96, Gia04, HS05, Le 94, Myc95, Van97]. processes [Sha22]. processing [oCPA96, Kal86, Mir93, SAS90e, SAS90k, SAS92e]. produccion [dV84]. product [And89, SAS82c, SAS95-31]. productCd-1118762614.html [Kha15]. Production [Pop97]. Products [Ano00b, SAS90o, SAS92c, Uni92, Uni93]. Professional [AS91, Ast92, Ast93, AS96]. professionals [RHWP24]. Program [BMGT15, NZ88, Sou94, Zim07, GG82, HP06, Hoh85, JLC91, SAS83b, Bjø01]. Program-Driven [Zim07]. Programmer [Cla03, Gur08, Hol08, San03, Sta07b, Ano94b, Ano95c, AA90, Ast93, Dob09, HR79, SAS81c, SAS85d]. Programmers [RHWP24]. Programming [And86, Ano06d, Bai19, Cha04a, Cha04b, Hud89, Kha20, LaL92, LH07, LM86, Mor02, Ng12, Oli20, Rut12, Sea12, Spe08, Wei12, Zie96f, Zie98a, Zie00c, Zim07, AS91, AS96, CS85, CS87, CS91, CPI95, CS97, DiI91, DCT89a, Dud86, DCT89b, Hil93, Hsu86, Nat93, SAS90k, Spe93, Spe04, Til96, MT14, Zie99b]. Programs [Ami87, Bas03, SB01, VGDM^{+}18, CS93, Fan94, Sam12, SMHBR06]. Project [SAS93g, And89, Ano84a, SAS93-42, SAS93-40, SAS93-43]. projects [And89]. projects-at-a-glance [And89]. PROJMAN [SAS93-41]. proportional [XB10]. protect [Sam12]. Pseudorandom [AS04]. Pty [Ano00b]. Publication [MYM21, RMM18]. Publication-Quality [MYM21]. Publications [Til96]. Publishing [BMDB98, McF01].

QC [BSW89, SAS89h, SAS94-28, SAS95z, SAS95y]. Qingzhao [Laz24, Sha22]. QLS [KS10]. Quality [Kas98a, Kas98b, MYM21, Zie97e, SAS95y]. Quantal [AR91]. Quantitative [All06, AR91, Hoh85, Sau04a]. Quasi [KS10]. Quasi-Least [KS10]. Query [SAS95j, SAS93-52, SAS94u]. Quick [Bzi11, CSS95, DH96, Spi96, Zie96e, Zie97c]. Quick-Start [Bzi11].

R [And86, Ano09c, Bun07, Den12, GL14, Hec11, Hor09a, Jen09, Laz24, Liu12, Liu17, Lou06, Lu08, Lu19, Mar88, Mas88, McF01, Nie08, Rah24, Ree09, Rob07, Sha22, Ste11, Tur12, Utl05, Ven01, Zie89, Zie92a, Zie92b, Zie96a, Zie96f, Zie97d, Zie99d, BVFB19, BMGT15, Cha08, Dam09, FMF12, FCY20, GGS23, HH04, HH15, KH10, Mue09, NMvdH14, Pri05, Rec10b, SMHBR06, Sha22, SC07, Sho18, Tae14, TR14, TPAM07, Wei12, Ano08a, Bra15, Hu09, Kel10, Kha15, Lum08, O’B08, Oli17, Rah24, Sha22, Sta21, Ano09c, Mai12, Oom10, Ree09, Rob09, Sau11, Voe09]. R-geepack [NMvdH14]. R-multgee [NMvdH14]. R-repolr [NMvdH14]. R-Squared [Rec10b]. Radisson [All86, Bjø01]. Raleigh [Nor89]. Ralph [Bie08, Che08, Sta07c]. Ramon [Bos07, Hei90, Moo03, Zie99c, Zie01c]. Ran [Sha20]. Randall [Chi02, Puk07]. Random [VDT14, SH80c]. Randomization [Gre17, MJ97, NZ88, WWG09, ZK12]. Randomization-Based [ZK12]. Randomized [Gre17, TPAM07]. Rank [Fri06]. rapid [KB95]. Rasch [OC15, Wei82]. Raton [Ano06b, Kel10, Liu17, Lum08, Oli20, RHWP24, Rah24, Rob08, Sha22, Sta21, Woo11]. Ravindra [Ano99, Zie00a, Zie01b]. real [NTGD18, SAS94j, SHO^{+}96]. real-time [NTGD18]. real-world [SAS94j]. Receiver [Gho10, Hal08, Kar08, Len10, BR09]. Reduced [Fri06]. Reduced-Rank [Fri06]. Reduction [Zie01b, SHD88]. Redundancy [LV16, LV17]. Reference [SAS88e, SAS88f, Ano94b, Ano95c, Ast93, BSW89, SAS89c, SAS89e, SAS89f, SAS89a, SAS90m, SAS90n, SAS90p, SAS90r, SAS90f, SAS90t, SAS90u, SAS90i, SAS90l, SAS91k, SAS91l, SAS91q, SAS91w, SAS93q, SAS93r, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS93z, SAS93-29, SAS93-30, SAS93-44, SAS93-27, SAS93-31, SAS94t, SAS94w, SAS94x, SAS94v, SAS94o, SAS94n, SAS95l, SAS95m, SAS95n, SAS95r, SAS95i, SAS95j, SAS95z]. references [SAS95v]. referencing [SAS92c, Uni92, Uni93]. Reflexive [Che92]. Registered [KGHS84]. Regression [AS02, Ami87, Ano06a, BKL05, Dha10, Len09, MH05, MF02, MF03, Neu06, SA87, TR14, UE90, ZMG07, Zie96d, Zie00b, FL91, Gri93, HP06, MPV12, Par06, SAS95f, Sal81a, SMHBR06, GI94, Web12, Mas88, Zie92b, Zie01c, And10, AC04, Der10, Gun03]. relational [SAS94t]. Relations [FTP02]. Relationships [Pat02]. Release [Har93, SAS93e, SAS94-31, Ano94b, Ano95c, SAS82e, SAS88j, SAS88m, SAS88n, SAS88o, SAS88p, SAS88r, SAS88t, SAS88u, SAS88v, SAS88w, SAS88l, SAS88q, SAS88b, SAS88h, SAS88z, SAS89d, SAS89e, SAS90x, SAS91a, SAS91x, SAS91m, SAS91p, SAS91o, SAS91r, SAS91v, SAS91s, SAS91n, SAS92j, SAS93b, SAS93a, SAS93f, SAS93-28, SAS93o, SAS93-45, SAS93-46, SAS93-48, SAS93-49, SAS93-50, SAS93-51, SAS93-52, SAS94h, SAS94q, SAS94v, SAS94-30, SAS94-32, SAS94-34, SAS95d, SAS95g, SAS95t, SAS95w, SAS95-27, SAS95-28, SAS95-29, SAS95-30, SAS96c, Sto82]. releases [LJ95, SAS93-44]. Remark [Nov03]. Remote [SR07]. Reno [SUG85]. Repeated [KAK05]. REPLAY [Ano84d]. REPLAY-CICS [Ano84d]. Replicating [Dam09]. repolr [NMvdH14]. Report [Dob06, Nor89, SAS88r, SAS88s, SAS88t, SAS88u, SAS88v, SAS88w, Ano94b, Ano95c, Bir94, CL90, SAS94j, SAS94q, SAS94v, SAS94l, SAS94-30, SAS94-31, SAS94-32, SAS95s, SAS95-28, SAS95-29, SAS95-30, SAS96c, Mon97, SAS90f, SAS93-51, SAS95i]. report-writing [SAS94j]. Reporting [SAS94j, SAS94u, Ter81]. reports [oCPA96, GL79, Sou94, Uni94a, Uni94b, Yor94]. Republica [dV84]. Resampling [KAK05]. Research [ACM96, Ano03a, Ano12, Bar83, DL04, O’B13, Wal02, Zie96c, Zie99a, Zie99e, Gre81, SAS94e, SBJ85, Liu04, Zie98b]. Researchers [Mor02, RHWP24, Spe93, Til96, All06, Sau04a]. Resource [Din06]. Resources [Wil89]. Response [AR91, Ami87, MH05, NZ88]. Responses [GHR18]. Results [BN93a, BN93b, BN95b, Zie96e, BN95a, CSS95]. Review [Ahm08, Ale91, All06, And10, AC04, ACL04, And86, Ano93b, Ano93a, Ano99, Ano03b, Ano03a, Ano04a, Ano06a, Ano06b, Ano08a, Ano08b, Ano09a, Ano09b, Ano09c, Ano09d, Ano11a, Ano11b, Ano12, BR09, Bas03, Bie08, Bra15, Bun07, Bzi11, Cab04, Cha04a, Che08, Chi06, Chi02, Cla03, Cle03a, Cow04a, Cow04b, Den12, Der10, Die20, Dob09, DL04, Dow88, Dra11, Dur06, Gen11, Gho10, Gie11, GL14, Grö10, Gun03, Gup02a, Gup02b, Hac84, Har08, Hec90, Hec96, Hec11, Hei90, Hol08, Hor09a, Hor09b, Hu09, JS13, Jen08, Jen09, Jon07, Kar08, Kar09, Kar07, Kas98a, Ke11a, Kha15, Kha18, Kha20, Khu13, LaL92, LaL09, Lac99, Laz24, Len09, Len97, Len10, Len15, Lew07, Liu04, Liu12, Liu17, Lou06, Lu08, Lu19, Lum08, Mai12]. Review [Mar88, Mas88, McF01, MT14, Moo03, Mor02, Muk15, Nas08, Neu06, Ng12, Nie08, Nie11, O’B08, O’B13, Oli10, Oli20, Oli17, Oom10, Pin10, Puk07, RHWP24, Rah24, Ree09, Ric13, Ric08, Rob07, Rob08, Rob09, Ros15, Rut12, San03, Sau04a, Sau04b, Sau11, Sch06b, Sch06c, Sea11, Sea12, Sha20, Sho90, Sin08, Spa06, Sta21, Sta07d, Sta07a, Sta07c, Sta07b, Ste11, Ste06, Ste99, Tho92, Til96, Tur12, Utl05, Van04, Ven01, Voe09, Wan19, Web12, Wil09, Zha19, Zie92a, Zie92b, Zie93, Zie95b, Zie95c, Zie95a, Zie96a, Zie96g, Zie96c, Zie96b, Zie96d, Zie96e, Zie96f, Zie97e, Zie97d, Zie97g, Zie97a, Zie97c, Zie97b, Zie97f, Zie98a, Zie98b, Zie99b, Zie99d, Zie99c, Zie99a, Zie99e, Zie00c, Zie00b, Zie00a, Zie01b, Zie01a, Zie01c]. Review [Zie02, AC08, Fra81, Sha06, Sha22, War05]. reviewed [Sha09, Sha15]. Reviews [Ano06c, Ano06d, Bos07, Kel10, Sha15, Woo11, Zie89]. Richard [Utl05]. Richter [Kar07]. Rick [Sea12]. Ridge [Tru88]. risk [GSD04]. risks [BVFB19]. RMF [Mer83]. roadmaps [BK94, Zie95c]. Robert [Bzi11, Mai12, Sau11, Voe09]. ROM [War05, Chi02]. Ron [Cla03, Dob09, Hol08, Khu13, San03, Sta07b]. Ronald [Ano06d, Zie98a]. Rosenberg [SUG91]. RTEL [Ano00b]. RTERM [EK86]. Rudolf [Ale91, Moo03, Zie01c]. Russell [Bos07, Chi02, Liu12, Zha19]. Rutledge [Bzi11].

S [Ano06d, Ano06b, Ano09a, Kar09, Len97, Lew07, Oli10, Pin10, Ric13, Rob08, Sch06c, Ste06, Utl05, Zie89, Zie95c, Zie01a, Zie02, HH04, Pri05]. S-PLUS [Utl05, Pri05, HH04]. S. [Ano09b]. safety [CHM95]. Sage [Til96]. Sample [KAK05, SAS91d, SAS91e]. Sampling [dSM15, Opd13, Tam94]. Samuel [Zie99b].

San [Ano97, Cou03, SUG82, SUG89]. Sandra [Hei90, Zie99c]. Santa [Pal00]. SAS [Ale91, AV08, AC04, ACL04, And86, Ano99, Ano00a, Ano00b, Ano08a, Ano08b, Ano09c, Ano11a, Ano11b, Ano12, BR09, Bas03, Bjø01, Bos07, Bra15, Bun07, Cab04, Chi02, Cla03, CS96, Cou01, Cou03, CM11, DH94, Dob09, DL04, Gen11, Gia04, GL14, Gun03, Gup02a, Gup02b, Hac84, HS05, Hec90, HP02, Hol08, Hor09b, Hu09, Jon07, Kar09, Kar07, Kas98a, Kel10, Kha15, Kha20, Lac99, Le 94, Liu12, Lu19, Lum08, Mai12, McF01, MS12, Myc95, Nie11, NF07, O’B08, Ogl93, Oli17, Pal00, PS09, RHWP24, Rah24, Ree09, Ric08, Rob09, Ros15, Sau04a, Sau04b, Sau11, Sch06b, Sch06c, Sea11, Sea12, Sha20, Sho90, Sin08, Sta21, Ste06, Ste99, Tho92, Ven01, Voe09, Wan19, War05, Wei94, Yah11, Yi06, Zie89]. SAS [Zie95b, Zie95a, Zie96g, Zie96b, Zie96d, Zie96e, Zie97g, Zie97f, Zie98a, Zie99b, Zie99d, Zie99a, Zie99e, Bie08, Jon07, oCPA96, AS02, All82b, All82a, AK85, All95, AR91, And89, Ann05, Ano83, Ano84a, Ano84d, Ano84c, Ano84b, Ano88, Ano94a, Ano94b, Ano95c, Ano96a, Ano96b, Ano97, Ano98, Ano04b, Ano06d, Ano06e, AA90, AS91, Ast92, Ast93, Ast94a, Ast94b, AS96, BVFB19, BN93a, BN93b, BN95a, BN95b, BMDB98, Bai19, BSW89, Bar83, BS95, Bar76, Bar79, Bas84, Bee05, BH14, BBV17, Bir94, BI07, Bra08, BvBGOG03, BD86, BD03, Bro89, BMGT15, BK94, CS93, Can93, Car93, Cha04b, Cha08, Che88, Che92, CL90, Cle03b, CS85, CS87, CS91, CPI95, CS97, Cot06, Cou80, CH81, CP94, Cow04b, CA85, CB85, Dam09, DS95, DS98a]. SAS [DE02, DE05, DE08, Dha10, DiI91, DH96, Dix07, DCSD07, Dob06, Dri95, Dru79, DS98b, DCT89a, Dud86, DCT89b, EK86, Eas87, Ell95, Ell00, EM09, EM10, EN88, ED96, FMT02, Fan94, FZZZ09, Fer03, FCY20, FS13, Fow93, Fra14, Fra79, FL81, FLS86, FL91, Fri91a, Fri06, Fri91b, GHR18, GG82, Gil96, Gil99, Gil12, GGS23, GL76, GL79, Gos14, Göt92, GSD04, Gra96, GI94, Gre81, Gre17, Gri93, Gur08, Hal08, HP06, Har93, Har81, Hat94, HS94, HH04, Hel77, Hel78, HR79, Hel83, Hel85, Her90, Hil93, Hoh85, HP92, Hoy96, Hut92, Ing81, Jaf89, Jaf94, Jaf96, JLC91, JMS^{+}09, Joy85, Jur15, KCM08, KM16, Kal86, Kar87a, Kar87b, Kas98b, Ke11b, KN95, KN99, KS10, KH10, KHSG83, KGHS84, KB95, KW85, KK96, LaP81]. SAS [LVK^{+}11, Laf02, Laf81, Lan81, Lan82, LJ95, LW95, LS02, Law03, Law08, LSC09, Law10, LH07, Len15, LF83, LF87, LS98, LLS85, LFS91, LSF02, Lit06, LG11, Liu17, LM86, LV16, LV17, LS87, MK08, Mar83, MJ97, MPT06, MH05, Mer83, Mer84, Mil93, Mir93, Mir95, Mon97, Mue09, MF02, MF03, MB92, MYM21, Mwi11, Mwi12, Nat93, NTGD18, NZ88, NO06, NB18, NMvdH14, Nov03, OC15, OK80, Ona90, OGR89, Pan09, Pat02, PYC93, Pel95, Per10, PKC^{+}20, Pie87a, Pie87b, Pop97, Pra08, PG05, Pri05, QDG17, Rad94, Rai95, Ray82, Rec10a, Rec10b, Rei80, RMM18, Rod11, Ros83, SUG82, SUG83, SUG84, SUG85, SUG87, SUG89, SUG91, Ser93b, Ser93a, Ser93c, Ser93d, Saf83, Sal78a, Sal78b, Sal79a, Sal79b, Sal81b, Sal81a, Sal82, Sal88, SMHBR06, SB96, Sch06a]. SAS [SED14, SL87, SG80, SH80c, SH80a, SH80b, Sea80, Sel11, Sha06, Sha09, Sha15, Shi96, SC07, Sho18, SH09, Sil96, SS98, SR07, SB01, Spe93, Spe08, Spe04, Spi96, SM82, Sta86, Sta94, Ste83, Ste86, SLMN86, SDK95, Sto82, SMCW18, SA87, Sun98, Tae14, Tam94, TR14, Tie88, Tie89, TPAM07, Til94, Tru88, Uni92, Uni93, UM14, UE90, dV84, VDT14, VGDM^{+}18, Wal02, WWG09, WWM^{+}24, Wei12, Wei82, WB85, WHG88, Whi94, Wie06, Wie08, Wil89, Wit92, Wul07, XB10, XM14, Yor94, Yua11, Zel02, ZCI^{+}16, Zie96c, Zie96d, Zie97a, Zie97f, Zim07, ZK12, dSM15, Ahm08, And10, Ano06b, Ano09a, Ano09b, Cle03a, Dra11, Gho10, Gie11, Hor09a, Kar08, Laz24, Len97, Len10, Lew07, Lu08, Nas08, Neu06, Ng12, Nie08, O’B13, Oli10, Oli20, Pin10]. SAS [Puk07, Ric13, Rob08, Sch06b, Sha22, Utl05, Web12, Wil09, Woo11, Zie97c, Zie97b, Zie02, All06, AC08, Ano93b, Ano93a, Ano08b, Ano09d, Cha04a, Cow04a, Den12, Der10, Die20, Dow88, Grö10, Har08, Hec11, Hei90, JS13, Ke11a, Kha18, LaL92, Len09, Mar88, Mas88, Mor02, Muk15, Oom10, Rob07, Spa06, Ste11, Til96, Tur12, Zha19, Zie92a, Zie92b, Zie93, Zie95c, Zie96f, Khu13]. SAS-AF [SAS88b]. SAS-Based [DS98b, FS13, SS98]. SAS-C [SAS88c, SAS88d, SAS88e, SAS88f, SAS88g]. SAS-FSP [SAS88h]. SAS-GENMOD [NMvdH14]. SAS-GRAPH [SAS88i, SAS88j, SAS88w]. SAS-IML [SB01, SAS88l]. SAS-Oriented [Lac99, Ste99]. SAS-PAD [GG82]. SAS-STAT [SAS88q]. SAS-System [MB92]. SAS-World [SAS88a]. SAS/ [Spi96]. SAS/ACCESS [Ano04b, Ano06e, SAS90m, SAS90n, SAS90o, SAS90p, SAS91k, SAS91l, SAS91m, SAS92h, SAS93q, SAS93r, SAS93t, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS94b, SAS94q, SAS94r, SAS94s, SAS94t, SAS95l, SAS95m, SAS95n, SAS96c, Ser93d]. SAS/AF [SAS85g, SAS89c, SAS91n, SAS93z, SAS93-27, SAS95o, SAS95p]. SAS/ASSIST [SAS90q, SAS91c, SAS92a, SAS94u, SAS95q, SAS96b]. SAS/C [SAS89d, SAS89e, SAS91p, SAS91o, SAS93b, SAS94v]. SAS/CALC [SAS91b, SAS91q]. SAS/CONNECT [SAS90r, SAS93-28, SAS94w]. SAS/CPE [SAS93-29, SAS94x]. SAS/DB2 [Bro89]. SAS/EIS [SAS93d, SAS93-30, SAS93-31, SAS93-50, SAS94a, SAS95r, SAS95s]. SAS/ENGLISH [SAS93-32, SAS93-34, SAS93-33]. SAS/ETS [All82b, SAS84c, SAS85h, SAS88x, SAS93-35, SAS93-36, SAS94y, SAS95t, SAS95u]. SAS/FSP [SAS82c, SAS89f, SAS91n]. SAS/GIS [SAS95v]. SAS/Graph [Zie96g, Zie96e, Ano84c, CH81, Joy85, SAS81d, And89, CSS95, Eas87, Lan81, Lan82, Mir95, SAS82d, SAS82e, SAS83c, SAS90t, SAS90s, SAS91u, SAS91r, SAS91s, SAS91t, SAS93-37, SAS93-38, SAS94d, SAS94z, dV84]. SAS/IML [Gen11, Sea11, Sea12, AS02, HP06, KCM08, NZ88, NO06, Pan09, Per10, Sal88, ZK12, CA85, CB85, JMS^{+}09, SAS85i, SAS90u, SAS91v, SAS95w]. SAS/IMS [Sto82]. SAS/IMS-DL/I [Sto82]. SAS/INSIGHT [LJ95, SAS93-39, SAS94-27, SAS95x]. SAS/LAB [SAS92i]. SAS/OR [Ano84d, SAS89g, SAS93-42, SAS93-40, SAS93-41, SAS93-43]. SAS/PH [SAS93-44]. SAS/PH-Clinical [SAS93-44]. SAS/QC [BSW89, SAS89h, SAS94-28, SAS95z, SAS95y]. SAS/RTERM [EK86]. SAS/SHARE [SAS93-45, SAS93-46, SAS94-29, SAS94-30, SAS94-31, SAS94-32, SAS95-27, SAS95-28, SAS95-29, SAS95-30]. SAS/SPECTRAVIEW [SAS94c]. SAS/STAT [LS87, SAS85j, SAS90v, SAS92j, SAS93-47, SAS93-48, SAS93-49, SAS94-36, SAS94-33, SAS96d]. SAS/Tell [SAS76]. SAS/Tell-a-graf [SAS76]. SAS/TOOLKIT [SAS91w]. SAS/TRADER [SAS95-31]. SAS/TUTOR [Har93, SAS94-34]. SAS’97 [Van97]. SAS’98 [Lev98]. SAS’99 [CF99]. SASEPA [FTP02]. SASÒ [Sau04a]. SAS(R) [Bie08, Ano03a, Ano06c, Ano06d, Bzi11, Che08, Chi06, Dur06, Hec96, Jen08, Jen09, LaL09, San03, Sta07b, Van04, Zie96a, Zie96c, Zie97e, Zie97a, Zie98b, Zie00b, Zie00a, Zie01b, Zie01a, Ami87, Opd13, MT14, OH13, SWH^{+}11, Sta07a, Sta07d, Sta07c, BR09, Ano06a, Bos07, Lou06, Ano03b, Ano04a, Liu04, Moo03, Rut12, Zie97d, Zie99c, Zie00c, Zie01c]. SAS^{T M} [WJ19]. SASweave [LH07]. Scandinavia [Bjø01]. Schabenberger [Bos07, Lou06, Rob07]. Scheduling [FTP02]. Schlotzhauer [Hei90, Zie99c]. Schoffer [Ric08]. Schreier [Ke11b].

Science [Kha20, RHWP24, Rah24, All86]. Sciences [All86, LaP81, Laf81, Wal90]. scientists [Spe93, Mor02, Til96]. Scope [RHWP24]. Scott [Kar07]. Screen [SAS91j, SAS92f, SAS94n, Sta94, SAS82c, SAS90l, SAS90x, SAS91n]. seal [CS93]. Sean [All06, Ano03b, Sau04a]. Searle [Zie89]. Second [Bra15, Mai12, McF01, Myc95, Ros15, Ano99, DH94, WHG88]. secrets [AS91, AS96]. section [Dru79]. sections [GL76]. Security [Ano00b]. Select [Gos14]. selected [CM82]. Selection [AS02, SH09, SAS94-35]. Selective [Sha06]. selectivity [Mil93]. Semi [VGDM^{+}18]. Semi-Continuous [VGDM^{+}18]. Seoul [Yi06]. sep [War05]. September [CF99, CS96, CM11, DH94, HS05, HP02, Le 94, Lev98, MS12, Myc95, Van97, Yah11]. Sequential [BVFB19, JMD08]. Series [Oli17, Ven01, BD86, BD03, Dru79, QDG17, SAS95u, Sal79a, Ano04a, Dow88]. Server [SAS93y, SAS95a, SAS95b]. service [oCPA96]. Services [Ano00b, SAS93p]. session [Ser93b, Ser93a, Ser93c, Ser93d]. SESUG [DH94]. sets [Lan81, Lan82, SAS91s, SAS95c, Sal79b]. Seugi [Tie88, Tie89]. Seventeenth [All86, Hut92]. Seventh [SUG82]. several [Ano94a, TPAM07]. SGML [BS95, SHO^{+}96]. SHARE [SAS93-45, SAS93-46, SAS94-29, SAS94-30, SAS94-31, SAS94-32, SAS95-27, SAS95-28, SAS95-29, SAS95-30]. Sharpening [Cot06]. SHAZAM [Sil96, WHG88]. Shipp [Zie96e]. Shkedy [Oli17]. Short [Dra11, Khu13, Liu12, Mai12, O’B08, O’B13, Puk07, Ric13, Sta07d, Sta07a, Sta07c, Sta07b, Voe09]. Shostak [Ano12]. Shoukri [Ano08a, Lum08, Sta21]. SIGDOC [ACM96]. sigma [SG80, SH80c, Sea80]. Signs [NO06]. silico [NTGD18]. Simulating [AS04]. simulation [SAS93-35]. Simulations [Nov03]. Simultaneous [FTP02, SB01]. SIR [Gre81]. Sistema [Mar83]. Sivo [All06, Ano03b, Sau04a]. Sixteenth [SUG91]. Size [FS13]. Skew [LS02]. Skew-symmetry [LS02]. Skills [Cot06]. Slaughter [Hor09b]. SMF [Mer83]. SMF/RMF [Mer83]. Smith [Ano06d, And86, Zie98a]. Smoothing [LG11]. Smoothing-Spline-Based [LG11]. Snippet [Dha10, Gos14, Gre17, Jur15, KM16, LSC09, LG11, LV16, OC15, Pan09, PKC^{+}20, Rec10a, Rec10b, TR14, VDT14, WWG09, Wei12]. SNP NLMM [VDT14]. Social [Mor02, LaP81, Spe93, Til96, Wal90]. sociales [Laf81]. SOCR [Din06]. Software [Ano00b, DS98b, Fri06, Fri91b, GL14, JMD08, JMS^{+}09, KCM08, Sha06, SS98, Yua11, ZMG07, Zie95c, Ano94a, BSW89, Bir94, BK94, CSS95, Fra81, GG82, GGS23, HP06, Hoy96, JLC91, KN95, KN99, LJ95, LS05, Mir93, Mir95, MF02, MF03, NTGD18, PG05, SAS89c, SAS89f, SAS89h, SAS90q, SAS90r, SAS90t, SAS90w, SAS90u, SAS90s, SAS91a, SAS91c, SAS91m, SAS91u, SAS91r, SAS91v, SAS91s, SAS91t, SAS91n, SAS91q, SAS91w, SAS92a, SAS92j, SAS92g, SAS92i, SAS93d, SAS93z, SAS93-29, SAS93-28, SAS93-30, SAS93-32, SAS93-34, SAS93-35, SAS93-37, SAS93-38, SAS93-44, SAS93n, SAS93o, SAS93-27, SAS93-33, SAS93-40, SAS93-41, SAS93-31, SAS93-45, SAS93-46, SAS93-47, SAS93-48, SAS93-49, SAS93-50, SAS94-36, SAS94a, SAS94b, SAS94c, SAS94d, SAS94f, SAS94g, SAS94j, SAS94s, SAS94t, SAS94u, SAS94w, SAS94x, SAS94y, SAS94z, SAS94-27, SAS94-28, SAS94o, SAS94p, SAS94-29, SAS94-33, SAS95n, SAS95o, SAS95p, SAS95q, SAS95r, SAS95t]. software [SAS95u, SAS95v, SAS95w, SAS95z, SAS95y, SAS95k, SAS95-27, SAS95-31, SAS96b, SAS96d, Sam12, Spi96, Sta86, WB85, AC04, Ano00a, Gun03, Sea12, Zie96a, Zie96g, Zie96e, Zie99b, Zie99d, Zie00a, Zie01b, Ano99]. Solution [Ano00b, SAS95a]. solutions [Mir93]. solve [PYC93]. Solving [Che92, SAS88k]. Some [Mil92, Fra79, Gre81]. Sonja [Kha15]. Sons [Ano99, Kha15]. Source [SAS88g, RMM18]. South [DH94]. Southeast [DH94]. Space [Sel11]. Spain [AV08, HP02]. Spatial [Gri93, NO06, dSM15, Sha22]. Speaking [Sam12]. special [Fra79]. specification [LLS85]. specifications [JLC91]. Spector [Til96, Mar88, Mor02, Zie92a]. SPECTRAVIEW [SAS94c]. Spencer [Cha04a]. Spline [LG11]. spreadsheet [SAS91b]. Springer [Sha20]. SPSS [Ano09c, GL14, Mai12, Oli10, Ree09, Sau11, Ste06, Voe09, Bar83, Dri95, EN88, Fan94, Fri91b, Gre81, HP92, KW85, Laf81, Law03, Mue09, NMvdH14, Pel95, Sha06, Sha09, Ven01]. SPSS-GENLIN [NMvdH14]. SPSS/PC [Dri95, Fri91b]. SPSS^{T M} [WJ19]. SPSSX [Zie89, Bar83]. SQC [SAS95y]. SQL [Ann05, Ke11b, SAS89a, SAS90p, SAS93t, SAS93y, SAS93-52, SAS94q, SAS94s, SAS95j]. SQL/DS [SAS90p]. Square [LS02]. Squared [Rec10b, Har81]. Squares [KS10]. Sr. [Che08]. Staff [Zie96c, Zie96d, Zie97a, Zie97f]. Stalker [Ano00b]. standard [CL90]. Start [Bzi11, Zie97c, DH96]. started [SAS91c, SAS91b, SAS93d, SAS94b, SAS95e, SAS96b]. STAT [LS87, SAS85j, SAS88q, SAS90v, SAS92j, SAS93-47, SAS93-48, SAS93-49, SAS94-36, SAS94-33, SAS96d]. STATA [OGR89, SMHBR06, Dam09, VGDM^{+}18]. State [Nor89, Sel11, CHM95]. statements [CL90, GL79]. Static [Le 94, AV08, CF99, CS96, Cou01, Cou03, CM11, Gia04, HS05, HP02, Le 94, Lev98, MS12, Myc95, NF07, Pal00, PS09, Van97, Yah11, Yi06]. Statistical [Ano03a, Ano09a, Ano09b, Ano09d, Ano12, Bai19, Bar83, Den12, DE05, DE08, DL04, Fra81, Fri91a, Fri91b, GL14, Gup02a, Gup02b, Har08, Hec11, HH04, HH15, Hei90, Ing81, Kas98b, Ke11a, KH10, KGHS84, KW85, Liu17, Mar83, Mon97, Mwi12, Pin10, Pri05, Ste11, Tae14, Wal02, Wie06, Wil09, Wit92, Zie93, Zie97b, Zie99b, Zie99c, Zie99a, Zie02, DE02, DCT89a, DCT89b, ED96, FMT02, Fan94, KHSG83, LS05, Len97, MK08, Mil92, NTGD18, Pel95, SAS94f, SAS94g, Sou94, SL87, Sha20, Tur12, Wil89, Ano06d, Ano06b, Ano11a, Die20, Kas98a, Kha15, Laz24, Lew07, Ng12, Oli20, Rah24, Rob08, Sea12, Sha22, Spa06, Utl05, Zie97e]. Statistics [Ale91, ACL04, And86, Ano00a, Ano06d, Bra08, Bun07, Che08, Din06, DCSD07, JS13, Kar09, Khu13, Ric13, Sta07c, Ste99, Van04, Wal90, Zie96a, Zie98a, Zie00a, All86, Ano99, Bee05, Bie08, Can93, CS85, CS87, CS91, CS97, DCT89a, Dud86, DCT89b, FMF12, Fra14, Gri93, HS94, KN95, KN99, LF83, LF87, McF01, SAS82b, SAS85f, SAS94-36, SAS94g, Sha22, Ano06c, Ano06c, Kar07, Zie95a, LaL09, Sta07d]. Statistik [Wei94]. Statistische [Ric08, KW85]. Statpak [Fri91b]. Stein [Bie08, Bun07, Che08, Chi06, Dur06, Jon07, Sta07c]. Step [ACL04, Ano06c, Hec96, MT14, OH13, Ros15, Zie95a, Hat94, HS94, SAS96c, ACL04, Van04]. Step-by-Step [ACL04, Ano06c, Hec96, OH13, Ros15, Zie95a, Hat94, HS94, ACL04, Van04]. Stepanski [Ano06c, Ano06c, Zie95a]. Stephen [All06, Ano03b, Sau04a]. Stochastic [JXW^{+}11, Sha22]. Stock [SAS94-35, SAS95-32, Zie97f]. Stokes [Sch06c, Ste06, Zie96b, Zie01a]. Strategic [Zie99d]. strategies [CZZ13, KB95, Sta86]. strategy [How88]. stratigraphic [Hoh85]. Strengthening [Dri95]. Stroup [Bos07, Lou06, Moo03, Rob07, Zha19, Zie97d]. Structural [BKL05, Hec96, OH13, Pat02, Ros15, Che88, Hat94]. Structure [Gos14]. STS [Ano00b]. Student [ACL04, Van04, Dud86]. Students [GL14, RHWP24]. Studies [All06, CBDM07, KGHS84, SED14, KHSG83, Mil93, Sau04a, Tru88, Ano03b].

Studio [Ano08b, BI07, Pra08]. Study [LVK^{+}11, BVFB19, SWH^{+}11]. Stuttgart [Ano93a]. SUDAAN [Dam09]. SUGA [SUG87]. SUGI [Ano96a, Ano96b, Gra96, Hut92, Ogl93, Ona90, SUG82, SUG83, SUG84, SUG85, SAS88y, SAS91h, SUG91]. SUGI-SAS [SUG82]. suite [SWH^{+}11]. summarization [Bir94]. summary [SAS92g]. SUNZ [SUG87]. Super [Ano00b]. Super-Symmetric [Ano00b]. Supercomputing [ACM92]. supplement [Ano04b, Ano06e]. supplemental [Hel77, Rei80, SAS81b, SAS82a]. support [LLS85, SAS91d, SAS91e, SAS93f, SBJ85, SAS91e]. supporting [SAS94k]. supports [Ano94a]. surrogate [Oli17, TPAM07]. Survey [Kha18, CL90, GGS23, GL76, GL79, Zie99e]. Surveys [KAK05]. Survival [All95, Liu04, TR14, Wan19, ZCI^{+}16, Zie98b, GSD04, Laf02, Pel95, Ano11b, Lu19, Zie97g]. Susan [Hor09b]. Svolba [Nas08, Sta07a]. svy freqs [MYM21]. SYBASE [SAS93y]. Symmetric [Ano00b]. symmetry [LS02]. Symposium [All86, AV08, CEM82, CF99, CS96, Cou01, Cou03, CM11, HP02, Le 94, Lev98, MS12, Myc95, NF07, Pal00, PS09, Van97, Yah11, Yi06, Gia04, HS05]. Syntax [SAS93-47, SAS90h, SAS92f, SAS94y]. System [All95, Ano84b, Dow88, FL91, Fri91a, Hec96, Hei90, Jaf96, JLC91, Kas98b, LFS91, Mar88, Mar83, Mas88, Pop97, SAS88z, SAS90q, SAS90w, SAS91c, SAS91f, SAS91g, SAS91x, SAS92b, SAS93g, SAS93p, SAS94e, SAS94k, SAS94-35, SAS95e, SAS96b, TD85, Zie92a, Zie92b, Zie93, Zie95a, Zie97d, Zie97f, Zie99c, Zie01c, Ano88, Ano93b, AA90, BI07, BD86, BK94, Can93, CHM95, Fow93, FLS86, Har93, Hat94, HS94, Her90, Jaf89, Jaf94, JXW^{+}11, LLS85, Mer84, MB92, PYC93, SAS86, SAS87, SAS88r, SAS88t, SAS88u, SAS88v, SAS93a, SAS93e, SAS93-41, SAS94h, SAS94i, SAS94-28, SAS94-34, SAS95a, SAS95b, SAS95d, SAS95f, SAS95g, SAS95h, SAS95u, SAS95-32, Saf83, SL87, SDK95, Til94, MB92, Pra08, SAS95y, Ste06, Ale91, Ano08b, Chi02, Hec90, Sch06c, Sho90, Zie95b, Zie96c, Zie96b, Zie96d, Zie97e, Zie97g]. System [Zie97a, Zie99e, Zie00b, Zie01a, McF01, Kas98a]. systemization [Fan94]. Systems [Ano00b, Che92, KW85, Sil96, SAS84a, SAS88z, SAS89i, SAS93-29].

T [Ano93a, Har81]. Table [KB95]. Table-driven [KB95]. Tables [LS02, Mon97]. TABULATE [Mon97]. Tabulations [MYM21]. Taeger [Kha15]. Tai [Sha20]. Tanya [Zie99b]. tasks [Fan94]. Taylor [Ano06b, Oli20, RHWP24]. TC2 [TD85]. Technical [SAS88r, SAS88s, SAS88t, SAS88u, SAS88v, SAS88w, Zie97f, Ano94b, Ano95c, SAS94q, SAS94v, SAS94-30, SAS94-31, SAS94-32, SAS95s, SAS95-32, SAS95-28, SAS95-29, SAS95-30, SAS96c]. Techniques [Dam09, Liu04, Zie98b, Zie99b, Nat93, SAS94m, SAS95b, Bas03]. Technological [ACM96]. Technologies [Ano00b]. Tell-a-graf [SAS76]. Templars [GB82]. Tennessee [Ona90]. Tenth [SUG85]. terrain [Tru88]. test [MJ97, Sal79a]. Testing [Kha15, SB01, CZZ13, Tae14]. Tests [Ami87, Chi02, PKC^{+}20, Dri95, Opd13, NB18]. Textbook [RHWP24]. their [Har81]. Theophile [Oli17]. Theory [Ahm08, And10, Bra15, Der10, Hu09, O’B08, Web12, Zim07, Cha08, Hil93, Kel10, Len15, Rob09, WHG88]. Third [CS96, Wil09]. thirteenth [SAS88y]. Thomas [Zie96g]. Three [MYM21, Pel95, Sha22]. Three-Way [MYM21]. tie [WB85]. Time [Ano04a, Dow88, GHR18, TR14, Ven01, BD86, BD03, Dru79, NTGD18, QDG17, SAS95u, Sal79a]. time-series [Sal79a]. Time-to-Event [GHR18]. Time-Varying [TR14]. Timm [Ahm08]. tips [SAS90k, SAS94m, SAS95b]. TN [Ano98]. Tobias [Nie08, Hor09a, Lu08, Puk07]. Tomasz [Oli17]. tool [And89, Ano84a]. toolkit [BvBGOG03, SAS91w]. Tools [Bas03]. TotalView [Ano00b]. Toulouse [CEM82]. TPFY [Sou94]. tracer [Tru88]. Tracking [And89]. Tracy [Tru88]. Trademark [KGHS84]. TRADER [SAS95-31]. transactions [oCPA96]. Transitioning [Dam09]. Translating [Bas84]. Transportation [JLC91, SM82]. Transporting [SAS89i, SAS88s]. trawl [Mil93]. Trees [Jen08]. trial [WWM^{+}24]. Trials [Chi06, Dur06, WWG09, BVFB19, Jon07, TPAM07]. Triangle [ACM96]. Trip [Bir94]. Truong [RHWP24]. Tschiersch [Ric08]. TSCSREG [Dru79]. TSO [SAS84a]. TSP [Sil96]. Tumor [Ami87]. Tumor-Prevalence [Ami87]. Tuning [Rai95]. TUTOR [Har93, SAS94-34]. Tutorial [TR14, WWM^{+}24]. Twenty [Ano96a, Ano96b, Gra96]. Twenty-First [Ano96a, Ano96b, Gra96]. Twin [FZZZ09]. Two [Ami87, DS98b, Rec10b, SED14, SS98, Wei82]. Two-Group [DS98b, SS98]. Two-Level [Rec10b]. Two-Phase [SED14]. Tzu [Sha20].

U.S.A [Chi02]. Uehlinger [Hac84]. UK [HS05, Myc95]. Unbalanced [Zie89, SH80a, SH80b, Wit92]. UniLogistic [Dha10]. Univariable [Dha10]. Univariate [Ano06c, Zie95a, HS94, Ahm08]. University [Nie08]. Unix [Ano94a, Har93, SAS88v, SAS88z, SAS90d, SAS91d, SAS91x, SAS93f, SAS93l, SAS93m, SAS94-32, SAS94-34, SAS95-30]. Unreplicated [Law08]. Unstratified [Gre17]. Update [Sun98, Sal78b]. updates [SAS82d, SAS82e]. updating [Bir94]. UpSet [GGS23]. US$34.95 [Sch06b]. US$79.95/£ [Ano06b]. USA [Cou03, Pal00, PS09, ACM96, Ano95a, Ano96b]. usage [SAS89c, SAS89f, SAS89a, SAS89b, SAS90m, SAS90n, SAS90p, SAS90r, SAS90f, SAS90u, SAS91k, SAS91l, SAS91u, SAS91i, SAS91j, SAS91q, SAS91w, SAS93q, SAS93r, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS93z, SAS93-29, SAS93-27, SAS94w, SAS94x, SAS94v, SAS95m, SAS95j, SAS95v, SAS95z]. USD [Kha15]. use [Gre81, WHG88]. USEPA [Ing81]. User [ACM96, CH81, Joy85, SAS88g, SAS88j, SAS88l, SAS88q, SAS88b, SAS88h, SAS93-50, All82b, All82a, Ano94b, Ano95c, Ast92, Bar76, Bar79, CB85, EK86, Hel77, JLC91, Ray82, Rei80, SAS81d, SAS82c, SAS82b, SAS84c, SAS85g, SAS85h, SAS85i, SAS85e, SAS85f, SAS88x, SAS89d, SAS89g, SAS90v, SAS91p, SAS91o, SAS92i, SAS93m, SAS93-34, SAS93-36, SAS93-39, SAS93-42, SAS93-43, SAS95x]. User-Oriented [ACM96]. User-written [SAS93-50]. Users [Ano96a, Ano96b, Ano97, Ano98, DH94, Gil99, Gra96, Hut92, Ogl93, Ona90, SUG82, SUG83, SUG84, SUG85, SUG87, SAS88y, SUG89, SUG91, Spe08, Tie88, Tie89, Fra79, Gil96, Mue09, SAS90w, Whi94, Ano09c, Mai12, Ree09, Rut12, Sau11, Voe09]. Using [AS02, All95, AR91, AC04, ACL04, Ann05, Ano06a, Ano06c, Ano06d, Ano06b, Ano08b, Ano09a, Ano09b, Ano11a, Ano11b, BMDB98, BH14, Bra08, Bra15, Bun07, Cab04, CBDM07, Che08, Chi06, Chi02, Cle03a, Cle03b, DE05, DE08, DCSD07, Dur06, FZZZ09, FCY20, Gun03, Hec96, Hu09, JS13, Jen08, Jen09, Kar09, Kas98a, Kas98b, Ke11b, KS10, LaL09, Laz24, LH07, Lew07, LS98, Muk15, Mwi11, Nas08, Neu06, O’B08, OH13, Pat02, Pop97, Pra08, RHWP24, Rec10a, Rec10b, Ric13, Ros15, SAS88z, SAS90x, SAS93-51, SAS93-52, Sal79b, Sch06a, Sch06b, Sch06c, Sel11, SH09, Sin08, Spa06, Sta07d, Sta07a, Sta07c, Ste06, SA87, TR14, Van04, WWG09, Wan19, War05, WB85, Wie06, Wie08, Wil09, Yua11, Zel02, Zie95a, Zie96c, Zie96b, Zie96d, Zie97e]. Using [Zie97g, Zie97a, Zie97b, Zie97f, Zie99d, Zie99e, Zie00b, Zie01a, Zie02, Bie08, BI07, Can93, Cha08, DE02, Die20, Dri95, ED96, Fer03, FMF12, Fra14, GGS23, Gri93, Gup02a, Gup02b, Hat94, HS94, Ing81, Opd13, JLC91, Jon07, Kel10, LaP81, LJ95, LS02, Len97, Len15, LF83, LF87, MPT06, Mer84, MF02, MF03, Mwi12, NTGD18, Pin10, QDG17, Rob08, Rob09, Ros83, SAS91c, SAS93g, SAS93-38, SAS94-36, SAS94e, SAS94-35, SAS95f, SAS95-32, SAS96b, SMHBR06, SB96, Sha22, Spi96, Sta86, SDK95, TPAM07, WWM^{+}24, WHG88, Wil89, PYC93]. Uso [dV84].

v.5.7 [Ano00b]. Valencia [AV08].

validation [Whi94]. Value [BS95]. Values [SB01]. VARCOMP [SG80]. Variable [AS02]. Variables [LS02]. Variance [NO06, Oli10, Zie89, SG80, SH80c, Sea80]. various [RHWP24]. Varying [TR14]. VBA [Sha20]. Venice [CF99, Yah11]. Verbeke [Lac99, Ste99]. Verfahren [Ric08]. Verlag [Ano93b, Ano93a, Sha20]. Verona [Gia04]. vers [Ser93c]. Version [Ano93a, Ano00b, Göt92, Joy85, SAS85f, SAS88x, SAS90s, SAS94c, SAS94m, SAS95f, SAS95-33, SLMN86, AK85, BSW89, CA85, EK86, Eas87, Kal86, LS87, SAS84c, SAS85j, SAS85b, SAS85i, SAS85c, SAS87, SAS88i, SAS88k, SAS88s, SAS89c, SAS89f, SAS89b, SAS89h, SAS89g, SAS90m, SAS90n, SAS90p, SAS90q, SAS90r, SAS90d, SAS90e, SAS90f, SAS90t, SAS90u, SAS90h, SAS90i, SAS90j, SAS90l, SAS90v, SAS91c, SAS91b, SAS91f, SAS91k, SAS91l, SAS91u, SAS91i, SAS91j, SAS91q, SAS91w, SAS92a, SAS92f, SAS92i, SAS93g, SAS93q, SAS93r, SAS93v, SAS93s, SAS93u, SAS93w, SAS93x, SAS93y, SAS93z, SAS93h, SAS93j, SAS93i, SAS93k, SAS93-29, SAS93l, SAS93m, SAS93-30, SAS93-32, SAS93-34, SAS93-35, SAS93-36, SAS93-39, SAS93-42, SAS93-27, SAS93-33, SAS93-40, SAS93-41, SAS93-31, SAS93-47, SAS93-43, SAS94a, SAS94b, SAS94f, SAS94g, SAS94s, SAS94t, SAS94w, SAS94x, SAS94y, SAS94l, SAS94-28, SAS94o, SAS94n, SAS95c, SAS95l, SAS95m, SAS95o, SAS95p, SAS95r, SAS95u, SAS95j]. version [SAS95v, SAS95x, SAS95z, SAS95-31, SAS96b, Ser93b, Ser93c, Sal79a, Shi96, Sto82]. versions [CM82]. vi [Ano06b]. via [Jur15]. views [Spi96, Ste83]. vii [Jon07, Sch06b]. Ville [Jen08]. Visual [Ano00b]. Visualizing [Sau04b]. VM [Saf83]. VM/CMS [Saf83]. VMS [SAS93-38]. vol [SAS88e, SAS88f]. Volume [Pop97]. Vonesh [Len15]. VS [SAS90a]. VSAM [SAS92e]. VSE [SAS88r, SAS88t, SAS92d, SAS95-27]. VSE-POWER [SAS88r].

W [Ano09d, Bos07, Lou06, Rob07, Zha19, Zie97d]. Walker [Ano12, Ano03a, Zie99a, DL04]. Wallis [MJ97]. Walter [Bos07, Chi06, Dur06, Jon07, Moo03, Zha19, Ric08]. Warehousing [Zie99d]. warning [Whi94]. Washington [ACM92]. Way [MYM21]. Web [Sun98]. weighted [SAS83b]. weights [GGS23]. Welbrock [Zie99d]. Wellington [SUG87]. Wendy [Ano08b]. Westfall [Chi02]. wheat [Bir94]. WHIST [JLC91]. WHIST-MOD [JLC91]. white [Ano00b]. Wicklin [Sea12]. Wiley [Ano99, Kha15]. WileyCDA [Kha15]. WileyTitle [Kha15]. William [Ke11a, O’B13]. Williams [WWG09]. Wilson [Rah24]. Wim [Oli17]. WinBUGS [CBDM07, SR07]. WinBUGSio [SR07]. window [SAS93-52, Har93, SAS95j]. Windows [Mon97, SAS93j, SAS94-34, Shi96, Gil96, Gil99, Har93, SAS93e, SAS93h, SAS93i, SAS94h, SAS94-30, SAS94-34, SAS95g, SAS95-28]. Winfred [Cle03a]. Winston [Cle03a]. Within [Ke11b, Sun98]. Wolfinger [Bos07, Chi02, Lou06, Rob07, Zha19, Zie97d]. WordMage [Ano00b]. Working [Gos14, Til94]. works [Her90, Tho92]. World [SAS88a, Lan81, SAS94j, SHO^{+}96].

Worldwide [JLC91]. Writing [Dob06, SAS94j, SAS94l]. written [SAS93-50]. WWW [Hoy96]. www.wiley.com [Kha15].

x [BR09, Ano06b, Har93, Rob08, Sch06b]. xii [Bos07]. xiii [Kha20]. Xitao [All06, Ano03b, Sau04a]. xiv [Chi02, Kha15]. xix [Sha22]. xv [Oli20, Sha22, Sta21]. xvi [Nie08]. xx [Sha20]. xxi [Oli17, Rah24]. xxii [Kel10]. xxxii [Dob09].

Yaffee [Ven01]. Yerex [Zie89, Zie89]. yield [Bir94]. York [Ano99, Ogl93]. Yosef [Chi02]. Yu [Laz24, Sha22].

Zealand [SUG87]. Zelterman [Sch06b]. Ziv [Oli17].

References

Aronson:1990:SSP

[AA90] Monte Aronson and Alvera L. Aronson. SAS system: a programmer’s guide. J. Ranade IBM series. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-002467-7. xiii + 322 pp. LCCN QA276.4 .A76 1990. Includes index.

Anderson-Cook:2004:BRBa

[AC04] Christine M. Anderson-Cook. Book review: Regression and ANOVA: An Integrated Approach Using SAS Software, by Keith E. Muller; Bethel A. Fetterman. The American Statistician, 58(2):172–173, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                                    tronic). URL http://www.
                                                    jstor.org/stable/27643532.

Anderson-Cook:2008:BRBd

[AC08] Christine M. Anderson-Cook.

Book review: Optimum Experimental Designs, With SAS. Journal of the American Statistical Association, 103(483): 1327–1328, September 2008. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Anderson-Cook:2004:BRBb

[ACL04] Christine M. Anderson-Cook and Russell V. Lenth. Book review: Step-by-Step Basic Statistics Using SAS: Student Guide, by Larry Hatcher; Stepby-Step Basic Statistics Using SAS: Exercises by Larry Hatcher. The American Statistician, 58(3):262, August 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/27643572.

ACM:1992:CPI

[ACM92] ACM, editor. Conference proceedings / 1992 International Conference on Supercomputing, July 19–23, 1992, Washington, DC. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

ACM:1996:SAI

[ACM96] ACM, editor. SIGDOC ’96.

The 14th Annual International Conference, Conference Proceedings. Marshaling New Technological Forces: Building a Corporate, Academic, and User-Oriented Triangle, October 20–23, 1996, Research Triangle Park, NC, USA. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Agresti:2013:CDA

[Agr13] Alan Agresti. Categorical data analysis, volume 792 of Wiley series in probability and statistics. John Wiley and Sons, Inc., New York, NY, USA, third edition, 2013. ISBN 0-470-46363-5 (hardcover). xvi + 714 pp. LCCN QA278 .A353 2013.

Ahmed:2008:BRBd

[Ahm08]  S. E. Ahmed.  Book re-

view: Univariate and Multivariate General Linear Models: Theory and Applications with SAS by K. Kim; N. Timm. Technometrics, 50(1): 97–98, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25471448.

Allen:1985:SLG

[AK85] Alice T. Allen and Brenda C. Kalt. SAS language guide for personal computers, version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-80-3. xvi + 429 pp. LCCN QA276.4 S19 1985. Edited by Alice T. Allen and Brenda C. Kalt. Includes index.

Alexander:1991:BRB

[Ale91] Melvin T. Alexander. Book review: Learning Statistics with the SAS System by Rudolf Freund. Technometrics, 33(1):110– 112, February 1991. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1269017.

Allen:1982:SUG

[All82a] Alice T. Allen. SAS user’s guide: basics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. ISBN 0-917382-36-6. 923 pp. LCCN QA276.4 .S27 1982. Includes bibliographical references and index.

Allen:1982:SEU

[All82b] Alice T. Allen. SAS/ETS user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. ISBN 0-917382-38-2 (paperback). 398 pp. LCCN HA30.3 .S28 1982, QA276.4 .S234 1982. Edited by Alice Allen Ray. Includes bibliographies and index.

Allen:1986:CSS

[All86] David M. Allen, editor. Computer science and statistics: Proceedings of the Seventeenth Symposium on the Interface of Computer Sciences and Statistics, Radisson Plaza Hotel, Lexington, Kentucky, March 17–19, 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/13455449.
html;   http://www.dtic.
mil/dtic/tr/fulltext/u2/
a166019.pdf;   http://zbmath.
org/?q=an:0647.62008.

Allison:1995:SAU

[All95] Paul David Allison. Survival Analysis Using the SAS System: a Practical Guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-279-X, 1-59047-532-1 (e-book). viii + 292 pp. LCCN QA276.4 .A394 1995.

Allen:2006:BRB

[All06] Jeff M. Allen. Book review: SAS for Monte Carlo Studies: A Guide for Quantitative Researchers, by Xitao Fan; Ákos Felsővályi; Stephen A. Sivo; Sean C. Keenan. Journal of the American Statistical Association, 101(476):1728– 1729, December 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/
stable/27639799.

Amini:1987:TMS

[Ami87] Saeid B. Amini. Two mainframe SAS(R) programs for nonparametric tests for homogeneity of dose-response growth curves and regression analysis of tumor-prevalence data. The American Statistician, 41 (3):237, August 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2685116.

Anderson:1986:BRA

[And86] A. J. B. Anderson. Book review: Applied Statistics and the SAS Programming Language, by R. P. Cody, J. K. Smith. Biometrics, 42(4):1004– 1005, December 1986. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/2530725.

Anderson:1989:TPA

[And89] B. W. Anderson. Tracking projects-at-a-glance: applying the SAS/GRAPH product as a project management tool. Computers & industrial engineering, 17(1):154–158, 1989. CODEN CINDDL. ISSN 0360-8352.

Andersen:2010:BRB

[And10] Per Kragh Andersen. Book review: Regression Modeling: Methods, Theory, and Computation with SAS, by M. Panik. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173(3):696, July 2010. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

jstor.org/stable/40666293.

Annis:2005:PSB

[Ann05] Izabela E. Annis. PROC SQL:

Beyond the basics using SAS.

The American Statistician, 59 (4):350, November 2005. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:1983:Sa

[Ano83] Anonymous. SAS. Computational Statistics & Data Analysis, 1(??):61–62, March 1983. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0167947383900683.

Anonymous:1984:SBP

[Ano84a]  Anonymous.  SAS busi-

ness planning and project management tool. Computational Statistics & Data Analysis, 2(3):262, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0167947384900215.

Anonymous:1984:SSN

[Ano84b] Anonymous. The SAS system now for minicomputers. Computational Statistics & Data Analysis, 2(3):260–261, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0167947384900203.

Anonymous:1984:SGE

[Ano84c]  Anonymous.  SAS/GRAPH

enhancements. Computational Statistics & Data Analysis, 2(3):263–264, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0167947384900239.

Anonymous:1984:SFR

[Ano84d] Anonymous. SAS/OR, FS-CALC, REPLAY-CICS. Computational Statistics & Data Analysis, 2(3):264, November 1984. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0167947384900240.

Anonymous:1988:NEG

[Ano88] Anonymous. New expanded guide of the SAS system for personal computers. Computational Statistics & Data Analysis, 7(1):101–102, August 1988. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0167947388900217.

Anonymous:1993:BRBh

[Ano93a]  Anonymous.  Book review:

SAS kompakt für die Version 6: Göttsche, T. (1992) Stuttgart: Gustav Fischer Verlag, pp. 192, ISBN 3-437-40281-1, DM 49.- . Computational Statistics & Data Analysis, 16(2):248, August 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
016794739390123B.

Anonymous:1993:BRBc

[Ano93b]  Anonymous.  Book review:

SAS system — eine praxisbezogene einführung: Frank Multrus and Michael Bleicher (1992): Hidelberg: Hüthig Buch Verlag, ISBN 3-7785-2157-8, pp. 241, DM 68. Computational Statistics & Data Analysis, 15(4):469, May 1993. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
016794739390182S.

Anonymous:1994:SIBb

[Ano94a] Anonymous. SAS Institute’s business analysis software now supports several Unix platforms. Open Systems Today, 161:8–??, October 1994. ISSN 1061-0839.

Anonymous:1994:SOD

[Ano94b] Anonymous. SAS ODBC driver technical report: user’s guide and programmer’s reference: release 6.10. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-630-2. viii + 49 pp. LCCN ????

Anonymous:1995:ACA

[Ano95a] Anonymous, editor. 20th Annual conference — April 1995, Orlando, FL, USA, SUGI 1995; 20th. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-211-0. LCCN ????

Anonymous:1995:Fh

[Ano95b] Anonymous. Framework. Datamation, 41(8):9–??, May 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:SOD

[Ano95c] Anonymous. SAS ODBC driver technical report: user’s guide and programmer’s reference: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-236-6. viii + 50 pp. LCCN ????

Anonymous:1996:PTA

[Ano96a] Anonymous, editor. Proceedings of the Twenty-First Annual SAS Users Group International Conference, SUGI 21. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. Two volumes.

Anonymous:1996:PTF

[Ano96b] Anonymous, editor. Proceedings of the Twenty-First Annual SAS Users Group International Conference, SUGI 21, Chicago, IL, USA, 10– 13 March 1996. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996.

ISBN ???? LCCN ???? Two volumes.

Anonymous:1997:SUG

[Ano97] Anonymous, editor. SAS Users Group: Annual international conference; 22nd — March 1997, San Diego, CA, SUGI 1997; CONF 22. SAS, ????, 1997. ISBN 1-55544-961-1. LCCN ????

Anonymous:1998:SUG

[Ano98]  Anonymous, editor.  SAS

Users Group: Annual international conference; 23rd — March 1998, Nashville, TN, SUGI 1998; CONF 23. SAS, ????, 1998. ISBN 1-58025-149-8. LCCN ????

Anonymous:1999:BRAq

[Ano99]  Anonymous.  Book review:

Applied multivariate statistics with SAS software, second edition: By Ravindra Khattree and Dayanand Naik. SAS Institute, Inc./John Wiley & Sons, Cary, NC/New York. (1999). 338 pages. $59.95. Computers and Mathematics with Applications, 38(5–6):255, September 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122199901578.

Anonymous:2000:AMS

[Ano00a] Anonymous. Applied Multivariate Statistics with SAS Software. The American Statistician, 54(2):154–??, May 2000.

CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Anonymous:2000:NPI

[Ano00b] Anonymous. New products:

IQ2 NetCONNECT, BAS-COM Global Internet Services, Inc.; OpenLinux eServer 2.3, Caldera Systems, Inc.; CommuniGate Pro, Super-Symmetric Dynamic Cluster Version, Stalker Software, Inc.; Embedix Linux 1.0, Lineo, Inc.; white dwarf linux 1.0, EMJ Embedded Systems; Linuxbased Internal G.lite Modem, SAS; IntraLinux, STS International, Inc.; Linux End-to-End Security Solution, JAWS Technologies; JetForm Central for Linux, JetForm Corporation; JWAVE version 3.0, Visual Numerics, Inc.; RTEL, Lantronix; Japanese Word-Mage v.5.7, Lava Software Pty. Ltd.; TotalView 4.0 Parallel Debugger, Etnus; iNUX Consumer Linux Personal Computer systems, iNUX Inc. Linux Journal, 73:??, May 2000. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2003:BRBt

[Ano03a]  Anonymous.  Book review:

Common Statistical Methods for Clinical Research with SAS(R) Examples by Glenn A. Walker. Technometrics, 45(2): 182–183, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/25047038.

Anonymous:2003:BRBal

[Ano03b]  Anonymous.  Book review:

SAS(R) for Monte Carlo Studies by Xitao Fan; Ákos Felsővályi; Stephen A. Sivo; Sean C. Keenan. Technometrics, 45(4):373, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/25047113.

Anonymous:2004:BRBg

[Ano04a]  Anonymous.  Book review:

SAS(R) for Forecasting Time Series by John C. Brocklebank; David A. Dickey. Technometrics, 46(1):120–121, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/25470788.

Anonymous:2004:SAS

[Ano04b] Anonymous. SAS/ACCESS 9.1.2 supplement for MySQL: SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2004. ISBN 1-59047-458-9 (paperback). iv + 23 pp. LCCN QA76.9.D3 S25463 2004.

Anonymous:2006:BRBf

[Ano06a]  Anonymous.  Book review:

Fixed Effects Regression Methods for Longitudinal Data Using SAS(R) by Paul D. Allison. Technometrics, 48(1):157– 158, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25471144.

Anonymous:2006:BRSc

[Ano06b]  Anonymous.  Book review:

Statistical Analysis of Medical Data Using SAS, by Der, G. and Everitt, B. S., Taylor & Francis/CRC Press, Boca Raton, Florida, 2006. vi + 428 pp. US$79.95/£39.99, ISBN 1-58488-469-X. Biometrics, 62 (2):634–635, June 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2006:BRBo

[Ano06c] Anonymous. Book reviews:

A Step-by-Step Approach to Using SAS(R) for Univariate & Multivariate Statistics by Norm O’Rourke; Larry Hatcher; Edward J. Stepanski; JMP(R) for Basic Univariate and Multivariate Statistics by Ann Lehman; Norm O’Rourke; Larry Hatcher; Edward J. Stepanski. Technometrics, 48(2):314, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/25471186.

Anonymous:2006:BRBz

[Ano06d] Anonymous. Book reviews: Applied Statistics and the SAS(R) Programming Language by Ronald P. Cody; Jeffrey K. Smith; Statistical Analysis of Medical Data Using SAS by Geoff Der; Brian S. Everitt. Technometrics, 48(3):449, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/25471224.

Anonymous:2006:SAS

[Ano06e] Anonymous. SAS/ACCESS 9.1.3 supplement for MYSQL: SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2006. ISBN 1-59047-924-6. 23 pp. LCCN MLCM 2006/42439 (T).

Anonymous:2008:BRAb

[Ano08a]  Anonymous.  Book review:

Analysis of Correlated Data with SAS and R, 3rd edition. by Shoukri, M. M. and Chaudhary, M. A. Biometrics, 64 (2):660–661, June 2008. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2008:BRBe

[Ano08b]  Anonymous.  Book review:

SAS Graphics for Java: Examples Using SAS AppDev Studio and the Output Delivery System by Wendy Bohnenkamp; Jackie Iverson. Technometrics, 50(3): 412, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/25471510.

Anonymous:2009:BRBm

[Ano09a]  Anonymous.  Book review:

Handbook of Statistical Analyses Using SAS (3rd ed.) by Geoff Der; Brian S. Everitt. Technometrics, 51(4): 488, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/40586671.

Anonymous:2009:BRH

[Ano09b]  Anonymous.  Book review:

Handbook of Statistical Analyses Using SAS, 3rd edition by Der, G. and Everitt, B. S. Biometrics, 65(4):1312– 1313, December 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRRb

[Ano09c]  Anonymous.  Book review:

R for SAS and SPSS Users, by Muenchen, R. A. Biometrics, 65(4):1313, December 2009. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2009:BRSb

[Ano09d]  Anonymous.  Book review:

SAS for Data Analysis: Intermediate Statistical Methods, by Marasinghe, M. G. and Kennedy, W. J. Biometrics, 65(2):672, June 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2011:BRBb

[Ano11a] Anonymous. Book review: Statistical Data Mining Using SAS Applications by George Fernandez. Technometrics, 53(1): 107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/40997313.

Anonymous:2011:BRSa

[Ano11b]  Anonymous.  Book review:

Survival Analysis Using SAS: a Practical Guide, by Allison, P. D. Biometrics, 68(3): 1182, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:2012:BRC

[Ano12]  Anonymous.  Book review:

Common Statistical Methods for Clinical Research with SAS Examples, by Walker, G. A. and Shostak, J. Biometrics, 68(1):330, March 2012. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Altenburg:1991:AQQ

[AR91] H.-P. Altenburg and G. Rosenkranz. Analysis of quantal and quantitative dose-response data using SAS. The American Statistician, 45(2):159, May 1991. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/2684392.

Aster:1991:PSP

[AS91] Rick Aster and Rhena Seidman.

Professional SAS programming secrets. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-7612-

0. 577 pp. LCCN QA276.4 .A84 1991. US$34.95. Includes index.

Aster:1996:PSP

[AS96] Rick Aster and Rhena Seidman.

Professional SAS programming secrets. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-913095-X (paperback). ???? pp. LCCN ????

Al-Subaihi:2002:VSM

[AS02] Ali A. Al-Subaihi. Variable selection in multivariable regression using SAS/IML. Journal of Statistical Software, 7(12): 1–20, ???? 2002. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v07/i12;   http://www.
jstatsoft.org/v07/i12/mv.
pdf.

Al-Subaihi:2004:SCM

[AS04] Ali A. Al-Subaihi. Simulating correlated multivariate pseudorandom numbers. Journal of Statistical Software, 9(4):1– 20, ???? 2004. CODEN JS-SOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/
counter.php?id=85&url=v09/
i04/paper.pdf&ct=1.

Aster:1992:PSU

[Ast92] Rick Aster. Professional SAS user interfaces. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1992. ISBN 0-8306-3717-6 (paperback). 460 pp. LCCN QA276.4 .A85 1992. US$34.95. Includes index.

Aster:1993:PSP

[Ast93] Rick Aster. Professional SAS programmer’s pocket reference. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1993. ISBN 0-8306-4317-6 (p). 188 pp. LCCN QA276.4 .A83 1993. Includes index.

Aster:1994:LAS

[Ast94a] Rick Aster. A look at SAS files, 1994. 2 sound cassettes.

Aster:1994:SFI

[Ast94b] Rick Aster. SAS foundations: from installation to operation. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1994. ISBN 0-8306-4316-8. xii + 279 pp. LCCN QA276.4 .A86 1994. US$28.95.

Alpuente:2008:SAI

[AV08] María Alpuente and Germán Vidal, editors. Static Analysis: 15th International Symposium, SAS 2008, Valencia, Spain, July 16–18, 2008. Proceedings, volume 5079 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2008. CO-DEN LNCSD9. ISBN 3-540-69163-4 (print), 3-540-69166-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.

springerlink.com/content/
978-3-540-69166-2.

Bailer:2019:SPSb

[Bai19] A. John Bailer. Statistical Programming in SAS. Chapman and Hall/CRC, Boca Raton, FL, USA, second edition, 2019. ISBN 0-367-35800-X (hardcover), 0-367-35797-6 (paperback), 0-429-34184-9 (e-book), 1-000-73492-7 (PDF e-book), 1-000-73517-6 (Mobipocket ebook), 1-000-73542-7 (ePub ebook). LCCN QA276.4 .B355 2020.

Barr:1976:UGS

[Bar76] Anthony J. Barr. A user’s guide to SAS, 76. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1976. ISBN 0-917382-01-3. 329 pp. LCCN QA276.4 .U93 1976. Includes bibliographical references and index.

Barr:1979:SUG

[Bar79] Anthony J. Barr. SAS user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1979. ISBN 0-917382-06-3 (??invalid checksum??). 494 pp. LCCN QA276.4 .S230 1979. “Anthony J. Barr designed and implemented the SAS supervisor … ” Includes bibliographies and index.

Barcikowski:1983:CPR

[Bar83] Robert S. Barcikowski. Computer Packages and Research Design: With Annotations of Input and Output from the BMDP, SAS, SPSS, and SPSSX Statistical Packages. University Press of America, Lanham, MD, USA, 1983. ISBN 0-8191-3494-5 (v. 1: paperback), 0-8191-3495-3 (v. 2: paperback), 0-8191-3496-1 (v. 3: paperback). xiv + 781 pp. LCCN QA276.4.C576 1983.

Bass:1984:TSL

[Bas84] J. Bass. Translating the SAS language into BASIC. Byte Magazine, 9(1):417–434, January 1984. CODEN BYTEDJ. ISSN 0360-5280.

Bassett:2003:BRB

[Bas03] Roland Bassett. Book review: Debugging SAS Programs: A Handbook of Tools and Techniques, by Michele M. Burlew. Journal of the American Statistical Association, 98(464):1087, December 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http:

//miranda.asa.catchword.
org/vl=1515603/cl=66/nw=1/
rpsv/cw/asa/01621459/v98n464/
s40/p1087; http://www.jstor.
org/stable/30045361.

Berta:2017:CSM

[BBV17] Paolo Berta, Matteo Bossi, and Stefano Verzillo. %CEM: a SAS macro to perform coarsened exact matching. Journal of Statistical Computation and Simulation, 87(2):227–238, 2017. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Brocklebank:1986:SSF

[BD86] John C. Brocklebank and David A. Dickey. SAS system for forecasting time series. SAS series in statistical applications. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 1-55544-027-4. x + 240 pp. LCCN QA276.4.B7 1986. Includes bibliographical references (p. [233]- 234) and index.

Brocklebank:2003:SFT

[BD03] John C. Brocklebank and David A. Dickey. SAS for forecasting time series. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 1-59047-182-2 (SAS Institute), 0-471-39566-8 (Wiley). x + 398 pp. LCCN QA280 .B754 2003.

Beebe:2005:SSF

[Bee05] Nelson H. F. Beebe. SAS (statistics) FAQ. World-Wide Web frequently-asked question document., September 2005. URL https://

www.math.utah.edu/faq/sas/

. This report is updated frequently.

Berglund:2014:MIM

[BH14] Patricia Berglund and Steven Heeringa. Multiple Imputation of Missing Data Using SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2014. ISBN 1-61290-452-1, 1-62959-203-X, 1-62959-205-6 (e-book). xii + 150 pp. LCCN HA31.2 .B37 2014.

Bohnenkamp:2007:SGJ

[BI07] Wendy Bohnenkamp and Jackie Iverson. SAS Graphics for Java: examples using SAS AppDev studio and the Output delivery system. SAS Press series. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2007. ISBN 1-59047-693-X. xii + 342 pp. LCCN HA31 .B645 2007.

Biesheuvel:2008:BRP

[Bie08] Egbert Biesheuvel. Book review: Pharmaceutical statistics using SAS(R): a practical guide. Alex Dmitrienko, Christy Chuang-Stein and Ralph D’Agostino (eds), SAS(R) Press, Cary, NC, 2007. No. of pages: ix + 443. Price: $69.95. ISBN: 978-1-59047-886-8. Statistics in Medicine, 27 (10):1798–1799, May 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Birkett:1994:TRN

[Bir94] Thomas R. Birkett. Trip report for November 20 to December 15, 1994: analysis of the 1993-94 wheat data and 1994-95 cotton data: updating the wheat and cotton yield forecasting models: listing, with documentation in comments, of SAS software for summarization and modeling and plots of ASF cotton and wheat models and data. USDA/NASS, Washington, DC, USA, December 18, 1994. various pp.

Bjornaraa:2001:PPN

[Bjø01] Tore I. Bjørnarå, editor. Program and proceedings for the Nordic MATLAB Conference 2001: October 17–18, 2001, Radisson SAS Scandinavia Hotel, Oslo. Comsol, Trondheim, Norway, 2001. ISBN 82-995955-0-9. LCCN ????

Burch:1994:SSR

[BK94] Laurie Burch and SherriJoyce King. SAS software roadmaps: your guide to discovering the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-606-X. xii + 317 pp. LCCN ????

Brezger:2005:BAB

[BKL05] Andreas Brezger, Thomas Kneib, and Stefan Lang. BayesX: Analyzing Bayesian structural additive regression models. Journal of Statistical Software, 14(11):1–22, September 9, 2005. CO-DEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/counter.php?
id=136&url=v14/i11&ct=2;
http://www.jstatsoft.org/
counter.php?id=136&url=v14/
i11/v14i11.pdf&ct=1.

Bahler:1998:SHH

[BMDB98] C. C. Bahler, S. Muller, D. Doolittle, and A. Barrios.

SAS and HTML — HTML publishing using SAS. In Anonymous [Ano98], pages 229–237. ISBN 1-58025-149-8. LCCN ????

Bunouf:2015:SPC

[BMGT15] Pierre Bunouf, Geert Molenberghs, Jean-Marie Grouin, and Herbert Thijs. A SAS program combining R functionalities to implement patternmixture models. Journal of Statistical Software, 68(8):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/
jss/article/view/v068i08;
 http://www.jstatsoft.org/
index.php/jss/article/view/
v068i08/v68i08.pdf.

Backstrom:1993:CRSa

[BN93a] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Research Report LiTH-IDA-R-93-34, Department of Computer and Information Science, Linköping University, Linköping, Sweden, Linköping, Sweden, 1993. 39 pp. A revised version of this paper has been published in Computational Intelligence.

Backstrom:1993:CRSb

[BN93b] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. In Proceedings of the 13th International Joint Conference on Artificial Intelligence, pages 1430–1435. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, August 1993. A long version of this paper has been published in Computational Intelligence.

Backstrom:1995:CRS

[BN95a] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Comput. Intelligence, 11(4):625– 655, 1995. ISSN 0824-7935.

Baeckstroem:1995:CRS

[BN95b] Christer Bäckström and Bernhard Nebel. Complexity results for SAS^{+} planning. Computational Intelligence, 11(4): 625–655, 1995. CODEN COMIE6. ISSN 0824-7935 (print), 1467-8640 (electronic).

URL ftp://ftp.informatik.
uni-ulm.de/pub/papers/ki/
backstrom-nebel-ci-95.ps.

Boscardin:2007:BRS

[Bos07] W. John Boscardin. Book reviews: SAS(R) for mixed models (2nd edn). Ramon C. Littell, George A. Milliken, Walter W. Stroup, Russell D. Wolfinger and Oliver Schabenberger, SAS Institute, Cary, NC, 2006. No. of pages: xii + 814. Price: $89.95. ISBN 10: 1-59047-500-3; ISBN 13: 978-1-59047-500-3. Statistics in Medicine, 26(29): 5354–5355, December 20, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandos:2009:BRA

[BR09] Andriy I. Bandos and Howard E. Rockette. Book review: Analyzing receiver operating characteristic curves with SAS(R). Mithat Gönen, SAS Institute Inc., Cary, NC, 2007. No. of pages: x + 134. Price: $31.95. ISBN 978-1-59994-298-8. Statistics in Medicine, 28 (3):538–539, February 10, 2009. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bradstreet:2008:PSU

[Bra08] Thomas E. Bradstreet. Pharmaceutical statistics using SAS: a practical guide. The American Statistician, 62(2):183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Braun:2015:BRA

[Bra15] Thomas M. Braun. Book review: Adaptive Design Theory and Implementation Using SAS and R, Second Edition. International Statistical Review = Revue Internationale de Statistique, 83(3):512–513, December 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Brown:1989:GS

[Bro89] Diane E. Brown. Guide to SAS/DB2. The database experts’ series. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-008207-3. x + 444 pp.

LCCN QA76.9.D3 B768 1989. US$39.95.

Barnum:1995:AVS

[BS95] J. Barnum and C. Sampson.

Adding value to SAS data with SGML. In Anonymous [Ano95a], pages 750–759. ISBN 1-55544-211-0. LCCN ????

Ball:1989:SSR

[BSW89] Amy E. Ball, Sandra D.

Schlotzhauer, and Helen F. Weeks. SAS/QC software: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-327-3. xxii + 660 pp. LCCN QA276.4.S23888 1989. Writing and editing by Amy E. Ball, Sandra D. Schlotzhauer, Helen F. Weeks. Includes bibliographies and index.

Buncher:2007:BRP

[Bun07] C. Ralph Buncher. Book review: Pharmaceutical Statistics Using SAS — A Practical Guide, edited by A. Dmitrienko, C. Chuang-Stein and R. D’Agostino. Biometrics, 63(4):1301, December 2007. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burgard:1992:MMY

[Bur92] Mike Burgard. Making the Most Out of Your Data. UNIX/world, 9(9):85–??, September 1992. ISSN 0739-5922.

Brand:2003:TSE

[BvBGOG03] Jaap P. L. Brand, Stef van Buuren, Karin Groothuis-Oudshoorn, and Edzard S. Gelsema. A toolkit in SAS for the evaluation of multiple imputation methods. Statistica Neerlandica, 57(1):36–45, February 2003. CODEN ???? ISSN 0039-0402 (print), 1467-9574 (electronic). URL https:

//onlinelibrary.wiley.com/
doi/epdf/10.1111/1467-9574.
00219.

Baayen:2019:STC

[BVFB19] C. Baayen, C. Volteau, C. Flamant, and P. Blanche. Sequential trials in the context of competing risks: Concepts and case study, with R and SAS code. Statistics in Medicine, 38(19): 3682–3702, August 30, 2019. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bzik:2011:BRB

[Bzi11] Thomas Bzik. Book review: Just Enough SAS(R): A Quick-Start Guide to SAS(R) for Engineers by Robert A. Rutledge. Technometrics, 53(2): 216–217, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/23209742.

Crum:1985:SGP

[CA85] Larry D. Crum and Alice T. Allen. SAS/IML guide for personal computers, version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-96-X. xvi + 244 pp. LCCN QA276.4 S17 1985. Edited by Larry D. Crum and Alice T. Allen. Includes bibliographical references (p. [235]) and index.

Caby:2004:BRB

[Cab04] Errol C. Caby. Book review:

Data Mining Using SAS Applications by George Fernandez. Technometrics, 46(2): 260–261, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25470822.

Cantley:1993:GDS

[Can93] Ian Cantley. Generating descriptive statistics using the SAS system. University of Ulster, Newtownabbey, Co. Antrim, 1993. ISBN 1-85889-095-0 (plastic folder). LCCN ????

Cartier:1993:SML

[Car93] Jeff Cartier. SAS macro language course notes. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-539-X. xii + 372 pp. LCCN ????

Crum:1985:SUG

[CB85] Larry D. Crum and David D.

Baggett. SAS/IML user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition edition, 1985. ISBN 0-917382-86-2. xv + 300 pp. LCCN QA276.4 .S235 1985. “Edited by Larry D. Crum; copyedited by David D. Baggett.” Includes index.

Carrigan:2007:CMD

[CBDM07] Gretchen Carrigan, Adrian G. Barnett, Annette J. Dobson, and Gita Mishra. Compensating for missing data from longitudinal studies using Win-BUGS. Journal of Statistical Software, 19(7):1–17, June 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v19/i07.

Caussinus:1982:CSH

[CEM82] H. Caussinus, P. Ettinger, and J. R. Mathiew, editors. COMPSTAT 1982: 5th Symposium held at Toulouse, 1982. Physica-Verlag, Wien, Austria, 1982. ISBN 3-7051-0002-5 (vol. 1), 3-7051-0001-7 (vol. 2). LCCN QA276.4 .C194 1982. Two volumes.

Cortesi:1999:SAI

[CF99] Agostino Cortesi and Gilberto Filé, editors. Static analysis: 6th International Symposium, SAS’99, Venice, Italy, September 22–24, 1999: Proceedings, volume 1694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1999. CO-DEN LNCSD9. ISBN 3-540-66459-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1999. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1694.htm;   http://www.
springerlink.com/content/
978-3-540-66459-8;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1694.

Council:1981:SUG

[CH81] Kathryn A. Council and Jane T. Helwig. SAS/Graph User’s Guide: 1981 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. 126 pp. LCCN QA276.4 .S767 1981.

Chan:2004:BRB

[Cha04a] Grace Chan. Book review: SAS Programming: The One-Day Course, by Neil H. Spencer. The American Statistician, 58 (3):260, August 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/27643570.

Chan:2004:SPO

[Cha04b] Grace Chan. SAS programming: The one-day course. The American Statistician, 58(3): 260, August 2004. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Chang:2008:ADT

[Cha08] Mark Chang. Adaptive design theory and implementation using SAS and R. Chapman and Hall/CRC biostatistics series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2008. ISBN 1-58488-962-4. xxii + 418 pp. LCCN R853.C55 C42

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0801/2007011412-d.html;
 http://www.loc.gov/catdir/
toc/ecip0713/2007011412.
html.

Chen:1988:SDD

[Che88] Hsin-Chu Chen. The SAS domain decomposition method for structural analysis. Technical Report CSRD 754; UILU-ENG-88-8003, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1988. vi + 112 pp.

Chen:1992:PSM

[Che92] H.-C. Chen. Parallel SAS multicluster algorithms for solving linear systems with reflexive coefficient matrices. In ACM [ACM92], pages 447–455. ISBN 0-89791-485-6 (paperback), 0-89791-486-4. LCCN QA 76.88 I57 1992. Sponsored by ACM SIGARCH.

Chernick:2008:BRBa

[Che08] Michael R. Chernick. Book review: Pharmaceutical Statistics Using SAS(R): A Practical Guide by Alex Dmitrienko; Christy Chuang-Stein; Ralph D’Agostino Sr. Technometrics, 50(1):96, February 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/25471444.

Chiu:2002:BRM

[Chi02] Yen-Feng Chiu. Book review:

Multiple Comparisons and Multiple Tests. Using the SAS System. Peter H. Westfall, Randall D.Tobias, Dror Rom, Russell D. Wolfinger and Yosef Hochberg, SAS Institute, Cary, U.S.A. 2000. No. of pages: xiv + 397. Price: DKK 412.00. ISBN 1-58025-397-0. Statistics in Medicine, 21(10):1499– 1500, May 30, 2002. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiang:2006:BRB

[Chi06] Alan Y. Chiang. Book review:

Analysis of Clinical Trials Using SAS(R): A Practical Guide by Alex Dmitrienko; Geert Molenberghs; Christy Chuang-Stein; Walter Offen. Technometrics, 48(2):311, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/25471180.

Council:1995:HSI

[CHM95] Forrest M. Council, Elizabeth G. Hamilton, and Yusuf Mohamedshah. Highway safety information system guidebook for the North Carolina state data files. Federal Highway Administration, Washington, DC, USA, 1995. 106 pp.

Clark:1990:IDN

[CL90] Hal G. Clark and Leonard Lorensen. Illustrations of departures from the new standard auditor’s report on financial statements of business enterprises: a survey of the application of SAS no. 58. Financial report survey; 41. American Institute of Certified Public Accountants, New York, NY, USA, 1990. ISBN 0-87051-083-5. 136 pp. LCCN HF5667 .C56 1990. Includes bibliographical references (p. 133-136).

Clarke:2003:BRB

[Cla03] William R. Clarke. Book review: Longitudinal Data and SAS: A Programmer’s Guide, by Ron Cody. Journal of the American Statistical Association, 98(463):773, September 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

                                        jstor.org/stable/30045316.

Cleary:2003:BRB

[Cle03a] Richard J. Cleary. Book review: Conducting Meta Analysis Using SAS, by Winfred Arthur, Jr.; Winston Bennett, Jr.; Allen I. Huffcutt. Journal of the American Statistical Association, 98(461):254, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.

                                        jstor.org/stable/30045221.

Cleary:2003:CMA

[Cle03b]  Richard J. Cleary.  Con-

ducting meta analysis using SAS. Journal of the American Statistical Association, 98(461):254–??, March 2003. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://

fidelio.ingentaselect.com/
cgi-bin/linker?ini=asa&reqidx=
/cw/asa/01621459/v98n461/
s34/p254.

Ciampi:1982:EVS

[CM82] Constantino Ciampi and A. A.

Martino, editors. Edited versions of selected papers from the International Conference on “Logic, Informatics, Law,” Florence, Italy, April 6–10, 1981. Elsevier Science Publishers, Amsterdam, The Netherlands, 1982. ISBN 0-444-86413-X (set), 0-444-86414-8 (vol. 1), 0-444-86415-6 (vol. 2). LCCN K662.I4 I58. Two volumes. Vol. 1: Artificial intelligence and legal information systems. Vol. 2: Deontic logic, computational linguistics, and legal information systems.

Cousot:2011:SAI

[CM11] Radhia Cousot and Matthieu Martel, editors. Static Analysis: 17th International Symposium, SAS 2010, Perpignan, France, September 14–16, 2010. Proceedings, volume 6337 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-15768-8 (print), 3-642-15769-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-642-15769-1.

Cotsonis:2006:SYS

[Cot06] George Cotsonis. Sharpening your SAS skills. The American Statistician, 60(4):346, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Council:1980:SAG

[Cou80] Kathryn A. Council. SAS applications guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1980. ISBN 0-917382-16-1. 201 pp. LCCN QA276.4.S232 1980. Includes index.

Cousot:2001:SAI

[Cou01] Patrick Cousot, editor. Static analysis: 8th International Symposium, SAS 2001,Paris, France, July 16–18, 2001: Proceedings, volume 2126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2001. CO-DEN LNCSD9. ISBN 3-540-42314-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 2001; QA267.A1 L43 no.2126. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t2126.htm;   http://www.

springerlink.com/content/
978-3-540-42314-0;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=2126.

Cousot:2003:SAI

[Cou03] Radhia Cousot, editor. Static analysis: 10th International Symposium, SAS 2003, San Diego, CA, USA, June 11–13, 2003: Proceedings, volume 2694 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2003. CODEN LNCSD9. ISBN 3-540-40325-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 2003. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t2694.htm;   http://www.
springerlink.com/content/
978-3-540-40325-8;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=2694.

Cowles:2004:BRBa

[Cow04a] Mary Kathryn Cowles. Book review: SAS for Linear Models (4th ed.). The American Statistician, 58(1):89– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v58n1/s21/p89;
 http://oberon.ingentaselect.
com/cgi-bin/linker?ini=asa&

                                        reqidx=/cw/asa/00031305/
                                        v58n1/s22/p89.

Cowles:2004:BRSb

[Cow04b] Mary Kathryn Cowles. Book review: SAS for Linear Models (4th ed.). The American Statistician, 58(1):89– ??, February 2004. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://oberon.
                                        ingentaselect.com/cgi-bin/
                                        linker?ini=asa&reqidx=/cw/
                                        asa/00031305/v58n1/s22/p89.

Couret:1994:LCA

[CP94] Dominique Couret and Françoise Pelletier. De l’enquête à la carte avec SAS. Collection Didactiques, 1142-2580. ORSTOM Editions, Paris, France, 1994. ISBN 2-7099-1179-5. 191 pp. LCCN ????

Cody:1995:SPE

[CPI95] Ronald P. Cody, Raymond Pass, and SAS Institute. SAS programming by example. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-681-7. viii + 337 pp. LCCN ????

Cody:1985:ASS

[CS85] Ronald P. Cody and Jeffrey K. Smith. Applied statistics and the SAS programming language. North-Holland, Amsterdam, The Netherlands, 1985. ISBN 0-444-00889-6 (paperback). viii + 187 pp. LCCN QA276.4 .C53 1985. Includes index.

Cody:1987:ASS

[CS87] Ronald P. Cody and Jeffrey K.

Smith. Applied statistics and the SAS programming language. North-Holland, Amsterdam, The Netherlands, second edition, 1987. ISBN 0-444-01192-7 (paperback). vii + 280 pp. LCCN QA276.4 .C53 1987. Includes index.

Chen:1989:DDM

[CS89] Hsin-Chu Chen and Ahmed Sameh. A domain decomposition method for 3D elasticity problems. Technical Report CSRD 890, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, September 1989. 18 pp.

Cody:1991:ASS

[CS91] Ronald P. Cody and Jeffrey K.

Smith. Applied statistics and the SAS programming language. North-Holland, Amsterdam, The Netherlands, third edition, 1991. ISBN 0-444-01619-8. xviii + 403 pp. LCCN QA276.4 .C53 1991. Includes index.

Cadigan:1993:SPF

[CS93] Noel G. Cadigan and Peter A.

Shelton. SAS programs for fitting a seal population dynamics model. Canadian technical report of fisheries and aquatic sciences; no. 1927. Science Branch, Dept. of Fisheries and Oceans, St. John’s, NF, Canada, 1993. iii + 33 pp.

Cousot:1996:SAT

[CS96] Radhia Cousot and David A. Schmidt, editors. Static analysis: Third International Symposium, SAS ’96, Aachen, Germany, September 24–26, 1996: proceedings, volume 1145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1996. CODEN LNCSD9. ISBN 3-540-61739-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1996. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1145.htm;   http://www.
springerlink.com/content/
978-3-540-61739-6;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1145.

Cody:1997:ASS

[CS97] Ronald P. Cody and Jeffrey K. Smith. Applied statistics and the SAS programming language. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, fourth edition, 1997. ISBN 0-13-743642-4. xv + 445 pp. LCCN QA276.4.C53 1997.

Carpenter:1995:QRS

[CSS95] Arthur L. Carpenter, Charles E. Shipp, and SAS Institute, Inc. Quick results with SAS/ GRAPH software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-683-3. xvi + 249 pp. LCCN ????

Cherniak:2013:OSB

[CZZ13] Andrii Cherniak, Huma Zaidi, and Vladimir Zadorozhny. Optimization strategies for A/B testing on HADOOP. Proceedings of the VLDB Endowment, 6(11):973–984, August 2013. CODEN ???? ISSN 2150-8097.

Damico:2009:TRR

[Dam09] Anthony Damico. Transitioning to R: Replicating SAS, Stata, and SUDAAN analysis techniques in health policy data. The R Journal, 1(2):37–44, December 2009. CODEN ???? ISSN 2073-4859. URL http:

//journal.r-project.org/
archive/2009-2/RJournal_
2009-2_Damico.pdf.

Dmitrienko:2007:PSU

[DCSD07] Alex Dmitrienko, Christy Chuang-Stein, and Ralph B. D’Agostino, editors. Pharmaceutical Statistics Using SAS: a Practical Guide. SAS Press series. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2007. ISBN 1-59047-886-X, 1-59994-357-3 (e-book). ix + 444 pp. LCCN RM301.27 .P48

2007. URL http://proquest.
safaribooksonline.com/9781590478868.

Dudewics:1989:MEP

[DCT89a] Edward J. Dudewics, Pinyuen Chen, and Baldeo K. Taneja. Modern elementary probability and statistics, with statistical programming in SAS, MINITAB, BMDP. American Sciences Press, Syracuse, NY, USA, 1989. ISBN 0-935950-19-2. ix + 375 pp. LCCN QA273 .D793 1989. US$34.50. Includes index. Includes bibliographical references.

Dudewicz:1989:MEP

[DCT89b] Edward J. Dudewicz, Pinyuen Chen, and Baldero Kumar Taneja. Modern elementary probability and statistics: with statistical programming in SAS, MINITAB, and BMDP. American Sciences Press, Columbus, OH, USA, 1989. ISBN 0-935950-19-2. ix + 375 pp. LCCN QA273 .D793 1989. US$34.50. Includes index. Includes bibliographical references.

Der:2002:HSA

[DE02] Geoff Der and Brian S. Everitt. A handbook of statistical analyses using SAS. Chapman and Hall/CRC, Boca Raton, FL, USA, 2002. ISBN 1-58488-245-X. xi + 360 pp. LCCN QA276.4 .D47 2002.

Der:2005:SAM

[DE05] Geoff Der and Brian S. Everitt. Statistical Analysis of Medical Data Using SAS. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2005. ISBN 1-58488-469-X (hardcover), 1-4200-5722-7 (e-book). 443 pp. LCCN RA409.5 .D465 2006.

Der:2008:HSA

[DE08] Geoff Der and Brian Everitt. A Handbook of Statistical Analyses Using SAS. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2008. ISBN 0-429-14697-3 (e-book), 1-322-63794-6, 1-58488-784-2 (hardback), 1-58488-785-0 (e-book). xv + 391 pp. LCCN QA276.4 .E88 2009.

Denny:2012:BRB

[Den12] Frances Denny. Book review:

SAS and R: Data Management, Statistical Analysis, and Graphics, by K. Kleinman; N. J. Horton. Journal of the Royal Statistical Society. Series A (Statistics in Society), 175 (3):826–827, July 2012. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/
stable/23251186.

Derquenne:2010:BRB

[Der10] Christian Derquenne. Book review: Regression Modeling: Methods, Theory, and Computation with SAS. Journal of Statistical Software, 36(BR-5): ??, September 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v36/b05.

DiIorio:1994:PSA

[DH94] Frank C. DiIorio and Debbie Hoppe, editors. Proceedings of the second annual Southeast SAS Users Group Conference:

SESUG ’94: the Omni Hotel at Charleston Place, Charleston, South Carolina, September 18– 20, 1994. The Group, United States, 1994.

DiIorio:1996:QSD

[DH96] Frank C. DiIorio and Kenneth A. Hardy. Quick start to data analysis with SAS. Duxbury Press, Belmont, CA, USA, 1996. ISBN 0-534-23760-6. xvi + 301 pp. LCCN ????

Dhand:2010:CSU

[Dha10] Navneet K. Dhand. Code snippet: UniLogistic: a SAS macro for descriptive and univariable logistic regression analyses. Journal of Statistical Software, 35(CS-1):??, July 2010. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v35/c01.

Dietz:2020:BRS

[Die20] Sebastian Dietz. Book review: Statistical Data Analysis using SAS. Journal of the Royal Statistical Society. Series A (Statistics in Society), 183 (4):1829, October 2020. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

DiIorio:1991:SAP

[DiI91] Frank C. DiIorio. SAS applications programming: a gentle introduction. The Duxbury series in statistics and decision sciences. PWS-KENT Pub. Co., Boston, MA, USA, 1991. ISBN 0-534-92390-9. xiv + 684 pp.

LCCN QA276.4 .D55 1991. Includes index.

Dinov:2006:SSO

[Din06] Ivo D. Dinov. SOCR: Statistics online computational resource. Journal of Statistical Software, 16(11):1–16, October 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v16/i11.

Dixon:2007:LSB

[Dix07] Philip Dixon. The little SAS book for enterprise guide 3.0. The American Statistician, 61 (3):277, August 2007. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobler:2004:BRBe

[DL04] Carolyn Pillers Dobler and Russell V. Lenth. Book review: Common Statistical Methods for Clinical Research with SAS Examples, by Glenn A. Walker. The American Statistician, 58 (2):175, May 2004. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/27643539.

Dobler:2006:SGR

[Dob06] Carolyn Pillers Dobler. SAS guide to report writing: Examples. The American Statistician, 60(3):294, August 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Dobbins:2009:BRL

[Dob09] Timothy Dobbins. Book review: Learning SAS by example: a programmer’s guide. Ron Cody, SAS Institute Inc: Cary, NC, 2007. No. of pages: xxxii + 626 (CD included). Price: $69.95. ISBN: 978-1-59994-165-3. Statistics in Medicine, 28(9): 1427–1428, April 30, 2009. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Downing:1988:BRB

[Dow88] Darryl J. Downing. Book review: SAS System for Forecasting Time Series by John C. Bocklebank; David A. Dickey. Technometrics, 30(1): 118, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270332.

Draper:2011:SBRb

[Dra11] Norman R. Draper. Short book review: Design and Analysis of Experiments with SAS by John Lawson. International Statistical Review = Revue Internationale de Statistique, 79 (1):118, April 2011. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/
stable/41306171.

Driscoll:1995:SPA

[Dri95] Daniel Driscoll. Strengthening policy analysis: econometric tests using SPSS/PC+, SAS PC, and GAUSS-386. Microcomputers in policy research; 2. International Food Policy Research Institute, Washington, DC, USA, 1995. ISBN 0-89629-330-0. ???? pp. LCCN ????

Drummond:1979:TSP

[Dru79] Douglas J. Drummond. TSC-SREG: a SAS procedure for the analysis of time series crosssection data. SAS technical report; S-106. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1979. 55 pp. Includes bibliographies.

Delwiche:1995:LSB

[DS95] Lora D. Delwiche and Susan J.

Slaughter. The little SAS book: a primer. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-215-3. ix + 228 pp. LCCN QA276.4.D45 1995.

Delwiche:1998:LSB

[DS98a] Lora D. Delwiche and Susan J.

Slaughter. The Little SAS Book: a primer. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1998. ISBN 1-58025-239-7. xiii + 267 pp. LCCN ????

DuarteSilva:1998:NTG

[DS98b] A. Pedro Duarte Silva and Antonie Stam. Nonparametric two-group classification: Concepts and a SAS-based software package. The American Statistician, 52(2):183–??

(or 185–??), May 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts_98/silva.html.

daSilva:2015:SMA

[dSM15] Alan Ricardo da Silva and Iracema Veiga Madeira Mauriz. A SAS macro for adaptive spatial sampling. International Journal of Statistics and Probability, 4(4):20–??, ???? 2015. CODEN ???? ISSN 1927-7032 (print), 1927-7040 (elec-

tronic). URL http://www.
ccsenet.org/journal/index.
php/ijsp/article/view/51693.

Dudewicz:1986:MEP

[Dud86] Edward J. Dudewicz. Modern elementary probability and statistics with SAS programming: student manual for Mathematics 127, Mathematics 128. ????, ????, second edition, 1986. ISBN 0-935950-12-5 (paperback). various pp. LCCN QA276.12 .D82 1986.

Durkalski:2006:BRB

[Dur06] Valerie L. Durkalski. Book review: Analysis of Clinical Trials Using SAS(R): A Practical Guide, by Alex Dmitrienko; Geert Molenberghs; Christy Chuang-Stein; Walter Offen. Journal of the American Statistical Association, 101(474): 858, June 2006. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/
stable/27590761.

Villasenor:1984:USG

[dV84] Yolanda Fernandez de Villaseñor. Uso de SAS/GRAPH para produccion de mapas de la Republica Mexicana. Centro de Estadistica y Calculo, Chapingo, Mexico, 1984. ISBN 968-839-006-5 (??invalid ISBN??). 30 pp. LCCN GA102.4.E4 V55 1984.

Easter:1987:SGP

[Eas87] Gretel Easter. SAS/GRAPH guide for personal computers, version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1987. ISBN 1-55544-068-1. xviii + 534 pp. LCCN QA 276.4 S27 1987. Edited by Gretel Easter. Includes bibliographies and index.

Everitt:1996:HSA

[ED96] Brian Everitt and Geoff Der. A handbook of statistical analyses using SAS. Chapman and Hall, Ltd., London, UK, 1996. ISBN 0-412-71050-1 (paperback). vi + 158 pp. LCCN QA276.4 .E88 1996.

Easter:1986:SUG

[EK86] Gretel Easter and Mike Kalt.

SAS/RTERM user’s guide, version 1. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 0-917382-99-4. x + 50 pp. LCCN QA276.4 S195 1986. Includes bibliographies and index.

Elliott:1995:LSC

[Ell95] Rebecca J. Elliott. Learning SAS in the computer lab. Duxbury Press, Belmont, CA, USA, 1995. ISBN 0-534-23442-9. x + 175 pp. LCCN QA276.4 .E44 1995.

Elliott:2000:LSC

[Ell00] Rebecca J. Elliott. Learning SAS in the computer lab. Duxbury Press, Belmont, CA, USA, second edition, 2000. ISBN 0-534-35925-6. xi + 236 pp. LCCN QA276.4 .E44 1999.

Elliott:2009:LSC

[EM09] Rebecca J. Elliott and Christopher H. Morrell. Learning SAS in the Computer Lab. Duxbury Press, Belmont, CA, USA, third edition, 2009. ISBN 0-495-55968-7. 298 (est.) pp. LCCN ????

Elliott:2010:LSC

[EM10] Rebecca J. Elliott and Christopher H. Morrell. Learning SAS in the computer lab. Brooks/Cole Cengage Learning, Boston, MA, USA, third edition, 2010. ISBN 0-495-55968-7. xi + 298 pp. LCCN QA276.4 .E44 2010.

Evans:1988:CSP

[EN88] Gerald Evans and William A. Newman. A comparison of SPSS PC, SAS PC, and BMDP PC. Collegiate Microcomputer, 6(2):97–106, May 1988. CO-DEN CMICDL. ISSN 0731-4213.

Enright:1994:ACA

[EWC94] Aaron Garth Enright, Linda M.

Wilkens, and James T. Canning. An alternative computer architecture course. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(4):9–12, December 1994. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Fan:1994:SSD

[Fan94] Yihua Fan. A systemization of statistical data management/ manipulation tasks and a comparison of statistical data management/manipulation capabilities of SAS and SPSS base programs. Thesis (ph. d.), University of Virginia, Charlottesville, VA, USA, 1994. xii + 130 pp.

Fisher:2020:USM

[FCY20] Paige Fisher, Wentao Cao, and Qingzhao Yu. Using SAS macros for multiple mediation analysis in R. Journal of Open Research Software, 8(1):22–??, October 07, 2020. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
277/.

Fernandez:2003:DMU

[Fer03] George Fernandez. Data mining using SAS applications.

Chapman and Hall/CRC, Boca Raton, FL, USA, 2003. ISBN ???? xv + 367 pp. LCCN HF1017 .F476 2003.

Freund:1981:SLM

[FL81] Rudolf Jakob Freund and Ramon C. Littell. SAS for linear models: a guide to the ANOVA and GLM procedures. SAS series in statistical applications. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. ISBN 0-917382-31-5. 231 pp. LCCN QA279 .F7. Includes index.

Freund:1991:SSR

[FL91] Rudolf Jakob Freund and Ramon C. Littell. SAS System for regression. SAS series in statistical applications. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1991. ISBN 1-55544-429-6. xvi + 210 pp. LCCN QA276.4 .F74 1991. “56141”–Cover. Includes bibliographical references (p. 203-204) and index.

Freund:1986:SSL

[FLS86] Rudolf Jakob Freund, Ramon C. Littell, and Philip C. Spector. SAS system for linear models. SAS series in statistical applications. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 1-55544-028-2 (paperback). x + 210 pp. LCCN QA76.6.G4S32417 1986. Includes index.

Fitzmaurice:2011:ALA

[FLW11] Garrett M. Fitzmaurice, Nan M.

Laird, and James H. Ware. Applied longitudinal analysis. Wiley series in probability and statistics. Wiley-Blackwell, Oxford, UK, second edition, 2011. ISBN 0-470-38027-6 (hardcover), 1-118-55179-6. xxv + 701 pp. LCCN ????

Field:2012:DSU

[FMF12] Andy P. Field, Jeremy Miles, and Zoë Field. Discovering statistics using R. Sage, London, UK, 2012. ISBN 1-4462-0045-0 (hardcover), 1-4462-0046-9. xxxiv + 957 pp. LCCN HA32 .F537 2012.

Falk:2002:FSA

[FMT02] Michael Falk, Frank Marohn, and Bernward Tewes. Foundations of statistical analyses and applications with SAS. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2002. ISBN 3-7643-6893-4, 0-8176-6893-4. ix + 399 pp. LCCN QA276.4 .F3513 2002.

Fowler:1993:SSG

[Fow93] Linda K. Fowler. The SAS system guide. College custom series. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-020189-7. 20 pp. LCCN ????

Frane:1979:PII

[Fra79] James W. Frane. Part I–The 1978 interface between BMDP-77 and SAS and some BMDP features of interest to SAS users. Part II–Recent developments in BMDP of special interest to SAS users. Technical report / BMDP Statistical Software; no. 40. BMDP Statistical Software, Department of Biomathematics, University of California, Los Angeles, CA, USA, 1979. 4 + [2] + 8 pp. Includes bibliography.

Francis:1981:SSC

[Fra81] Ivor Francis. Statistical software: a comparative review. Elsevier North-Holland, Inc., New York, NY, USA, 1981. ISBN 0-444-00658-3. xx + 542 pp. LCCN QA276.4 .F7 1981. US$70.

Fragkos:2014:AMS

[Fra14] Konstantinos C. Fragkos. Applied medical statistics using SAS. Journal of Applied Statistics, 41(8):1857–1858, August 2014. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Friendly:1991:SSS

[Fri91a] Michael Friendly. SAS System for Statistical Graphics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-441-5. xviii + 697 pp. LCCN QA276.4 .F75 1991. URL http:/

/www.sas.com/service/doc/
pubcat/uspubcat/ind_files/
56143.html.

Friesen:1991:CGS

[Fri91b]  Blake Friesen.  A Concise

Guide to Statistical Software Packages: An Introduction to Minitab, SPSS/PC+, SAS, NWA Statpak, and Lotus 1-2-3. Harcourt Brace Jovanovich Canada, Toronto, Ontario, Canada, 1991. ISBN 0-03-997801-X. 87 pp.

Friendly:2006:DEP

[Fri06] Michael Friendly. Data ellipses, HE plots and reduced-rank displays for multivariate linear models: SAS software and examples. Journal of Statistical Software, 17(6):1–43, November 2006. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v17/i06.

Fontana:2013:MSM

[FS13] Roberto Fontana and Sabrina Sampò. Minimum-size mixedlevel orthogonal fractional factorial designs generation: A SAS-based algorithm. Journal of Statistical Software, 53 (10):??, May 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v53/i10.

Fernandez:2002:SSA

[FTP02] C. Fernández, F. Torres, and S. T. Puente. SASEPA: Simultaneous allocation and scheduling with exclusion and precedence relations algorithm. Lecture Notes in Computer Science, 2328:65–??, 2002.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/
service/series/0558/bibs/
2328/23280065.htm;   http:
//link.springer-ny.com/link/
service/series/0558/papers/
2328/23280065.pdf.

Feng:2009:ATD

[FZZZ09] Rui Feng, Gongfu Zhou, Meizhuo Zhang, and Heping Zhang. Analysis of twin data using SAS. Biometrics, 65 (2):584–589, June 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilmour-Bryson:1982:CDT

[GB82] A. Gilmour-Bryson. Coding of the depositions of the Templars. In Ciampi and Martino [CM82], pages 451–467. ISBN 0-444-86413-X (set), 0-444-86414-8 (vol. 1), 0-444-86415-6 (vol. 2). LCCN K662.I4 I58. From Computing Reviews: “The article reports on the progress of an historical study using a computer for statistical analysis. The subject of the study is a mass of trial depositions of the Knights Templar. The purpose of the project is to find statistical regularities in the large amount of testimony and to create a model for similar studies. The Order of the Knights Templar, founded around 1100, was the first Christian military order.

After the Crusades, in the early fourteenth century, 127 articles of accusation were brought against the order, including charges of idolatry, sacrilege, and sodomy. The data being studied are the responses of 900 men to each of 127 accusations. The article details the accusations and the way in which the responses are being coded. The statistical package SAS (Statistical Analysis System) will be used. Examples of the results to be sought are: tallies of guilty/innocent responses, tallies of offenses committed versus seen versus heard about, and correlations between, for example, age and other responses. Depositions that differ markedly from the average response will be identified.”.

Gentle:2011:BRS

[Gen11] James E. Gentle. Book review: A SAS/IML Companion for Linear Models, by Perrett, J. J. Biometrics, 67(1): 327–328, March 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gautier:1982:SPL

[GG82] J.-M. Gautier and J.-M. Grosbras. SAS-PAD program library for data analysis integrated with SAS software. In Caussinus et al. [CEM82], pages 107–108. ISBN 3-7051-0002-5 (vol. 1), 3-7051-0001-7 (vol. 2). LCCN QA276.4 .C194 1982. Two volumes.

Gomez:2023:CUP

[GGS23] Camilo Gomez, Alexander V. Goponenko, and Julia N. Soulakova. Constructing Up-Set plot for survey data with weights using SAS and R software. Communications in Statistics: Simulation and Computation, 52(6):2320–2326, 2023. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2010:BRBb

[Gho10] Subir Ghosh. Book review: Analyzing Receiver Operating Characteristic Curves With SAS by Mithat Gonena. Technometrics, 52(3):365, August 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/27867264.

Garcia-Hernandez:2018:JSM

[GHR18] Alberto Garcia-Hernandez and Dimitris Rizopoulos. %JM: A SAS macro to fit jointly generalized mixed models for longitudinal data and time-to-event responses. Journal of Statistical Software, 84(??):??, ???? 2018. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

                                        jstatsoft.org/index.php/
                                        jss/article/view/v084i12;
                                         https://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v084i12/v84i12.pdf.

Graybill:1994:SLM

[GI94] Franklin A. Graybill and Hariharan K. Iyer. SAS laboratory manual to accompany Regression analysis: concepts and applications. Duxbury Press, Belmont, CA, USA, 1994. vii + 293 pp.

Giacobazzi:2004:SAI

[Gia04] Roberto Giacobazzi, editor.

Static analysis: 11th international symposium, SAS 2004, Verona, Italy, August 26– 28, 2004. proceedings, volume 3148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2004. CODEN LNCSD9. ISBN 3-540-22791-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.
springerlink.com/content/
978-3-540-22791-5;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3148;   http:/
/www.springerlink.com/openurl.
asp?genre=volume&id=doi:
10.1007/b99688.

Giesbrecht:2011:BRD

[Gie11] Francis Giesbrecht. Book review: Design and Analysis of Experiments with SAS, by Lawson, J. Biometrics, 67(4):1676– 1677, December 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gilmore:1996:PWB

[Gil96] Jodie Gilmore. Painless Windows 3.1: a beginner’s handbook for SAS users. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN 1-55544-878-X. xiv + 290 pp. LCCN QA76.76.W56 G555 1996.

Gilmore:1999:PWH

[Gil99] Jodie Gilmore. Painless Windows: a Handbook for SAS Users. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1999. ISBN 1-58025-238-9. xvi + 332 pp.

Gilmour:2012:JLD

[Gil12] Steven G. Gilmour. John Lawson: Design and analysis of experiments with SAS. Statistics and Computing, 22 (4):995–996, July 2012. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

                                        URL http://link.springer.
                                        com/accesspage/article/10.
                                        1007/s11222-011-9238-z.

Goodman:1976:IDP

[GL76] Hortense Goodman and Leonard Lorensen. Illustrations of the disclosure of “pro forma” calculations: a survey of the application of certain sections of APB opinion nos. 15, 16, and 20, and SAS no. 1. Financial report survey; 11. American Institute of Certified Public Accountants, New York, NY, USA, 1976. 100 pp. LCCN HF5681 .B2G66. Includes index.

Goodman:1979:IAR

[GL79] Hortense Goodman and Leonard Lorensen. Illustrations of auditor’s reports on comparative financial statements: a survey of the application of SAS no.

15. Financial report survey; 18. American Institute of Certified Public Accountants, New York, NY, USA, 1979. 130 pp. LCCN HF5667 .G628. Includes index.

Godfrey:2014:SRS

[GL14] A. Jonathan R. Godfrey and M. Theodor Loots. Software review: Statistical Software (R, SAS, SPSS, and Minitab) for Blind Students and Practitioners. Journal of Statistical Software, 58(SR-1):??, July 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v58/s01.

Gosho:2014:CSC

[Gos14] Masahiko Gosho. Code snippet: Criteria to select a working correlation structure in SAS. Journal of Statistical Software, 57(CS-1):??, March 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v57/c01.

Gottsche:1992:SKV

[Göt92] Thomas Göttsche. SAS kompakt für die Version 6. (German) [SAS compact for Version 6]. G. Fischer, Stuttgart, Germany, 1992. ISBN 3-437-40281-1. 192 pp. LCCN

                                        ????  URL http://d-nb.
                                        info/920401988/04;   http:
                                        //www.gbv.de/dms/hebis-darmstadt/
                                        toc/19748795.pdf.

Grant:1996:SPT

[Gra96] Paul Grant, editor. SUGI 21: proceedings of the Twenty-First Annual SAS Users Group International Conference, Chicago, Illinois, March 10–13, 1996, SUGI 1996; CONF 21. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN 1-55544-747-3. LCCN ????

Greenberg:1981:CSA

[Gre81] Edward A. Greenberg. Comparison of some available packages for use in research data management: [SIR, SAS, and SPSS]. ACM Press, New York, NY 10036, USA, 1981. 7 pp. Offprint.

Greene:2017:CSS

[Gre17] Erich J. Greene. Code snippet: A SAS macro for covariateconstrained randomization of general cluster-randomized and unstratified designs. Journal of Statistical Software, 77 (??):??, ???? 2017. CO-DEN JSSOBK. ISSN 1548-

                                        7660.  URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v077c01;
                                         https://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v077c01/v77c01.pdf.

Griffith:1993:SRA

[Gri93] Daniel A. Griffith. Spatial regression analysis on the PC: spatial statistics using SAS. Resource publications in geography. Association of American Geographers, Washington, DC, USA, 1993. ISBN 0-89291-213-8. xi + 130 pp. LCCN G70.3 .G747 1993.

Gromping:2010:BRB

[Grö10] Ulrike Grömping. Book review:

Design and Analysis of Experiments with SAS. Journal of Statistical Software, 37(BR-2): ??, December 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v37/b02.

Govindarajulu:2004:SMC

[GSD04] Usha Govindarajulu, Lisa Sullivan, and Ralph B. D’Agostino. SAS macros for confidence interval estimation for CHD risk prediction by different survival models. Computational Statistics & Data Analysis, 46(3):571–592, June 15, 2004. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947303001750.

Gunst:2003:BRB

[Gun03] Richard F. Gunst. Book review:

Regression and ANOVA: An Integrated Approach Using SAS Software by Keith E. Muller; Bethel A. Fetterman. Technometrics, 45(2):170–171, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047015.

Gupta:2002:BRB

[Gup02a] Ghanshyam Gupta. Book review: A Handbook of Statistical Analyses using SAS, 2nd ed. The American Statistician, 56(4):334–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s22/p334.

Gupta:2002:BRH

[Gup02b] Ghanshyam Gupta. Book review: A Handbook of Statistical Analyses using SAS, 2nd ed. The American Statistician, 56(4):334–??, November 2002. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://oberon.
ingentaselect.com/cgi-bin/
linker?ini=asa&reqidx=/cw/
asa/00031305/v56n4/s22/p334.

Gurwit:2008:LSE

[Gur08] Margie R. Gurwit. Learning SAS by example: a programmer’s guide. The American Statistician, 62(2):182–183, May 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hackl:1984:BRD

[Hac84] P. Hackl. Book review: Datenverarbeitung und Datenanalyse mit SAS: Eine problemorientierte Einführung, by H.-M. Uehlinger. Biometrics, 40(2): 572–573, June 1984. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/2531430.

Hall:2008:ARO

[Hal08] Daniel B. Hall. Analyzing receiver operating characteristic curves with SAS. The American Statistician, 62(4): 362, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Hardison:1981:SFC

[Har81] C. David Hardison. SAS functions to compute chi-squared, F, and t distributions and their inverses, allowing for noncentrality and fractional degrees of freedom. SAS technical report; S-122. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. 7 pp. Includes bibliography.

Hardin:1993:STF

[Har93] Reid Hardin. SAS/TUTOR:

fundamentals of the SAS system course guide, OS/2 and Windows (Release 6.08), Open-VMS and UNIX (X Window Environments), Windows NT (Release 6.09). SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. vi + 139 pp.

Hartmann:2008:BRB

[Har08] Wolfgang M. Hartmann. Book review: SAS for Data Analysis: Intermediate Statistical Methods. Journal of Statistical Software, 28(BR-1):??, October 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v28/b01.

Hatcher:1994:SAUa

[Hat94] Larry Hatcher. A step-by-step approach to using the SAS system for factor analysis and structural equation modeling. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-643-4. xiv + 588 pp. LCCN ????

Heckler:1990:BRB

[Hec90] Charles E. Heckler. Book review: Mastering the SAS System by Jay A. Jaffe. Technometrics, 32(3):350–351, August 1990. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/1269125.

Heckler:1996:BRB

[Hec96] Charles E. Heckler. Book review: A Step-by-Step Approach to Using the SAS(R) System for Factor Analysis and Structural Equation Modeling by Larry Hatcher. Technometrics, 38(3): 296–297, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270628.

Heckler:2011:BRB

[Hec11] Charles E. Heckler. Book review: SAS and R: Data Management, Statistical Analysis and Graphics by Ken Kleinman; Nicholas J. Horton. Technometrics, 53(2):221, May 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/23209749.

Heitkamp:1990:BRB

[Hei90] Norman Heitkamp. Book review: SAS System for Elementary Statistical Analysis by Sandra D. Schlotzhauer; Ramon C. Littell; Cary. Technometrics, 32(2):230–231, May 1990. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1268879.

Helwig:1977:SSL

[Hel77] Jane T. Helwig. SAS supplemental library user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1977. ISBN 0-917382-03-3 (??invalid checksum??). 171 pp. LCCN QA276.4 .S234 1977. Includes bibliographical references and index.

Helwig:1978:SIG

[Hel78] Jane T. Helwig. SAS introductory guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1978. ISBN 0-917382-05-3 (??invalid checksum??). 83 pp. LCCN QA278 .H45. Includes index.

Helwig:1983:SIG

[Hel83] Jane T. Helwig. SAS introductory guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, revised edition, 1983. ISBN 0-917382-43-9. 93 pp. LCCN QA 276.4 H44 1983. Includes index.

Helwig:1985:SIG

[Hel85] Jane T. Helwig. SAS introductory guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1985. ISBN 0-917382-73-0. 99 pp. LCCN QA276.4 .S561 1985. Includes index.

Herzberg:1990:HSW

[Her90] Paul A. Herzberg. How SAS works: a comprehensive introduction to the SAS system. Springer-Verlag Inc., New York, NY, USA, second edition, 1990. ISBN 0-387-97291-9 (Springer-Verlag New York), 3-540-97291-9 (Springer-Verlag Heidelberg), 0-921801-28-9 (Captus Press). xvii + 193 pp. LCCN QA276.4 .H47 1990. Includes index.

Heiberger:2004:SAD

[HH04] Richard M. Heiberger and Burt Holland. Statistical analysis and data display: an intermediate course with examples in S-plus, R, and SAS. Springer texts in statistics. Springer-Verlag Inc., New York, NY, USA, 2004. ISBN 0-387-40270-5. xxiv + 729 pp. LCCN QA276.4 .H445 2004.

Heiberger:2015:SAD

[HH15] Richard M. Heiberger and Burt Holland. Statistical Analysis and Data Display: an Intermediate Course with Examples in R. Springer texts in statistics. Springer-Verlag Inc., New York, NY, USA, second edition, 2015. ISBN 1-4939-2121-5 (hardcover), 1-4939-2122-3 (ebook). xxxi + 898 pp. LCCN QA276.4 .H445 2015.

Hill:1993:LSC

[Hil93]  R. Carter Hill.  Learning

SAS: a computer handbook for econometrics: a guide to programming for Griffiths/Hill/ Judge, Learning and practicing econometrics, and Judge/ Hill/Lutkepohl/Lee, Introduction to the theory and practice of econometrics, 2nd ed. John Wiley and Sons, Inc., New York, NY, USA, 1993. ISBN 0-471-58553-X (paperback). 388 pp. LCCN HB139.H547 1993.

Hilbe:2014:MCD

[Hil14] Joseph M. Hilbe. Modeling Count Data. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-02833-7 (hardcover), 1-107-61125-3 (paperback). xv + 283 pp. LCCN QA278 .H56 2014.

Hohn:1985:SPQ

[Hoh85] M. E. Hohn. SAS program for quantitative stratigraphic correlation by principal components. Computers and Geosciences, 11(4):471–477, 1985. CODEN CGOSDN. ISSN 0098-3004.

Holt:2008:BRB

[Hol08] Melinda Holt. Book review: Learning SAS by Example: A Programmer’s Guide by Ron Cody. Technometrics, 50(1): 91–92, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/25471436.

Hornikova:2009:BRBa

[Hor09a] Adriana Horníková. Book review: Optimum Experimental Designs, With SAS by A. C. Atkinson; A. N. Donev; R. D. Tobias. Technometrics, 51(3): 341–342, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/40586631.

Hornikova:2009:BRBc

[Hor09b] Adriana Horníková. Book review: The Little SAS Book: A Primer (4th ed.) by Lora D. Delwiche; Susan J. Slaughter. Technometrics, 51(4): 485, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/40586666.

Howard:1988:CAC

[How88] John A. Howard. Computer applications for consumer behavior in marketing strategy. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-169731-5. iii + 212 pp. LCCN HF5415.2 .H69 1988.

Hoyle:1996:SSW

[Hoy96] L. Hoyle. SAS software and the WWW — what next? In Anonymous [Ano96a], pages 867–874 (vol. 1). Two volumes.

Holzer:1992:MCP

[HP92] Christian Holzer and Manfred Precht. Multiple comparison procedures for normally distributed ANOVA models in SAS, SPSS, BMDP, and MINITAB. Computational Statistics & Data Analysis, 13(3):351–358, April 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0167947392901412.

Hermenegildo:2002:SAI

[HP02] Manuel Hermenegildo and German Puebla, editors. Static Analysis: 9th International Symposium, SAS 2002, Madrid, Spain, September 17–20, 2002. Proceedings, volume 2477 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2002. CODEN LNCSD9. ISBN 3-540-44235-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 2002. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t2477.htm;   http://www.
springerlink.com/content/
978-3-540-44235-6;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=2477.

Hammill:2006:SIS

[HP06] Bradley G. Hammill and John S. Preisser. A SAS/ IML software program for GEE and regression diagnostics. Computational Statistics & Data Analysis, 51 (2):1197–1212, November 15, 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947305003099.

Helwig:1979:SPG

[HR79] Jane T. Helwig and Patti S. Reinhardt. SAS programmer’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1979. 184 pp. Includes index.

Hatcher:1994:SAUb

[HS94] Larry Hatcher and Edward J. Stepanski. A step-by-step approach to using the SAS system for univariate and multivariate statistics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-634-5. xiv + 552 pp. LCCN ????

Hankin:2005:SAI

[HS05] Chris Hankin and Igor Siveroni, editors. Static analysis: 12th international symposium, SAS 2005, London, UK, September 7–9, 2005, proceedings, volume 3672 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2005. CODEN LNCSD9. ISBN 3-540-28584-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN

????  URL http://www.
springerlink.com/content/
978-3-540-28584-7;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=3672.

Hsu:1986:MPL

[Hsu86] Jeffrey Hsu. Microcomputer programming languages. John Wiley and Sons, Inc., New York, NY, USA, 1986. ISBN 0-471-62574-4. 404 pp. LCCN QA76.7.H78 1987. US$16.95. From Computing Reviews: “Finally, two ‘statistical languages’ are covered: SPSS and SAS.”.

Hu:2009:BRA

[Hu09] Feifang Hu. Book review:

Adaptive Design Theory and Implementation Using SAS and R, by Chang, M. Biometrics, 65(2):668–669, June 2009. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hudak:1989:CEA

[Hud89]  Paul Hudak.  Conception,

evolution, and application of functional programming languages. ACM Computing Surveys, 21(3):359–411, September 1989. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0360-
                                        0300/72554.html.

Hutchinson:1992:SUG

[Hut92] E. Jeffrey Hutchinson, editor. SAS Users Group International. Conference (17th: 1992: Honolulu, Hawaii) SUGI 17: proceedings of the Seventeenth Annual SAS Users Group International Conference, Honolulu, Hawaii, April 12–15, 1992. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-505-5. LCCN QA276.4 .S2397 1992.

Ingwersen:1981:SAU

[Ing81] James B. Ingwersen. Statistical analysis using SAS at the USEPA National Computer Center. USDA Forest Service, 1981. 51 pp. Cover title.

Jaffe:1989:MSS

[Jaf89] Jay Jaffe. Mastering the SAS system. Van Nostrand Reinhold Co., New York, NY, USA, 1989. ISBN 0-442-24015-5. xv + 629 pp. LCCN QA276.4 .J341 1989. US$44.95. Includes index.

Jaffe:1994:MSS

[Jaf94] Jay Jaffe. Mastering the SAS system. VNR computer library. Van Nostrand Reinhold Co., New York, NY, USA, second edition, 1994. ISBN 0-442-01295-0. x + 574 pp. LCCN QA276.4 .J341 1994. US$32.95. Includes index.

Jaffe:1996:MSS

[Jaf96] J. Jaffe. Mastering the SAS System. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, 1996. ISBN 1-85032-831-5. ???? pp. LCCN ???? US$39.95.

Jensen:2008:BRB

[Jen08] Willis A. Jensen. Book review: Decision Trees for Business Intelligence and Data Mining: Using SAS(R) Enterprise Miner^{T M} by Barry De Ville. Technometrics, 50(3): 409–410, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25471505.

Jensen:2009:BRBb

[Jen09] Willis A. Jensen. Book review:

Data Mining Using SAS(R) Enterprise Miner^{T M} by R. Matignon. Technometrics, 51 (1):105, February 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/40586577.

James:1991:PSU

[JLC91] Teresa L. James, Jane P. Loftis, and Tai-Lun Chiang. The program specifications for the user interface using SAS software for the Worldwide Household Goods Information System for Transportation Modernization (WHIST-MOD). Oak Ridge, TN, USA, 1991. 135 pp. Cover title. Microfiche. Springfield, Va.: National Technical Information Service.

Johnson:2008:SIS

[JMD08] Kjell Johnson, Abhyuday Mandal, and Tan Ding. Software for implementing the sequential elimination of level combinations algorithm. Journal of Statistical Software, 25(6):1–13, March 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v25/i06.

Johnson:2009:PSI

[JMS^{+}09] Jacqueline L. Johnson, Keith E. Muller, James C. Slaughter, Matthew J. Gurka, Matthew J. Gribbin, and Sean L. Simpson. POWERLIB: SAS/IML software for computing power in multivariate linear models. Journal of Statistical Software, 30(5):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v30/i05.

Jones:2007:BRA

[Jon07]  Michael Jones.  Book re-

view: 1. Analysis of clinical trials using SAS(R): a practical guide. Alex Dmitrienko, Geert Molenberghs, Christy Chuang-Stein and Walter Offen, SAS Press, Cary, 2005. No. of pages: vii + 419. Price: $69.95. ISBN 1-59047-504-6. Statistics in Medicine, 26(12):2587– 2588, May 30, 2007. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joyner:1985:SUG

[Joy85] Stephenie P. Joyner. SAS/ Graph User’s Guide: Version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-68-4. xviii + 596 pp. LCCN QA 276.4 S22 1985. Includes index.

Jasen:2013:BRB

[JS13] Kelcey Jasen and Luigi Salmaso.

Book review: Applied Medical Statistics Using SAS. Journal of Statistical Software, 52(BR-2):??, February 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v52/b02.

Juretig:2015:CSP

[Jur15] Francisco Juretig. Code snippet: MLGA: A SAS macro to compute maximum likelihood estimators via genetic algorithms. Journal of Statistical Software, 66(CS-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/index.php/
jss/article/view/v066c02;
 http://www.jstatsoft.org/

                                        index.php/jss/article/view/
                                        v066c02/v66c02.pdf.

Jampani:2011:MCD

[JXW^{+}11] Ravi Jampani, Fei Xu, Mingxi Wu, Luis Perez, Chris Jermaine, and Peter J. Haas. The Monte Carlo database system: Stochastic analysis close to the data. ACM Transactions on Database Systems, 36(3):18:1– 18:??, August 2011. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

Knoke:2005:ARM

[KAK05] James D. Knoke, Christy M. Anderson, and Gary G. Koch. Analyzing repeated measures marginal models on sample surveys with resampling methods. Journal of Statistical Software, 15(8):1–13, December 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v15/i08.

Kalt:1986:SGM

[Kal86] Brenda C. Kalt. SAS guide to macro processing: version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 1-55544-039-8 (??duplicate ISBN??). xvi + 300 pp. LCCN QA276.4 .K35 1986.

Karjala:1987:ACP

[Kar87a] E. Karjala. APL and SAS in co-operation. In Leroy J. Dickey and Lynne C. Shaw, editors, APL 88 Conference proceedings: APL, its past, its present, its future: 1–5 February 1988, Sydney, Australia, Members US$21.00, pages 202– 208. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-253-5. LCCN QA 76.73 A27 A195 v.18 no.2 1988. US$28.00.

Karjala:1987:ASC

[Kar87b] E. Karjala. APL and SAS in co-operation. In Leroy J. Dickey and Lynne C. Shaw, editors, APL 88 Conference proceedings: APL, its past, its present, its future: 1–5 February 1988, Sydney, Australia, Members US$21.00, pages 202– 208. ACM Press, New York, NY 10036, USA, 1987. ISBN 0-89791-253-5. LCCN QA 76.73 A27 A195 v.18 no.2 1988. US$28.00.

Karlsson:2007:BRS

[Kar07] Andreas Karlsson. Book review: A SAS Companion for Nonparametric Statistics, by Scott J. Richter and James J. Higgins. Computational Statistics, 22(1):187–188, April 2007. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/s00180-007-0032-5.

Karagrigoriou:2008:BRB

[Kar08] Alex Karagrigoriou. Book review: Analysing Receiver Operating Characteristic Curves with SAS, by M. Gönen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(4):1036, October 2008. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/30130797.

Karagrigoriou:2009:BRB

[Kar09]  Alex Karagrigoriou.  Book

review: Basic Statistics Using SAS Enterprise Guide, a Primer, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172(2): 530–531, April 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/
                                        stable/20622514.

Kastner:1998:BRB

[Kas98a] William Kastner. Book review: Statistical Quality Control Using the SAS System. The American Statistician, 52 (1):85, February 1998. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/2685576.

Kastner:1998:SQC

[Kas98b] William Kastner. Statistical quality control using the SAS system. The American Statistician, 52(1):85–??, February 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Kolosova:1995:TSR

[KB95] Tanya Kolosova and Samuel Berestizhevsky. Table-driven strategies for rapid SAS applications development. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-245-5. viii + 259 pp. LCCN ????

Kairalla:2008:GSI

[KCM08] John A. Kairalla, Christopher S. Coffey, and Keith E. Muller. GLUMIP 2.0: SAS/ IML software for planning internal pilots. Journal of Statistical Software, 28(7):??, November 2008. CODEN JS-SOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/
v28/i07.

Ke:2011:BRB

[Ke11a] Weiming Ke. Book review: SAS for Data Analysis — Intermediate Statistical Methods by Mervyn Marasinghe; William Kennedy. Technometrics, 53 (1):99–100, February 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/40997298.

Ke:2011:PSE

[Ke11b] Weiming Ke. PROC SQL by example — using SQL within SAS by Howard Schreier. Technometrics, 53(1):105, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40997307.

Kelly:2010:BRA

[Kel10] Patrick Kelly. Book reviews:

Adaptive design theory and implementation using SAS and R. Mark Chang, Chapman and Hall/CRC, Boca Raton, 2008. No. of pages: xxii + 418. Price: $89.95. ISBN 10: 1-58488-962-4, ISBN 13: 978-1-58488-962-5. Statistics in Medicine, 29 (9):1058, April 30, 2010. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kodell:1984:CST

[KGHS84] Ralph L. Kodell, David W. Gaylor, M. Gary Haskin, and Gary W. Shaw. CHRONIC: An SAS[registered trademark] procedure for statistical analysis of carcinogenesis studies. The American Statistician, 38 (2):158, May 1984. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2683262.

Kleinman:2010:SRD

[KH10] Ken Kleinman and Nicholas J. Horton. SAS and R: Data Management, Statistical Analysis, and Graphics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4200-7057-6 (hard back). xix + 323 pp. LCCN QA76.73.S27 K54 2010.

Khademi:2015:BRSb

[Kha15] Abdolvahab Khademi. Book review: Statistical Hypothesis Testing with SAS and R, Dirk Taeger and Sonja Kuhnt, John Wiley & Sons, Chichester, 2014. ISBN 978-1-119-95021-9. xiv + 292 pp. USD 97.95 (H).

http://www.wiley.com/WileyCDA/WileyTitle/productCd-1118762614.html.

Journal of Statistical Software, 68(BR-2):??, ???? 2015. CO-DEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/index.php/
jss/article/view/v068b02;
 http://www.jstatsoft.org/
index.php/jss/article/view/
v068b02/v68b02.pdf.

Khademi:2018:BRCa

[Kha18] Abdolvahab Khademi. Book review: Complex Survey Data Analysis with SAS. Journal of Statistical Software, 84 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

7660.  URL https://www.
jstatsoft.org/index.php/
jss/article/view/v084b02;
 https://www.jstatsoft.org/
index.php/jss/article/view/
v084b02/v84b02.pdf.

Khattree:2020:BRE

[Kha20] Ravi Khattree. Book review:

End-to-End Data Science with SAS: A Hands-On Programming Guide. James Gearheart SAS Press, 2020, xiii + 363 pages, £23.99/$29.99, e-book ISBN: 978-1-64295-806-5. International Statistical Review = Revue Internationale de Statistique, 88(3):809–812, December 2020. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kodell:1983:CSP

[KHSG83] R. L. Kodell, M. G. Haskin, G. W. Shaws, and D. W. Gaylor. Chronic: a SAS procedure for statistical analysis of carcinogenesis studies. Journal of Statistical Computation and Simulation, 16(3/4): 287–310, January 1983. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1563-5163 (electronic).

Khuri:2013:SBRa

[Khu13] André I. Khuri. Short book review: SAS Statistics by Example by Ron Cody. International Statistical Review = Revue Internationale de Statistique, 81 (1):164–165, April 2013. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Kuss:1996:MAM

[KK96] Oliver Kuss and Armin Koch. Meta-analysis macros for SAS. Computational Statistics & Data Analysis, 22(3):325– 333, July 15, 1996. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0167947396814563.

Kalema:2016:CSG

[KM16] George Kalema and Geert Molenberghs. Code snippet: Generating correlated and/or overdispersed count data: A SAS implementation. Journal of Statistical Software, 70 (??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660.  URL https://www.
jstatsoft.org/index.php/
jss/article/view/v070c01;
https://www.jstatsoft.org/
index.php/jss/article/view/
v070c01/v70c01.pdf.

Khattree:1995:AMS

[KN95] Ravindra Khattree and Dayanand N.

Naik. Applied multivariate statistics with SAS software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-239-0. xiv + 396 pp. LCCN ????

Khattree:1999:AMS

[KN99] Ravindra Khattree and Dayanand N.

Naik. Applied multivariate statistics with SAS software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1999. ISBN 1-58025-357-1. xiv + 338 pp. LCCN MLCM 99/2921 (Q).

Kim:2010:PQS

[KS10] Hanjoo Kim and Justine Shults.

%QLS SAS macro: a SAS macro for analysis of correlated data using quasi-least squares. Journal of Statistical Software, 35 (2):??, July 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v35/i02.

Kuffner:1985:DSA

[KW85] Helmuth Kuffner and Reinhard Wittenberg. Datenanalysessyteme fur statistische Auswertungen: e. Einführung in SPSS, BMDP und SAS. (German) [Data Analysis Systems for Statistical Evaluations: An Introduction to SPSS, BMDP, and SAS]. G. Fischer, Stuttgart, Germany, 1985. 292 pp.

Lachenbruch:1999:BRL

[Lac99] P. A. Lachenbruch. Book review: Linear Mixed Models in Practice: a SAS-Oriented Approach, by G. Verbeke, G. Molenberghs. Biometrics, 55 (1):327, March 1999. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (elec-

tronic). URL https://www.
jstor.org/stable/2533932.

Laforge:1981:AMP

[Laf81] Hubert Laforge. Analyse multivariée: pour les sciences sociales et biologiques avec applications des logiciels BMD, BMDP, SPSS, SAS. Etudes vivantes, Saint-Lambert, Québec, 1981. ISBN 2-7607-0075-5. xiii + 317 pp.

Lafler:2002:PSS

[Laf02] Kirk Paul Lafler. Power SAS: a survival guide. Apress, Berkeley, CA, USA, 2002. ISBN 1-59059-066-X. xxxiv + 291 pp. LCCN QA276.4 .L34 2002.

LaLonde:1992:BRB

[LaL92] Steven M. LaLonde. Book review: SAS Applications Programming, a Gentle Introduction by Frank C. DiIorio. Technometrics, 34(3):366– 367, August 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270061.

LaLonde:2009:BRB

[LaL09] Steven M. LaLonde. Book review: Statistics Using SAS(R) Enterprise Guide by James B. Davis. Technometrics, 51(1): 103, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/40586574.

Langston:1981:SWM

[Lan81] Richard D. Langston. SAS/ GRAPH world map data sets. SAS technical report; P-117. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. 18 pp.

Langston:1982:CYO

[Lan82] Richard D. Langston. Creating your own SAS/GRAPH map data sets with a digitizer. SAS technical report; A-107. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. 22 + 3 pp. This report, dated Mar. 10, 1982, has the same report no. as another report.

LaPlante:1981:ICA

[LaP81] Josephine M. LaPlante. An introduction to computer analysis in the social sciences and business using SAS. Learning packages in the policy sciences; PS-21. Policy Studies Associates, Croton-on-Hudson, NY, USA, 1981. ISBN 0-936826-16-9. v + 53 pp. LCCN QA276.4 .L37 1981.

Lawal:2003:CDA

[Law03] Bayo Lawal. Categorical data analysis with SAS and SPSS applications. Lawrence Erlbaum Associates, Mahwah, NJ, USA, 2003. ISBN 0-8058-4605-0. vii + 561 pp. LCCN QA278 .L384 2003. Includes CD-ROM.

Lawson:2008:SMA

[Law08] John Lawson. SAS macros for analysis of unreplicated 2^{k} and 2^{k−p} designs with a possible outlier. Journal of Statistical Software, 25(13):1–17, May 2008. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v25/i13.

Lawson:2010:DAE

[Law10] John Lawson. Design and Analysis of Experiments with SAS. Chapman and Hall/CRC texts in statistical science series. Chapman and Hall/CRC, Boca Raton, FL, USA, 2010. ISBN 1-4200-6060-0 (hardback). xiii + 582 pp. LCCN QA279 .L36 2010.

Lazic:2024:BRS

[Laz24] Stanley E. Lazic. Book review: Statistical Methods for Mediation, Confounding and Moderation Analysis Using R and SAS, by Qingzhao Yu and Bin Li, CRC Press, 2022. 294pp, ISBN: 978-0-367-36547-9, £125.00. Journal of the Royal Statistical Society. Series A (Statistics in Society), 187 (1):265, January 2024. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://academic.oup.
com/jrsssa/article/187/1/
265/7235727.

LeCharlier:1994:SAF

[Le 94] Baudouin Le Charlier, editor. Static analysis: First International Static Analysis Symposium, SAS ’94, Namur, Belgium, September 28– 30, 1994: proceedings, volume 864 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1994. CODEN LNCSD9. ISBN 3-540-58485-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1994. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t0864.htm;   http://www.
springerlink.com/content/
978-3-540-58485-8;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=864. Spine title:

Static analysis symposium.

Lenz:1997:BRBa

[Len97] Hans J. Lenz. Book review:

A handbook of statistical analyses using SAS: B. S. Everitt and G. Der (1996) London: Chapman & Hall, ISBN 0-412-71050-1, pp. 158, £19.99. Computational Statistics & Data Analysis, 24(4):504, June 5, 1997. CODEN CSDADW.

ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167947397827636.

Lenth:2009:BRB

[Len09] Russell V. Lenth. Book review: Design and Analysis of Experiments: Classical and Regression Approaches with SAS. Journal of Statistical Software, 29(BR-5):??, January 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v29/b05.

Lenz:2010:BRG

[Len10] Sylvia Tamara Lenz. Book review: Gönen, Mithat (2007): Analyzing Receiver Operating Characteristic Curves with SAS. Statistical Papers, 51(3): 755–756, September 2010. CO-DEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.
                                        com/accesspage/article/10.
                                        1007/s00362-008-0156-x.

Lenz:2015:BRE

[Len15] Sylvia Tamara Lenz. Book review: Edward F. Vonesh: Generalized linear and nonlinear models for correlated data, theory and applications using SAS^{(R)}. Statistical Papers, 56(1):269–270, February 2015. CODEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.
                                        com/accesspage/article/10.
                                        1007/s00362-013-0563-5.

Levi:1998:SAI

[Lev98] G. (Giorgio) Levi, editor.

Static analysis: 5th International Symposium, SAS’98, Pisa, Italy, September 14– 16, 1998: Proceedings, volume 1503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1998. CODEN LNCSD9. ISBN 3-540-65014-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.I5733 1998. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1503.htm;   http:
//www.springerlink.com/content/
978-3-540-65014-0;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1503.

Lewis:2007:BRB

[Lew07] Steff Lewis. Book review:

Statistical Analysis of Medical Data Using SAS, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170 (4):1180–1181, October 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (elec-

tronic). URL http://www.
jstor.org/stable/4623234.

Lewis:1983:BSU

[LF83] Bruce R. Lewis and Richard K.

Ford. Basic statistics using SAS. West Pub. Co., St. Paul, 1983. ISBN 0-314-70619-4 (paperback). viii + 144 pp. LCCN QA276.4.L49 1983. US$7.95. Includes index.

Lewis:1987:BSU

[LF87] Bruce R. Lewis and Richard K. Ford. Basic statistics using SAS. West Pub. Co., St. Paul, MN, USA, second edition, 1987. ISBN 0-314-34736-4 (paperback). x + 148 pp. LCCN QA276.4 .L49 1987. Includes index.

Littell:1991:SSL

[LFS91] Ramon C. Littell, Rudolf J. Freund, and Philip C. Spector. SAS System for linear models. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1991. ISBN 1-55544-430-X. xviii + 329 pp. LCCN QA276.4 .L58 1991.

Liu:2011:CSF

[LG11] Ziyue Liu and Wensheng Guo. Code snippet: fmixed: a SAS macro for smoothing-spline-based functional mixed effects models. Journal of Statistical Software, 43(CS-1):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v43/c01.

Lenth:2007:SLP

[LH07] Russell V. Lenth and Søren Højsgaard. SASweave: Literate programming using SAS. Journal of Statistical Software, 19 (8):1–20, May 2007. CODEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v19/i08.

Littell:2006:SMM

[Lit06] Ramon C. Littell. SAS for mixed models. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 2006. ISBN 1-59047-500-3. xii + 814 pp. LCCN QA276.4 .S22254 2006.

Liu:2004:BRB

[Liu04]  Shin Ta Liu.  Book re-

view: SAS(R) Survival Analysis Techniques for Medical Research by Alan B. Cantor. Technometrics, 46(2):261, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470823.

Liu:2012:SBRb

[Liu12] Shuangzhe Liu. Short book review: Maximum Likelihood Estimation and Inference: With Examples in R, SAS, and ADMB by Russell B. Millar. International Statistical Review = Revue Internationale de Statistique, 80(2):346, August 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Liu:2017:BRK

[Liu17] Xulei Liu. Book review: Kleinman, Ken and Horton, Nicholas J., SAS and R: Data Management, Statistical Analysis, and Graphics. Boca Raton: CRC Press. Biometrics, 73 (4):1470, December 2017. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Latour:1995:IDA

[LJ95] Dom Latour and Marty Johnson. Interactive data analysis using SAS/INSIGHT software, releases 6.10-6.11: course notes. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. x + 238 pp.

Lissandre:1985:SSS

[LLS85] M. Lissandre, P. Lagier, and A. Skalli. SAS: a specification support system. In Teichroew and David [TD85], pages 221–238. ISBN 0-444-87731-2. LCCN QA76.9.S88 I327 1983.

Long:1986:PSC

[LM86] Susan B. Long and Hubert B. Mueller. Programming SAS for causal analysis. The American Statistician, 40(3):233– 234, August 1986. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2684552.

Loughin:2006:BRS

[Lou06] Thomas M. Loughin. Book review: SAS(R) for Mixed Models, 2nd edition, edited by Littell, R. C., Milliken, G. A., Stroup, W. W., Wolfinger, R. D., and Schabenberger, O. Biometrics, 62(4):1273– 1274, December 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Luginbuhl:1987:SGP

[LS87] Regina C. Luginbuhl and Sandra D. Schlotzhauer. SAS/ STAT guide for personal computers, version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1987. ISBN 1-55544-064-9. xviii + 1028 pp. LCCN QA 276.4 S34 1987. Edited by Regina C. Luginbuhl and Sandra D. Schlotzhauer. Includes bibliographical references and index.

Li:1998:UHO

[LS98] C. X. Li and J. Sun. Using hyperlink to organize SAS HTML output. In Anonymous [Ano98], pages 986–990. ISBN 1-58025-149-8. LCCN ????

Lawal:2002:GFV

[LS02] H. Bayo Lawal and Richard A. Sundheim. Generating factor variables for asymmetry, non-independence and skew-symmetry models in square contingency tables using SAS. Journal of Statistical Software, 7(8):1–23, 2002. CODEN JSSOBK. ISSN ???? URL http://

www.jstatsoft.org/v07/i08;
http://www.jstatsoft.org/
v07/i08/generate.pdf;   http:
//www.jstatsoft.org/v07/
i08/GENERATE.txt;   http:
//www.jstatsoft.org/v07/
i08/updates.

Lemmon:2005:DSS

[LS05] David R. Lemmon and J. L.

(Joseph L.) Schafer. Developing statistical software in Fortran 95. Statistics and computing. Springer-Verlag Inc., New York, NY, USA, 2005. ISBN 0-387-28123-1, 0-387-23817-4 (paperback). xv + 323 pp. LCCN QA76.5 .L453

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0662/2004061447-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0818/2004061447-
t.html.

Lawson:2009:CSB

[LSC09] John Lawson, G. Bruce Schaalje, and Bruce Jay Collings. Code snippet: Blocking mixed-level factorials with SAS. Journal of Statistical Software, 32(CS-1): ??, November 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v32/c01.

Littell:2002:SLM

[LSF02] Ramon C. Littell, Walter W. Stroup, and Rudolf J. Freund. SAS for linear models. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, fourth edition, 2002. ISBN 0-471-22174-0, 1-59047-023-0. xii + 466 pp. LCCN QA276.4 .L58 2002.

Lu:2008:BRB

[Lu08] Z. Q. John Lu. Book review: Optimum Experimental Designs, with SAS, by A. C. Atkinson; A. N. Donev; R. D. Tobias. Journal of the Royal Statistical Society. Series A (Statistics in Society), 171(2):504, April 2008. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

URL http://www.jstor.org/
stable/30130770.

Lu:2019:BRS

[Lu19] Minggen Lu. Book review:

Survival Analysis with Interval-Censored Data: A Practical Approach with Examples in R, SAS, and BUGS. The American Statistician, 73(2):211– 212, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//amstat.tandfonline.com/
doi/full/10.1080/00031305.
2019.1603477.

Lumley:2008:BRA

[Lum08] Thomas Lumley. Book review:

Analysis of correlated data with SAS and R (3rd edn). Mohamed M. Shoukri and Mohammad A. Chaudhary, Chapman and Hall/CRC, Boca Raton, 2007. No. of pages: 295. Price: $89.95. ISBN: 1-58488-619-6. Statistics in Medicine, 27 (22):4610–4611, September 30, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lovaglio:2016:CSE

[LV16] Pietro Giorgio Lovaglio and Gianmarco Vacca. Code snippet: %ERA: A SAS macro for extended redundancy analysis. Journal of Statistical Software, 74(??):??, ???? 2016. CO-DEN JSSOBK. ISSN 1548-

7660.  URL https://www.

                                        jstatsoft.org/index.php/
                                        jss/article/view/v074c01;
                                         https://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v074c01/v74c01.pdf.

Lovaglio:2017:GSM

[LV17] Pietro Giorgio Lovaglio and Gianmarco Vacca. %Gra: an SAS macro for generalized redundancy analysis. Journal of Statistical Computation and Simulation, 87(5):1048–1060, 2017. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Laaksonen:2011:SMC

[LVK^{+}11] Maarit A. Laaksonen, Esa Virtala, Paul Knekt, Hannu Oja, and Tommi Härkänen. SAS macros for calculation of population attributable fraction in a cohort study design. Journal of Statistical Software, 43(7):??, July 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

                                        www.jstatsoft.org/v43/i07.

Lautato:1995:SFP

[LW95] Theresa Lautato and Kim Willis. SAS fundamentals: a point-and-click approach: course notes. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. xvi + 574 pp.

Maindonald:2012:SBRb

[Mai12] John H. Maindonald. Short book review: R for SAS and SPSS Users, Second Edition, by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 80(1):200–201, April 2012. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

jstor.org/stable/23257195.

MartinezGarza:1983:IAS

[Mar83] Angel Martinez Garza. Introduccion al SAS: Sistema para Analisis Estadistico = Statistical Analysis System. Centro de Estadistica y Calculo, Chapingo, Mexico, 1983. 166 pp. LCCN QA276.4 .M27 1983.

Marasinghe:1988:BRB

[Mar88] Mervyn G. Marasinghe. Book review: SAS System for Linear Models by R. Freund; R. Littell; P. Spector. Technometrics, 30(1):116–117, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1270330.

Mason:1988:BRB

[Mas88] Robert L. Mason. Book review: SAS System for Regression by R. J. Freund; R. C. Littell. Technometrics, 30(1):117– 118, February 1988. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270331.

Multrus:1992:SSP

[MB92] Frank Multrus and Michael Bleicher. SAS-System: eine praxisbezogene Einführung. (German) [SAS system: a practical introduction]. Hüthig, Heidelberg, Germany, 1992. ISBN 3-7785-2157-8. xix + 241 pp. LCCN ????

McFarland:2001:BRA

[McF01]  Richard McFarland.  Book

review: Applied multivariate statistics with the SAS system, Second Edition: R. Khattree and D. N. Naik; SAS Publishing, Cary, NC, 1999, 338pp., $39.95 ISBN 1-58025-357-1. Journal of Statistical Planning and Inference, 96(2):437–438, July 1, 2001. CODEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0378375800003104.

Merrill:1983:MGC

[Mer83] H. W. Barry Merrill. Merrill’s guide to computer performance evaluation: analysis of SMF/RMF data with SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1983. 340 pp. LCCN QA76.9.E94 M47 1983. Includes index. On spine: Merrill’s guide to CPE.

Merrill:1984:MEG

[Mer84] H. W. Barry Merrill. Merrill’s Expanded guide to computer performance evaluation using the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1984. ISBN 0-917382-54-4. x + 866 pp. LCCN QA76.9.E94 M47 1984.

Muller:2002:RAI

[MF02] Keith E. Muller and Bethel A.

Fetterman. Regression and ANOVA: an integrated approach using SAS software. John Wiley and Sons, Inc., New York, NY, USA, 2002. ISBN 0-471-46943-2 (Wiley paperback), 1-58025-890-5 (SAS). xi + 566 pp. LCCN QA278.2 .M82 2002.

Muller:2003:RAI

[MF03] Keith E. Muller and Bethel A.

Fetterman. Regression and ANOVA: an integrated approach using SAS software. John Wiley and Sons, Inc., New York, NY, USA, 2003. ISBN 0-471-46943-2. xi + 566 pp. LCCN QA278.2 .M82 2003.

McMillan:2005:SMB

[MH05] Garnett McMillan and Timothy Hanson. SAS macro BDM for fitting the Dale regression model to bivariate ordinal response data. Journal of Statistical Software, 14 (2):1–12, May 1, 2005. CO-DEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/counter.php?
id=127&url=v14/i02&ct=2;
http://www.jstatsoft.org/
counter.php?id=127&url=v14/
i02/v14i02.pdf&ct=1.

Miller:1992:SNS

[Mil92] A. J. Miller. Some notes on statistical packages for PCs. Technical report, CSIRO Division of Mathematics & Statistics, Melbourne, Australia, October 1992. 55 pp.

Millar:1993:ATS

[Mil93]  Russell B. Millar.  Analy-

sis of trawl selectivity studies (addendum): implementation in SAS. Fisheries Research, 17(3–4):373–377, August 1993. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        016578369390137V.

Miron:1993:SSS

[Mir93] Thomas Miron. SAS software solutions, basic data processing. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-536-5. iv + 234 pp. LCCN ????

Miron:1995:HBS

[Mir95] Thomas Miron. The howto book for SAS/GRAPH software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-233-1. xii + 286 pp. LCCN ????

May:1997:SMM

[MJ97] Warren L. May and William D. Johnson. A SAS^{R} macro for the multivariate extension of the Kruskal–Wallis test including multiple comparisons: Randomization and κ^{2} criteria. Computational Statistics & Data Analysis, 26(2):239–250, December 4, 1997. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S016794739782107X.

Marasinghe:2008:SDA

[MK08] Mervyn G. Marasinghe and William J. (William Jo) Kennedy. SAS for data analysis: intermediate statistical methods. Statistics and computing. Springer-Verlag Inc., New York, NY, USA, 2008. ISBN 0-387-77371-1 (hardcover). xii + 557 pp. LCCN QA276.4 .M266 2008.

Monti:1997:MST

[Mon97] Katherine L. Monti. Making statistical tables in windows: a comparison of Data Muncher and SAS (PROC TABULATE and PROC REPORT). The American Statistician, 51(1): 83–??, February 1997. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
amstat.org/publications/
tas/abstracts/monti.html.

Moore:2003:BRB

[Moo03] Terri L. Moore. Book review:

SAS(R) for Linear Models by Ramon Littell; Walter Stroup; Rudolf Freund. Technometrics, 45(2):172–173, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/25047018.

Morris:2002:BRBb

[Mor02] Clare Morris. Book review: SAS Programming for Researchers and Social Scientists, by P. E. Spector. Journal of the Royal Statistical Society. Series D (The Statistician), 51(4):586, ???? 2002. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://www.

                                        jstor.org/stable/3650378.

McMahon:2006:GMM

[MPT06] James M. McMahon, Enrique R. Pouget, and Stephanie Tortu. A guide for multilevel modeling of dyadic data with binary outcomes using SAS PROC NLMIXED. Computational Statistics & Data Analysis, 50(12):3663–3680, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167947305001878.

Montgomery:2012:ILR

[MPV12] Douglas C. Montgomery, Elizabeth A. Peck, and G. Geoffrey Vining. Introduction to linear regression analysis. Wiley series in probability and statistics. John Wiley and Sons, Inc., New York, NY, USA, fifth edition, 2012. ISBN 0-470-54281-0. xvi + 645 pp. LCCN QA278.2 .M65 2012.

                                        URL http://proquestcombo.
                                        safaribooksonline.com/9780470542811.

Mine:2012:SAI

[MS12] Antoine Miné and David Schmidt, editors. Static Analysis: 19th International Symposium, SAS 2012, Deauville, France, September 11–13, 2012. Proceedings, volume 7460 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33124-6 (print), 3-642-33125-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-642-33125-1.

Meng:2014:BRH

[MT14] Xiangxiang Meng and Wayne Thompson. Book review: Handbook of SAS(R) DATA Step Programming. Journal of Statistical Software, 56(BR-1): ??, January 2014. CODEN JS-SOBK. ISSN 1548-7660. URL

http://www.jstatsoft.org/
v56/b01.

Muenchen:2009:RSS

[Mue09] Robert A. Muenchen. R for SAS and SPSS users. Statistics and computing. Springer-Verlag Inc., New York, NY, USA, 2009. ISBN 0-387-09417-2. xvii + 470 pp.

Mukhopadhyay:2015:BRB

[Muk15] Purna Mukhopadhyay. Book review: Multiple Imputation of Missing Data Using SAS. International Statistical Review = Revue Internationale de Statistique, 83(2):326–327, August 2015. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Mwitondi:2011:DAU

[Mwi11] Kassim S. Mwitondi. Data analysis using SAS ENTER-PRISE GUIDE. Journal of Applied Statistics, 38(12):2993–??, December 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mwitondi:2012:SDM

[Mwi12] Kassim S. Mwitondi. Statistical data mining using SAS applications. Journal of Applied Statistics, 39(10):2302–??, October 2012. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Mycroft:1995:SAS

[Myc95] Alan Mycroft, editor. Static analysis: Second International Symposium, SAS ’95, Glasgow, UK, September 25–27, 1995: proceedings, volume 983 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1995. CODEN LNCSD9. ISBN 3-540-60360-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .I5733 1995. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t0983.htm;   http://www.
springerlink.com/content/
978-3-540-60360-3;   http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=983.

Muthusi:2021:SGS

[MYM21] Jacques Muthusi, Peter W.

Young, and Samuel Mwalili. %svy freqs: a generic SAS macro for creating publicationquality three-way cross-tabulations. Journal of Open Research Software, 9(1):30–??, October 22, 2021. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
318/.

Nash:2008:BRB

[Nas08] Maliha S. Nash. Book review: Data Preparation for Analytics Using SAS by Gerhard Svolba. Technometrics, 50(1): 92, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25471437.

NCHS:1993:ESP

[Nat93] National Center for Health Statistics. Effective SAS programming techniques. ARC seminars in SAS programming. ARC Professional Services Group, Rockville, MD, 1993. ii + 92 pp.

Neville:2018:MCT

[NB18] Zachariah Neville and Naomi C.

Brownstein. Macros to conduct tests of multimodality in SAS. Journal of Statistical Computation and Simulation, 88(17):

3269–3290, 2018. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Neuhaus:2006:BRB

[Neu06] John M. Neuhaus. Book review: Fixed Effects Regression Methods for Longitudinal Data Using SAS, by Paul D. Allison. Journal of the American Statistical Association, 101(475): 1308, September 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/
stable/27590804.

Nielson:2007:SAI

[NF07] Hanne Riis Nielson and Gilberto Filé, editors. Static Analysis: 14th International Symposium, SAS 2007, Kongens Lyngby, Denmark, August 22–24, 2007. Proceedings, volume 4634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2007. CODEN LNCSD9. ISBN 3-540-74060-0 (print), 3-540-74061-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-74061-2.

Ng:2012:BRB

[Ng12] Hon Keung Tony Ng. Book review: Statistical Programming in SAS by A. John Bailer. Technometrics, 54(4): 441, November 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/41714930.

Nie:2008:BRO

[Nie08] Lei Nie. Book review: Optimum experimental designs, with SAS. A. C. Atkinson, A. N. Donev and R. D. Tobias Oxford University Press, Oxford, 2007. No. of pages: xvi + 511. Price: £35.00 (paperback), £75.00 (hardback). ISBN 13: 978-0-19-929660-6, 978-0-19-929659-0. Statistics in Medicine, 27(28):6034–6035, December 10, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nielsen:2011:BRS

[Nie11] Søren Feodor Nielsen. Book review: SAS for data analysis. Journal of Applied Statistics, 38(8):1743–1744, August 2011. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic). See [MK08].

Nooraee:2014:GLO

[NMvdH14] Nazanin Nooraee, Geert Molenberghs, and Edwin R. van den Heuvel. GEE for longitudinal ordinal data: Comparing R-geepack, R-multgee, Rrepolr, SAS-GENMOD, SPSS-GENLIN. Computational Statistics & Data Analysis, 77(??):70–83, September 2014. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0167947314000863.

Nevalainen:2006:SIM

[NO06] Jaakko Nevalainen and Hannu Oja. SAS/IML macros for a multivariate analysis of variance based on spatial signs. Journal of Statistical Software, 16(5):1–17, May 2006. CODEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v16/i05.

NorthCarolina:1989:RSI

[Nor89] North Carolina.State Information Processing Needs and Cost Study Commission. Report of the State Information Processing Needs and Cost Study Commission: report to the 1989 General Assembly of North Carolina, Raleigh, NC. North Carolina State Government, Raleigh, NC, USA, 1989. iii + 9, A-85 pp. LCCN MLCM 93/01566 (Z). North Carolina state doc. no.: Y4 I43.

Novikov:2003:RES

[Nov03] Ilya Novikov. A remark on efficient simulations in SAS. Journal of the Royal Statistical Society. Series D (The Statistician), 52(1):83–86, ???? 2003. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/
                                        stable/4128171.

Nassiri:2018:EDS

[NTGD18] Mohammadreza Nassiri, Mahdi Elahi Torshizi, Shahrokh Ghovvati, and Mohammad Doosti. Evaluation of different statistical methods using SAS software: an in silico approach for analysis of real-time PCR data. Journal of Applied Statistics, 45(2): 306–319, February 2018. CO-DEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Nelson:1988:SIP

[NZ88] Dean E. Nelson and Gary O.

Zerbe. A SAS/IML program to execute randomization analysis of response curves with multiple comparisons. The American Statistician, 42(3): 231–232, August 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2685025.

OBrien:2008:SBRa

[O’B08] C. M. O’Brien. Short book review: Adaptive Design Theory and Implementation Using SAS and R by Mark Chang. International Statistical Review = Revue Internationale de Statistique, 76(2):300, August 2008. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (elec-

tronic). URL http://www.
jstor.org/stable/27919619.

OBrien:2013:SBRh

[O’B13] Carl M. O’Brien. Short book review: Applied Operational Research with SAS by Ali Emrouznejad, William Ho. International Statistical Review = Revue Internationale de Statistique, 81(3):474, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Olsbjerg:2015:CSP

[OC15] Maja Olsbjerg and Karl Bang Christensen. Code snippet:

                                        %lrasch mml: A SAS macro

for marginal maximum likelihood estimation in longitudinal polytomous Rasch models. Journal of Statistical Software, 67(CS-2):??, ???? 2015. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/index.php/
                                        jss/article/view/v067c02;
                                         http://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v067c02/v67c02.pdf.

AICPA:1996:ISN

[oCPA96] American Institute of Certified Public Accountants. Implementing SAS no. 70: reports on the processing of transactions by service organizations. Auditing procedure study. The Institute, Jersey City, NJ, USA, 1996. ISBN 0-87051-174-2. xii + 171 pp. LCCN ????

Oglesby:1993:SPE

[Ogl93] Jerry L. Oglesby, editor. SUGI 18: proceedings of the Eighteenth Annual SAS Users Group International Conference, New York, New York, May 9–12, 1993. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993.

ISBN 1-55544-550-0. LCCN ????

Ordway:1989:AFL

[OGR89] Lynn Ordway, Fred Gurzeler, and William Rogers. Almost a free lunch: SAS to STATA download procedure. Technical Report P-7578, Rand Corp., Santa Monica, CA, USA, 1989. xii + 19 pp. Cover title.

ORourke:2013:SSA

[OH13] Norm O’Rourke and Larry Hatcher. A Step-by-Step Approach to Using SAS(R) for Factor Analysis and Structural Equation Modeling. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 2013. ISBN 1-59994-230-5. xxii + 420 pp. LCCN QA276.4 H37 2013.

Olson:1980:SGP

[OK80] R. J. (Richard J.) Olson and K. D. Kumar. A SAS graphics procedure, DISPLA1. Springfield, VA, USA, 1980. ix + 39 pp.

Olive:2010:BRBa

[Oli10] David J. Olive. Book review:

Analysis of Variance Designs: A Conceptual and Computational Approach With SPSS and SAS by G. Gamst; L. S. Meyers; A. J. Guarino. Technometrics, 52(1):139, February 2010. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586688.

Olivier:2017:BRA

[Oli17] Jake Olivier. Book review: Applied surrogate endpoint evaluation methods with SAS and R. Ariel Alonso, Theophile Bigirumurame, Tomasz Burzykowski, Marc Buyse, Geert Molenberghs, Leacky Muchene, Nolen Joy Perualila, Ziv Shkedy, Wim Van der Elst, Chapman and Hall/CRC Biostatics Series, 2017, xxi + 373, 57.99 GBP, ISBN 13: 978-1-4822-4936-1 (Hardback). Statistics in Medicine, 36(24):3919–3920, October 30, 2017. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olive:2020:BRS

[Oli20] David J. Olive. Book review: Statistical Programming in SAS (2nd ed.), by A. J. Bailer. Boca Raton, FL: Chapman & Hall/ CRC, Taylor & Francis Group, 2020, xv + 362 pp., $79.95, ISBN: 978-0-367-35797-9. Technometrics, 62(2):1–288, 2020. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Ona:1990:PFA

[Ona90] Tony L. Ona, editor. Proceedings of the Fifteenth Annual SAS Users Group International Conference, Nashville, Tennessee, April 1–4, 1990: SUGI 15. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-403-2. LCCN QA276.4 .S2397 1990. Includes bibliographical references and indexes.

Ooms:2010:BRB

[Oom10] Jeroen Ooms. Book review:

SAS and R. Journal of Statistical Software, 37(BR-3): ??, December 2010. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v37/b03.

J:2013:BPT

[Opd13] John D. (“J. D.”) Opdyke.

Bootstraps, permutation tests, and sampling orders of magnitude faster using SAS(R). WIREs Computational Statistics, 5(5):391–405, September/ October 2013. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Palsberg:2000:SAI

[Pal00] Jens Palsberg, editor. Static analysis: 7th International Symposium, SAS 2000, Santa Barbara, CA, USA, June 29–July 1, 2000: Proceedings, volume 1824 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2000. CO-DEN LNCSD9. ISBN 3-540-67668-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1824. URL http://

link.springer-ny.com/link/
service/series/0558/tocs/
t1824.htm;   http://www.
springerlink.com/content/

                                        978-3-540-67668-3;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-
                                        9743&volume=1824.

Pan:2009:CSS

[Pan09] Wei Pan. Code snippet: a SAS/IML macro for computing percentage points of Pearson distributions. Journal of Statistical Software, 31(CS-2): ??, September 2009. CODEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v31/c02.

Pardoe:2006:ARM

[Par06] Iain Pardoe. Applied regression modeling: a business approach. Wiley-Interscience, New York, NY, USA, 2006. ISBN 0-471-97033-6. xx + 303 pp. LCCN QA278.2 .P363

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0741/2006044262-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0741/2006044262-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/fy0702/2006044262.
                                        html;   http://www.wiley.
                                        com/WileyCDA/WileyTitle/
                                        productCd-0471970336.html.

Patefield:2002:FNL

[Pat02]  Mike Patefield.  Fitting

non-linear structural relationships using SAS procedure NLMIXED. Journal of the Royal Statistical Society. Series D (The Statistician), 51 (3):355–366, ???? 2002. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/
stable/3650279.

Pelz:1995:ASD

[Pel95] Corey J. Pelz. Analysis of survival data: a comparison of three major statistical packages (SAS, SPSS, BMDP). Thesis, (m.s.), Medical College of Wisconsin, ????, 1995. 128 pp.

Perrett:2010:SIC

[Per10] Jamis J. Perrett. A SAS/IML Companion for Linear Models. Statistics and computing. Springer-Verlag Inc., New York, NY, USA, 2010. ISBN 1-4419-5556-9 (paperback), 1-4419-5557-7 (e-book). ISSN 1431-8784. xiv + 228 pp. LCCN QA276.4 .P468 2010.

Preisser:2005:ACF

[PG05] John S. Preisser and Daniel I.

Garcia. Alternative computational formulae for generalized linear model diagnostics: identifying influential observations with SAS software. Computational Statistics & Data Analysis, 48(4):755–764, April 1, 2005. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947304001008.

Pierson:1987:SMT

[Pie87a] Gil Pierson. SAS merged with TEX. TUGboat, 8(2):178, July 1987. ISSN 0896-3207.

Pierson:TB8-2-178

[Pie87b]  Gil Pierson.  SAS merged

with TEX. TUGboat, 8(2): 178–179, July 1987. ISSN 0896-3207. URL http://

                                        www.tug.org/TUGboat/tb08-
                                        2/tb18site.pdf.

Pinon:2010:BRB

[Pin10] Antonio Pinon. Book review: A Handbook of Statistical Analyses using SAS, by G. Der; B. S. Everitt. Journal of the Royal Statistical Society. Series A (Statistics in Society), 173 (2):464–465, April 2010. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/
                                        stable/40666262.

Pham:2020:CSA

[PKC^{+}20] Thanh V. Pham, Jeffrey D. Kromrey, Yi-Hsin Chen, Eun Sook Kim, Diep T. Nguyen, and Yan Wang. Code snip-

                                        pet: ANOVA robust: a SAS

macro for parametric tests of mean differences in onefactor ANOVA models. Journal of Statistical Software, 95 (??):??, ???? 2020. CO-DEN JSSOBK. ISSN 1548-

                                        7660.  URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v095c02;
                                         https://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v095c02/v95c02.pdf.

Pope:1997:USS

[Pop97] P. Pope. Using the SAS system for large volume HTML document production. In Anonymous [Ano97], pages 988–993. ISBN 1-55544-961-1. LCCN ????

Pratter:2008:SGJ

[Pra08] Frederick M. Pratter. SAS graphics for Java: Examples using SAS AppDev Studio and the Output Delivery System. The American Statistician, 62 (4):359, November 2008. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Prins:2005:SAD

[Pri05] Samantha C. Bates Prins. Statistical analysis and data display: An intermediate course with examples in S-Plus, R, and SAS. The American Statistician, 59(3):275–276, August 2005. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Palsberg:2009:SAI

[PS09] Jens Palsberg and Zhendong Su, editors. Static Analysis: 16th International Symposium, SAS 2009, Los Angeles, CA, USA, August 9–11, 2009. Proceedings, volume 5673 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2009. CO-DEN LNCSD9. ISBN 3-642-03236-2 (print), 3-642-03237-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-642-03237-0.

Pukelsheim:2007:SBR

[Puk07] Friedrich Pukelsheim. Short book review: Optimum Experimental Designs, with SAS by Anthony Atkinson, Alexander Donev, Randall Tobias. International Statistical Review = Revue Internationale de Statistique, 75(3):413, December 2007. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic). URL http://www.

                                        jstor.org/stable/41509884.

Payton:1993:CUS

[PYC93] Mark E. Payton, Linda J. Young, and J. P. Chandler. C404. Using SAS to solve a system of non-linear equations. Journal of Statistical Computation and Simulation, 47(1–2): 114–117, 1993. CODEN JSC-SAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Qumsiyeh:2017:BAT

[QDG17] Maher Qumsiyeh, Robert Deis, and Dalton Gannon. Bootstrapping an autoregressive time series model using SAS. Advances and Applications in Statistics, 50(6):435–467, June 2017. CODEN ???? ISSN 0972-3617. URL http://www.

                                        pphmj.com/abstract/10940.
                                        htm.

Radtke:1994:HGS

[Rad94] Michelle Radtke. A help guide for the SAS novice. Thesis (b.s.), California Polytechnic State University, Pomona, CA, USA, 1994. various pp.

Rahnavard:2024:BRS

[Rah24] Ali Rahnavard. Book review:

Statistical Analytics for Health Data Science with SAS and R Jeffrey R. Wilson, Ding-Geng Chen, and Karl E. Peace, Boca Raton, FL: Chapman & Hall/ CRC Press, 2023, xxi + 257 pp., $99.95(H), ISBN: 978-1-032-32562-0. Journal of the American Statistical Association, 119(545):786–787, 2024. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

Raithel:1995:TSA

[Rai95] Michael A. Raithel. Tuning SAS applications in the MVS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-278-1. viii + 303 pp. LCCN ????

Ray:1982:SUG

[Ray82] Alice Allen Ray. SAS user’s guide: basics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. xiii + 923 pp. Includes index.

Recchia:2010:CSC

[Rec10a] Anthony Recchia. Code snippet: Contiguity-constrained hierarchical agglomerative clustering using SAS. Journal of Statistical Software, 33(CS-2): ??, February 2010. CODEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v33/c02.

Recchia:2010:CSR

[Rec10b] Anthony Recchia. Code snippet: R-squared measures for two-level hierarchical linear models using SAS. Journal of Statistical Software, 32(CS-2): ??, January 2010. CODEN JS-SOBK. ISSN 1548-7660. URL

                                        http://www.jstatsoft.org/
                                        v32/c02.

Reese:2009:BRBa

[Ree09] R. Allan Reese. Book review: R for SAS and SPSS Users, by R. A. Muenchen. Journal of the Royal Statistical Society. Series A (Statistics in Society), 172 (3):697–698, June 2009. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/
                                        stable/20622537.

Reinhardt:1980:SSL

[Rei80] Patti S. Reinhardt. SAS supplemental library user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1980. ISBN 0-917382-11-3 (??invalid checksum??) (paperback). 202 pp. LCCN QA276.4 .S25 1980. Includes bibliographies.

Rahmansyah:2024:BRD

[RHWP24] Egi Rahmansyah, Nur Hidayah, Megawati Zein Waliulu, and Hawinda Restu Putri. Book review: Data Science and Machine Learning for Non-Programmers Using SAS Enterprise Miner, 1st ed. Dothang Truong, CRC Press Taylor & Francis Group, 2024. IX + 565 pages, £59,99 (hbk) ISBN: 978-0-367-75538-6 (hbk). Boca Raton and Abingdon. Scope: Textbook. Level: Students, lecturers, researchers, and industry professionals from various backgrounds. Technometrics, 66(3):493–494, 2024. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

Richter:2008:BRW

[Ric08] Bernd Richter. Book review:

Walter Krämer, Olaf Schoffer, Lars Tschiersch: Datenanalyse mit SAS — Statistische Verfahren und ihre grafischen Aspekte. Statistical Papers, 49(4): 799–800, October 2008. CO-DEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

URL http://link.springer.
com/accesspage/article/10.
1007/s00362-007-0091-2.

Richardson:2013:SBRb

[Ric13] Alice Richardson. Short book review: Applied Medical Statistics Using SAS by Geoff Der, Brian S. Everitt. International Statistical Review = Revue Internationale de Statistique, 81 (3):476–477, December 2013. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Righolt:2018:SCP

[RMM18] Christiaan H. Righolt, Barret A. Monchka, and Salaheddin M. Mahmud. From source code to publication: Code Diary, an automatic documentation parser for SAS. SoftwareX, 7(??):217–221, January/June 2018. CODEN ???? ISSN 2352-7110. URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S2352711018300669.

Robinson:2007:BRBa

[Rob07] Derek Robinson. Book review: SAS for Mixed Models, by R. C. Littell; G. A. Milliken; W. W. Stroup; R. D. Wolfinger; O. Schabenberger. Journal of the Royal Statistical Society. Series A (Statistics in Society), 170(1):257–258, January 2007. CODEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic). URL http://www.

                                        jstor.org/stable/4623152.

Robinson:2008:BRS

[Rob08] Derek Robinson. Book review: Statistical analysis of medical data using SAS. Geoff Der and Brian S. Everitt, Chapman and Hall/CRC, Boca Raton, FL, 2006. No. of pages: 428. Price: $79.95. ISBN 978-1-58488-469-X. Statistics in Medicine, 27 (8):1326–1327, April 15, 2008. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:2009:BRB

[Rob09] Andrew Robinson. Book review: Adaptive design theory and implementation using SAS and R. Journal of Applied Statistics, 36(6):701–702, June 2009. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Rodriguez:2011:S

[Rod11] Robert N. Rodriguez. SAS.

WIREs Computational Statistics, 3(1):1–11, January/February 2011. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

Rose:1983:DEU

[Ros83] R. F. Rose. A “data engine”

using SAS and INQUIRE. J. Med. Syst., 7(3):257–266, June 1983.

Ross:2010:IPM

[Ros10]  Sheldon M. Ross.  Intro-

duction to probability models. Academic Press, New York, NY, USA, tenth edition, 2010. ISBN 0-12-375686-3 (hardcover). xv + 784 pp. LCCN QA273 .R84X 2010; QA273 .R84 2010. URL http://www.

sciencedirect.com/science/
book/9780123756862.

Ross:2014:IPM

[Ros14] Sheldon M. Ross. Introduction to probability models. Academic Press, New York, NY, USA, eleventh edition, 2014. ISBN 0-12-407948-2. xv + 767 pp. LCCN QA273 .R84 2014.

Rosenblad:2015:BRB

[Ros15] Andreas Rosenblad. Book review: A Step-by-Step Approach to Using SAS^{(}R) for Factor Analysis and Structural Equation Modelling, Second Edition. International Statistical Review = Revue Internationale de Statistique, 83(2): 325–326, August 2015. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Rutherford:2012:BRB

[Rut12] Jim Rutherford. Book review: SAS(R) Programming for Enterprise Guide(R) Users (2nd ed.) by Neil Constable. Technometrics, 54(1):104, February 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/41714846.

Suissa:1987:GLR

[SA87] Samy Suissa and Jeanne Adam. Generalized linear regression for discrete data using SAS. The American Statistician, 41 (3):241, August 1987. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2685122.

Saffer:1983:SCV

[Saf83] Marian Saffer. SAS companion for the VM/CMS operating system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1983. ISBN 0-917382-50-1. vii + 161 pp. LCCN QA276.4 .S17 1983. Includes index.

Sall:1978:SMA

[Sal78a] John Sall. SAS merge applications. SAS technical report; A-101. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1978. 41 pp.

Sall:1978:SUA

[Sal78b] John Sall. SAS update applications. SAS technical report; A-103. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1978. 16 pp.

Sall:1979:SET

[Sal79a] John Sall. SAS econometric and time-series library: preliminary test version for SAS. SAS technical report; P-110. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1979. 95 pp.

Sall:1979:USL

[Sal79b] John Sall. Using SAS for large data sets. SAS technical report; A-104. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, rev edition, 1979. 16 pp.

Sall:1980:CP

[Sal80] John Sall. The CITIBASE procedure. SAS technical report; P-114. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1980. 6 pp. “December 18, 1980 … SAS/ ETS.”.

Sall:1981:SRA

[Sal81a] John Sall. SAS regression applications. SAS technical report; A-102. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, rev edition, 1981. various pp.

Sall:1981:SCC

[Sal81b] Virginia B. Sall. SAS CMS companion: 1981 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. ISBN 0-917382-19-6. 130 pp. LCCN QA276.4 .S629 1981. Includes index.

Sall:1982:SCC

[Sal82] Virginia B. Sall. SAS CMS companion. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. ISBN 0-917382-19-3 (??invalid checksum??). 130 pp. LCCN QA276.4 .S225 1982. Includes bibliographical references and index.

Sall:1988:SII

[Sal88] John Sall. SAS/IML: An interactive matrix language. The American Statistician, 42(3): 232–233, August 1988. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2685027.

Samuelson:2012:LSD

[Sam12] Pamela Samuelson. Legally speaking: Do software copyrights protect what programs do? Communications of the ACM, 55(3):27–29, March 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Sandry:2003:BRBa

[San03] Thomas D. Sandry. Book review: Longitudinal Data and SAS(R): A Programmer’s Guide by Ron Cody. Technometrics, 45(2):173–174, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047020.

SAS:1976:SI

[SAS76] SAS Institute. The SAS/Tella-graf interface. AUI Data Graphics, Washington, DC, USA, 1976. 20 pp. Contents: Introductory guide to programming – The TAG procedure: users’ guide.

SAS:1980:SSN

[SAS80]  SAS Institute.  SAS 79.3

and SAS 79.4: new features, changes and enhancements. SAS technical report; P-111. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1980. 11 pp.

SAS:1981:SCE

[SAS81a]  SAS Institute.  SAS 79.5

changes and enhancements. SAS technical report; P-115. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. 202 pp.

SAS:1981:SSP

[SAS81b] SAS Institute. SAS 79.5 supplemental procedures. SAS technical report; S-120x. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. various pp. Consists of SAS technical reports S-120 through S-129 under one cover.

SAS:1981:SPG

[SAS81c] SAS Institute. SAS programmer’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. ISBN 0-917382-17-X (paperback). 208 pp. LCCN QA276.4 .S233 1981. Includes index.

SAS:1981:SUG

[SAS81d]  SAS Institute.  SAS/graph

user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1981. 126 pp. Includes index. Bibliography: p. 122.

SAS:1982:SES

[SAS82a] SAS Institute. SAS 82.2 enhancements to the SAS supplemental library. SAS technical report; S-130. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. 22 pp.

SAS:1982:SUG

[SAS82b] SAS Institute. SAS user’s guide: statistics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. ISBN 0-917382-32-3 (paperback), 0-917382-37-4 (paperback: cover). 584 pp. LCCN QA276.4 .S29 1982. Includes bibliographies and index.

SAS:1982:SFU

[SAS82c] SAS Institute. SAS/FSP user’s guide: full-screen product. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. ISBN 0-917382-32-3. 81 pp. LCCN CM.1.2.08.

SAS:1982:SGEa

[SAS82d] SAS Institute. SAS/GRAPH enhancements and updates for 79.6. SAS technical report; P-119. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. 23 pp.

SAS:1982:SGEb

[SAS82e] SAS Institute. SAS/GRAPH enhancements and updates for 82.2 release. SAS technical report; P-121. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. 76 pp.

SAS:1983:ESM

[SAS83a] SAS Institute. Enhancements to the SAS macro facility. SAS technical report; P-127. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1983. 30 pp.

SAS:1983:SPC

[SAS83b] SAS Institute. A SAS program for creating a weighted file. Health and Welfare Agency, Dept. of Health Services, 1983. 2 pp. Caption title. Includes bibliographical references.

SAS:1983:SUI

[SAS83c] SAS Institute. SAS/GRAPH under IBM CMS. Computing Services, University of California, 1983. 6 pp. Includes bibliography.

SAS:1984:SCO

[SAS84a] SAS Institute. SAS companion for OS operating systems and TSO. 1984 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1984. ISBN 0-917382-55-2. vi + 219 pp. LCCN QA276.4 .S22 1984.

SAS:1984:SUC

[SAS84b] SAS Institute. SAS under CMS. Computing Services, University of California, 1984. 10 pp. Includes bibliography.

SAS:1984:SUG

[SAS84c] SAS Institute. SAS/ETS user’s guide, version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1984. ISBN 0-917382-62-5 (paperback). xviii + 738 pp. LCCN HA30.3 .S28 1984. On t.p. the registered trademark symbol “TM” is superscript following “SAS/ETS” in the title. Includes bibliographies and index.

SAS:1985:SAG

[SAS85a] SAS Institute. SAS applications guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-16-1 (duplicate ISBN). 201 pp. LCCN ????

SAS:1985:SIG

[SAS85b] SAS Institute. SAS introductory guide for personal computers: version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-81-1. xiv + 111 pp. LCCN QA276.4 S18 1985. Includes index.

SAS:1985:SPGa

[SAS85c] SAS Institute. SAS procedures guide for personal computers: version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-79-X. xx + 373 pp. LCCN QA276.4 S193 1985. Includes bibliographies and index.

SAS:1985:SPGb

[SAS85d] SAS Institute. SAS programmer’s guide. SAS technical report; P-139. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition, 1985. ISBN 0-917382-17-3 (??invalid checksum??). 413 pp. LCCN ????

SAS:1985:SUGa

[SAS85e] SAS Institute. SAS user’s guide: basics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition, 1985. ISBN 0-917382-65-X (paperback). xx + 1290 pp. LCCN QA276.4 .S19 1985. Includes index. One folded reference card inserted.

SAS:1985:SUGb

[SAS85f] SAS Institute. SAS user’s guide: statistics, Version 5 Edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition, 1985. ISBN 0-917382-66-8. xvi + 956 pp. LCCN QA276.4 .S237 1985. Includes bibliographies and index.

SAS:1985:SAU

[SAS85g] SAS Institute. SAS/AF user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition, 1985. ISBN 0-917382-63-3 (paperback). xviii + 223 pp. LCCN QA276.4 .S238 1985. On t.p. the registered trademark symbol “TM” is superscript following “SAS/AF” in the title. Includes index.

SAS:1985:SEU

[SAS85h] SAS Institute. SAS/ETS user’s guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 5 edition, 1985. ISBN 0-917382-62-5. xviii + 738 pp. LCCN HA30.3 .S28 1984.

SAS:1985:SIU

[SAS85i] SAS Institute. SAS/IML user’s guide: version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. 300 pp.

SAS:1985:SGP

[SAS85j] SAS Institute. SAS/STAT guide for personal computers, version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. ISBN 0-917382-84-6. xvi + 378 pp. LCCN QA276.4 S197 1985. On t.p. the registered trademark symbol “TM” is superscript following “SAS/ STAT” in the title. Includes bibliographies and index.

SAS:1986:SCC

[SAS86] SAS Institute. SAS companion for the CMS operating system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 0-917382-50-1. xv + 101 pp. LCCN CM.1.3.09.

SAS:1987:MIS

[SAS87] SAS Institute. Master index to SAS system documentation, version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1987. ISBN 1-55544-047-9. vi + 224 pp. LCCN QA276.4 .M275 1987.

SAS:1988:FPF

[SAS88a]  SAS Institute.  FEMCAD-

88: proceedings of the Fourth SAS-World Conference, Paris, 17–19 October 1988. Technology transfer series. IITT-International, Gournay-sur-Marne, France, 1988. ISBN 2-907669-00-1 (set), 2-907669-01-X (v. 1), 2-907669-02-8 (v. 2). LCCN TA647 .S77 1988. Includes bibliographical references. Numerical and experimental analysis in structural optimization.

SAS:1988:SUGc

[SAS88b] SAS Institute. SAS-AF User’s Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-090-8. 305 pp. LCCN ???? US$25.95.

SAS:1988:SCCa

[SAS88c] SAS Institute. SAS-C Compiler. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-301-X. 680 pp. LCCN ???? US$29.95.

SAS:1988:SCCb

[SAS88d] SAS Institute. SAS-C Compiler and Library. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-078-9. 199 pp. LCCN ???? US$11.95.

SAS:1988:SCLa

[SAS88e] SAS Institute. SAS-C Library Reference, vol. 1. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-079-7. 257 pp. LCCN ???? US$13.95.

SAS:1988:SCLb

[SAS88f] SAS Institute. SAS-C Library Reference, vol. 2. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-081-9. 439 pp. LCCN ???? US$22.95.

SAS:1988:SCS

[SAS88g] SAS Institute. SAS-C Source Level Debugger User’s Guide. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-080-0. 206 pp. LCCN ???? US$12.95.

SAS:1988:SUGd

[SAS88h] SAS Institute. SAS-FSP User’s Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-092-4. 331 pp. LCCN ???? US$25.95.

SAS:1988:SGH

[SAS88i] SAS Institute. SAS-GRAPH hardware interfaces for personal computers: version 6 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-093-2. 353 pp. LCCN ???? US$12.95.

SAS:1988:SGU

[SAS88j] SAS Institute. SAS-GRAPH User’s Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-087-8. 549 pp. LCCN ???? US$34.95.

SAS:1988:SGP

[SAS88k] SAS Institute. SAS Guide to Problem Solving and Error Messages: version 5 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-302-8. 208 pp. LCCN ???? US$14.95.

SAS:1988:SUGa

[SAS88l] SAS Institute. SAS-IML User’s Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-094-0. 357 pp. LCCN ???? US$16.95.

SAS:1988:SIG

[SAS88m] SAS Institute. SAS introductory guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-096-7. xv + 111 pp. LCCN QA276.4 .S25 1988. US$11.95. Includes index.

SAS:1988:SLGa

[SAS88n] SAS Institute. SAS Language Guide for Personal Computers: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-099-1. 550 pp. LCCN ???? US$19.95.

SAS:1988:SLGb

[SAS88o] SAS Institute. SAS Language Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-091-6. xxii + 530 pp. LCCN QA276.4 .S26 1988. US$19.95.

SAS:1988:SPG

[SAS88p] SAS Institute. SAS Procedures Guide: release 6.03 edition.

SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-089-4. xxii + 441 pp. LCCN QA276.4 .S27 1988. US$15.95. Includes bibliographies and index.

SAS:1988:SUGb

[SAS88q]  SAS Institute.  SAS-STAT

User’s Guide: release 6.03 edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-094-0 (??duplicate ISBN??). 1028 pp. LCCN ???? US$33.95.

SAS:1988:STRa

[SAS88r] SAS Institute. SAS technical report: P-163 VSE-POWER accounting for release 5.16 of the SAS system. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-083-5. 34 pp. US$4.50.

SAS:1988:STRb

[SAS88s] SAS Institute. SAS technical report: P-173 transporting and converting version. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-084-3. 63 pp. US$3.50.

SAS:1988:STRc

[SAS88t] SAS Institute. SAS technical report: P-174 changes and enhancements to the SAS system release 5.16, for the VSE operating system. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-082-7. 392 pp. US$7.00.

SAS:1988:STRd

[SAS88u] SAS Institute. SAS technical report: P-175 changes and enhancements to the SAS system release 5.18, under OS and CMS. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-086-X. 307 pp. US$9.50.

SAS:1988:STRe

[SAS88v] SAS Institute. SAS technical report: P-177 system administrator’s guide to the SAS system, release 6.03, under UNIX. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-097-5. 96 pp. US$6.00.

SAS:1988:STRf

[SAS88w] SAS Institute. SAS technical report: P-178 additional SAS-GRAPH hardware interfaces, release 5.18. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-300-1. 55 pp. US$4.00.

SAS:1988:SEU

[SAS88x] SAS Institute. SAS/ETS user’s guide; Version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-325-7. xxx + 560 pp. LCCN CM.1.0.11. Includes bibliographical references.

SAS:1988:SOP

[SAS88y] SAS Institute. SUGI 13, Orlando: proceedings of the thirteenth annual SAS Users Group International Conference, Orlando, Florida, March 27–30, 1988. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. xxix + 1600 pp. Includes bibliographies and indexes.

SAS:1988:USS

[SAS88z] SAS Institute. Using the SAS System: release 6.03, under UNIX operating systems and derivatives. SAS technical report; P-176. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-098-3. viii + 91 pp. LCCN QA276.4 .U75 1988. Includes index.

SAS:1989:SGS

[SAS89a] SAS Institute. SAS guide to the SQL procedure: usage and reference. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-367-2. xviii + 210 pp. LCCN CM.1.0.06. Includes bibliographical references.

SAS:1989:SLP

[SAS89b] SAS Institute. SAS language and procedures: usage, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-371-0. xxx + 638 pp. LCCN QA276.4 .S232 1989. Includes bibliographical references and index.

SAS:1989:SAS

[SAS89c] SAS Institute. SAS/AF software, usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-369-9. xxx + 245 pp. LCCN QA276.4.S228 1989. Spine title: SAS/AF software. Includes bibliographical references and index.

SAS:1989:SCC

[SAS89d] SAS Institute. SAS/C compiler and library user’s guide: release 4.50. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1989. ISBN 1-55544-361-3. xiv + 390 pp. LCCN QA276.4 .S2385 1989.

SAS:1989:SCL

[SAS89e] SAS Institute. SAS/C library reference: release 4.50. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-359-1 (v. 1), 1-55544-360-5 (v. 2). various pp. LCCN QA276.4 .S239 1989.

SAS:1989:SFS

[SAS89f] SAS Institute. SAS/FSP software, usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-365-6. xl + 451 pp.

LCCN QA276.4.S23384 1989. Spine title: SAS/FSP software. Includes bibliographical references and index.

SAS:1989:SUG

[SAS89g] SAS Institute. SAS/OR user’s guide: version 6, first edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-324-9. xvii + 479 pp. LCCN T57.6.S1276 1989. The patent registration symbol is superscript following SAS/OR in title. Includes 1 fold. p. at back: ‘SAS/OR software reference card’. Spine title: SAS/ OR user’s guide. Includes bibliographical references and index.

SAS:1989:SSE

[SAS89h] SAS Institute. SAS/QC software examples: version 6. SAS technical report; P-188. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-349-4. 374 pp. LCCN QA276.4.S2 1989. Includes bibliographical references.

SAS:1989:TSF

[SAS89i] SAS Institute. Transporting SAS files between host systems. SAS technical report; P-195. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. ISBN 1-55544-374-5. xiv + 167 pp. LCCN CM.1.0.13.

SAS:1990:SCA

[SAS90a] SAS Institute. SAS companion for the AOS/VS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-407-5 (??duplicate ISBN??). xxiv + 271 pp. LCCN QA276.4 .S2214 1990.

SAS:1990:SCC

[SAS90b] SAS Institute. SAS companion for the CMS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-399-0. xxiv + 319 pp. LCCN CM.1.0.03.

SAS:1990:SCP

[SAS90c] SAS Institute. SAS companion for the PRIMOS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-407-5, 1-55544-406-7. xxiv + 270 pp. LCCN QA276.4 .S2215 1990.

SAS:1990:SCU

[SAS90d] SAS Institute. SAS companion for the UNIX environment and derivatives: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-420-2. xxii + 273 pp. LCCN QA276.4 .S222 1990. Includes index.

SAS:1990:SGM

[SAS90e] SAS Institute. SAS guide to macro processing: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1990. ISBN 1-55544-039-8. xxiv + 319 pp. LCCN QA276.4 .S2235 1990.

SAS:1990:SGR

[SAS90f] SAS Institute. SAS guide to the REPORT procedure: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-422-9. xxiv + 307 pp. LCCN QA276.4 .S224 1990. Spine title: SAS guide to the REPORT procedure. Includes bibliographical references and index.

SAS:1990:SLPa

[SAS90g] SAS Institute. SAS language and procedures: introduction. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-410-5. xvi + 124 pp. LCCN QA276.4 .S2316 1990. Includes index.

SAS:1990:SLPb

[SAS90h] SAS Institute. SAS language and procedures: syntax, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-400-8. xiv + 175 pp. LCCN QA276.4 .S2317 1990.

SAS:1990:SLR

[SAS90i] SAS Institute. SAS language:

reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-381-8. xxx + 1042 pp. LCCN QA276.4.S2344 1990. Includes bibliographies and index.

SAS:1990:SPG

[SAS90j] SAS Institute. SAS procedures guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1990. ISBN 1-55544-378-8. xi + 705 pp. LCCN QA276.4.S23486 1990. Includes index.

SAS:1990:SPT

[SAS90k] SAS Institute. SAS programming tips: a guide to efficient SAS processing. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-431-8. xvi + 155 pp. LCCN QA276.4 .S2333 1990.

SAS:1990:SSC

[SAS90l] SAS Institute. SAS screen control language: reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-386-9. xxii + 442 pp. LCCN QA276.4.S235 1990. Includes index.

SAS:1990:SAIa

[SAS90m] SAS Institute. SAS/ACCESS interface to ADABAS: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-411-3. xxii + 244 pp. LCCN QA276.4 .S2377 1990.

SAS:1990:SAIb

[SAS90n] SAS Institute. SAS/ACCESS interface to CA-DATACOM/ DB: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-412-1. xxii + 211 pp. LCCN QA276.4 .S215 1990.

SAS:1990:SAIc

[SAS90o] SAS Institute. SAS/ACCESS interface to OS/2 database products. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-432-6. xvi + 141 pp. LCCN QA276.4 .S216 1990.

SAS:1990:SAId

[SAS90p] SAS Institute. SAS/ACCESS interface to SQL/DS: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-385-0. xx + 191 pp. LCCN QA276.4 .S217 1990.

SAS:1990:SAS

[SAS90q] SAS Institute. SAS/ASSIST software: your interface to the SAS System: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-375-3. viii + 50 pp.

SAS:1990:SCS

[SAS90r] SAS Institute. SAS/CONNECT software, usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-408-3. xxvi + 287 pp. LCCN QA76.8.S27 S274 1990. Includes bibliographical references and index.

SAS:1990:SSI

[SAS90s] SAS Institute. SAS/GRAPH software. Introduction, Version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-395-8. xviii + 122 pp. LCCN QA276.4 .S2364 1990. Includes index.

SAS:1990:SGS

[SAS90t] SAS Institute. SAS/GRAPH software: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-379-6. LCCN T385.S26 1990, QA276.4.S2746 1990. Includes bibliographical references and indexes.

SAS:1990:SIS

[SAS90u] SAS Institute. SAS/IML software: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-377-X. xxviii + 501 pp. LCCN QA276.4.S2377 1990. Includes bibliographical references (p. 489-490) and index.

SAS:1990:SSU

[SAS90v] SAS Institute. SAS/STAT user’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, fourth edition, 1990. ISBN 1-55544-376-1. LCCN QA276.4 .S237 1990. Includes bibliographical references and indexes.

SAS:1990:SIG

[SAS90w] SAS Institute. System 2000 introductory guide for SAS software users. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-177-7. xii + 142 pp. LCCN QA276.4 .S96 1990.

SAS:1990:USS

[SAS90x] SAS Institute. Using SAS screen control language in release 6.06. SAS technical report; P-199. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-387-7. xxiv + 211 pp. LCCN CM.1.0.13.

SAS:1991:CEB

[SAS91a] SAS Institute. Changes and enhancements to base SAS software: release 6.07. SAS technical report; P-222. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-466-0. xxxvii + 344 pp. LCCN QA276.4.C483 1991. Includes bibliographical references and index.

SAS:1991:GSSb

[SAS91b] SAS Institute. Getting started with SAS/CALC spreadsheet applications, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-455-5. xx + 224 pp. LCCN QA276.4.G47 1991. Includes bibliographies and index.

SAS:1991:GSSa

[SAS91c] SAS Institute. Getting started with the SAS System using SAS/ASSIST software: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-435-0. xiv + 196 pp. LCCN QA276.4 .G48 1991. Includes index. Videography (p. 181-182).

SAS:1991:GSNa

[SAS91d] SAS Institute. A Guide to the SAS Notes, Sample Library, and online customer support facility for the UNIX environment and derivatives. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-434-2. iv + 34 pp. LCCN QA276.4 .G85 1991.

SAS:1991:GSNb

[SAS91e] SAS Institute. A Guide to the SAS Notes, Sample Library, SUPPORT application, and online customer support facility. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-474-1. iv + 36 pp. LCCN QA276.4 .G86 1991.

SAS:1991:ISS

[SAS91f] SAS Institute. Introducing the SAS System: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-449-0. v + 83 pp. LCCN QA276.4 .I59 1991. Includes index.

SAS:1991:MISa

[SAS91g] SAS Institute. Master index to SAS System documentation. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1991. ISBN 1-55544-423-7. vii + 606 pp. LCCN QA276.4 .M275 1991.

SAS:1991:MISb

[SAS91h] SAS Institute. Master index to SUGI proceedings. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-440-7. vi + 209 pp. LCCN QA276.4 .M276 1991.

SAS:1991:SLP

[SAS91i] SAS Institute. SAS language and procedures: usage 2, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-445-8. xxvi + 649 pp. LCCN QA276.4 S2318 1991. Includes bibliographical references and index.

SAS:1991:SSCa

[SAS91j] SAS Institute. SAS Screen Control Language: usage: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-444-X. xxxvi + 690 pp. LCCN QA276.4 .S2337 1991.

SAS:1991:SAIa

[SAS91k] SAS Institute. SAS/ACCESS interface to IMS-DL/I: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-456-3. 251 pp. LCCN QA276.4 .S2378 1991.

SAS:1991:SAIb

[SAS91l] SAS Institute. SAS/ACCESS interface to INGRES: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-463-6. xxii + 204 pp. LCCN QA276.4 .S2379 1991.

SAS:1991:SAS

[SAS91m] SAS Institute. SAS/ACCESS software: changes and enhancements, release 6.07. SAS technical report; P-221. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-465-2. 122 pp. LCCN QA276.4.S249 1991. Includes bibliographical references and index.

SAS:1991:SSS

[SAS91n]  SAS Institute.  SAS/AF

software, SAS/FSP software, and SAS screen control language:changes and enhancements: release 6.07. SAS technical report; P-216. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-468-7. xiv + 313 pp. LCCN QA276.4.S23144 1991. On t.p. the registered trademark symbol “R” is superscript following “SAS/AF”, “SAS/FSP”, and “SAS” in the title. Includes bibliographical references and index.

SAS:1991:SCCb

[SAS91o] SAS Institute. SAS/C CICS user’s guide: release 5.00. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-442-3. xviii + 134 pp. LCCN QA276.4 .S2384 1991.

SAS:1991:SCCa

[SAS91p] SAS Institute. SAS/C compiler and library user’s guide: release 5.00. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1991. ISBN 1-55544-447-4. xx + 442 pp. LCCN QA276.4 .S2386 1991.

SAS:1991:SSU

[SAS91q] SAS Institute. SAS/CALC software: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-455-5. xlii + 680 pp. LCCN QA276.4.S2375 1991. On t.p. the registered trademark symbol “TM” is superscript following “SAS/CALC” in the title. Includes bibliographical references (p. xxxivxxxv) and index.

SAS:1991:SGSb

[SAS91r] SAS Institute. SAS/GRAPH software, changes and enhancements: release 6.07. SAS technical report; P-215. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-461-X. xii + 67 pp. LCCN QA276.4.S233 1991. Includes bibliographical references and index.

SAS:1991:SSCb

[SAS91s] SAS Institute. SAS/GRAPH software: changes and enhancements to map data sets, release. SAS technical report; P-208. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-437-7. 98 pp. LCCN QA276.4.S25 1991. Includes bibliographical references and index.

SAS:1991:SSG

[SAS91t] SAS Institute. SAS/GRAPH software: graphics editor. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-419-9. xx + 174 pp. LCCN QA276.4 .S2228 1991. “56023”–Cover. Includes index.

SAS:1991:SGSa

[SAS91u] SAS Institute. SAS/GRAPH software: usage: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-421-0. xliv + 887 pp. LCCN QA276.4 .S2394 1991.

SAS:1991:SIS

[SAS91v] SAS Institute. SAS/IML software: changes and enhancements, release 6.07. SAS technical report; P-230. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-478-4. 53 + [1] pp. LCCN QA276.4.S2339 1991. Includes bibliographical references and indexes.

SAS:1991:STS

[SAS91w] SAS Institute. SAS/TOOLKIT software: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-453-9. xxix + 761 pp. LCCN QA276.4 .S2396 1991.

SAS:1991:SAG

[SAS91x] SAS Institute. System administrator’s guide to the SAS System, release 6.07, for the UNIX environment and derivatives. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-436-9. x + 93 pp. LCCN QA276.4 .S97 1991.

SAS:1992:DMS

[SAS92a] SAS Institute. Doing more with SAS/ASSIST software: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-502-0. xxiv + 368 pp. LCCN QA276.4 .D65 1992.

SAS:1992:MIS

[SAS92b] SAS Institute. Master index to SAS System documentation. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, fourth edition, 1992. ISBN 1-55544-503-9. x + 1074 pp. LCCN QA276.4 .M275 1992.

SAS:1992:PPD

[SAS92c]  SAS Institute.  Planning,

preparing, documenting, and referencing SAS products. Gaithersburg, MD (P.O. Box 6015, Gaithersburg 20884-6015), 1992. 49 pp.

SAS:1992:SCV

[SAS92d] SAS Institute. SAS companion for the VSE environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-525-X. xxviii + 558 pp. LCCN QA276.4 .S2223 1992.

SAS:1992:SGV

[SAS92e] SAS Institute. SAS guide to VSAM processing. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-504-7. xiv + 137 pp. LCCN QA276.4 .S2243 1993.

SAS:1992:SSC

[SAS92f] SAS Institute. SAS Screen Control Language: syntax: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-488-1. ix + 124 pp. LCCN QA276.4 .S2336 1992.

SAS:1992:SSSb

[SAS92g] SAS Institute. SAS software:

summary of changes and enhancements. SAS technical report; P-231. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-479-2. xvi + 107 pp. LCCN QA276.4 .S23384 1992. “59148”–Cover. Includes bibliographical references (p. 82–83) and index.

SAS:1992:SAI

[SAS92h] SAS Institute. SAS/ACCESS interface to AS/400 data. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-514-4. xiv + 98 pp. LCCN QA276.4 .S214 1992.

SAS:1992:SSU

[SAS92i] SAS Institute. SAS/LAB software user’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-486-5. xiv + 291 pp. LCCN QA276.4.S2314 1992. Spine title: SAS/LAB software user’s guide. Includes bibliographical references and index.

SAS:1992:SSSa

[SAS92j] SAS Institute. SAS/STAT software: changes and enhancements: release 6.07. SAS technical report; P-229. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1992. ISBN 1-55544-473-3. 620 pp. LCCN QA276.4 S26 1992. Includes indexes.

SAS:1993:CESb

[SAS93a] SAS Institute. Changes and enhancements to the SAS system for the OpenVMS environment: release 6.09. SAS technical report; P-251. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-568-3. xiv + 22 pp. LCCN ????

SAS:1993:CESa

[SAS93b] SAS Institute. Changes and enhancements to the SAS/C compiler, library, and debugger: release 5.50. SAS technical report; C-112. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-541-1. viii + 335 pp. LCCN ????

SAS:1993:CAE

[SAS93c] SAS Institute. Conjoint analysis examples. SAS technical report; P-109. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-581-0. 86 pp. LCCN ????

SAS:1993:GSS

[SAS93d] SAS Institute. Getting started with SAS/EIS software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-546-2. 64 pp. LCCN QA276.4.G473 1993. Includes index.

SAS:1993:IIS

[SAS93e] SAS Institute. Installation instructions for the SAS system, under Microsoft Windows: Release 6.08. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. 35 pp.

SAS:1993:ICS

[SAS93f] SAS Institute. International character support for UNIX environments: release 6.09. SAS technical report; P-257. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-569-1. x + 22 pp. LCCN ????

SAS:1993:PMU

[SAS93g] SAS Institute. Project management using the SAS System: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-557-8. 159 pp. LCCN ????

SAS:1993:SCMa

[SAS93h] SAS Institute. SAS companion for the Microsoft Windows environment: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-527-6. xxii + 356 pp. LCCN ????

SAS:1993:SCMc

[SAS93i] SAS Institute. SAS companion for the Microsoft Windows environment: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN ???? xxii + 566 pp. LCCN QA276.4 .S22147 1993.

SAS:1993:SCMb

[SAS93j] SAS Institute. SAS companion for the Microsoft Windows NT environment: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-562-4. xxiv + 382 pp. LCCN ????

SAS:1993:SCO

[SAS93k] SAS Institute. SAS companion for the OS/2 environment: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1993. ISBN 1-55544-517-9. xx + 318 pp. LCCN ????

SAS:1993:SCUb

[SAS93l] SAS Institute. SAS companion for the UNIX environments: language: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-565-9. xviii + 256 pp. LCCN ????

SAS:1993:SCUc

[SAS93m] SAS Institute. SAS companion for UNIX environments: user interfaces: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-560-8. xiv + 156 pp. LCCN ????

SAS:1993:SSCa

[SAS93n]  SAS Institute.  SAS soft-

ware: changes and enhancements. SAS technical report; P-252. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-567-5. xiv + 63 pp. LCCN ????

SAS:1993:SSCb

[SAS93o] SAS Institute. SAS software:

changes and enhancements, release 6.08. SAS technical report; P-242. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-522-5. xvi + 211 pp. LCCN QA276.4 .S233827 1993. Includes index.

SAS:1993:SSI

[SAS93p] SAS Institute. The SAS System: information delivery for banking and financial services. SAS series for business and industry. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-537-3. vii + 316 pp. LCCN HG1709 .S27 1993.

SAS:1993:SAIa

[SAS93q] SAS Institute. SAS/ACCESS interface to AS/400 data: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-530-6 (paperback). xxii + 213 pp. LCCN ????

SAS:1993:SAIb

[SAS93r] SAS Institute. SAS/ACCESS interface to AS/400 data: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-530-6. xxii + 213 pp. LCCN QA276.4 .S2144 1993.

SAS:1993:SAIe

[SAS93s] SAS Institute. SAS/ACCESS interface to AS/400 data: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-530-6. xxii + 213 pp. LCCN QA276.4 .S2144 1993.

SAS:1993:SAIc

[SAS93t] SAS Institute. SAS/ACCESS interface to ODBC: SQL procedure pass-through facility. SAS technical report; P-262. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-599-3. xii + 32 pp. LCCN ????

SAS:1993:SAIf

[SAS93u] SAS Institute. SAS/ACCESS interface to ORACLE: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1993. ISBN 1-55544-580-2. xxii + 261 pp. LCCN ????

SAS:1993:SAId

[SAS93v] SAS Institute. SAS/ACCESS interface to OS/2 database manager: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-529-2. xxii + 230 pp. LCCN QA276.4 .S2158 1993.

SAS:1993:SAIg

[SAS93w] SAS Institute. SAS/ACCESS interface to OS/2 database manager: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-529-2. xxii + 230 pp. LCCN ????

SAS:1993:SAIh

[SAS93x] SAS Institute. SAS/ACCESS interface to PC file formats, usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-531-4. xxiv + 292 pp. LCCN QA276.4 .S2163 1993.

SAS:1993:SAIi

[SAS93y] SAS Institute. SAS/ACCESS interface to SYBASE and SQL Server: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-552-7. xxii + 258 pp. LCCN ????

SAS:1993:SAS

[SAS93z] SAS Institute. SAS/AF software: FRAME entry usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-477-6. xxvi + 597 pp. LCCN ????

SAS:1993:SSF

[SAS93-27] SAS Institute. SAS/AF software: FRAME entry usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-477-6. xxvi + 597 pp. LCCN QA276.4.S2373 1993. Includes index.

SAS:1993:SCSb

[SAS93-28] SAS Institute. SAS/CONNECT software: changes and enhancements: release 6.09. SAS technical report; P-255. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-566-7. xviii + 109 pp. LCCN ????

SAS:1993:SCSa

[SAS93-29] SAS Institute. SAS/CPE software for open systems: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-574-8. xviii + 274 pp. LCCN ????

SAS:1993:SESa

[SAS93-30] SAS Institute. SAS/EIS software: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-470-9. xxiv + 340 pp. LCCN ????

SAS:1993:SSR

[SAS93-31] SAS Institute. SAS/EIS software: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-470-9. xxiv + 340 pp. LCCN QA276.4.S2225 1993. Includes bibliographical references and index.

SAS:1993:SESb

[SAS93-32] SAS Institute. SAS/ENGLISH software: knowledge base administrator’s guide, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-571-3. xix + 354 pp. LCCN ????

SAS:1993:SSK

[SAS93-33] SAS Institute. SAS/ENGLISH software: knowledge base administrator’s guide, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-571-3. xix + 354 pp. LCCN QA76.9.N38 S36 1993. On t.p. patent registration symbol “R” is superscript following “SAS/ ENGLISH” in title. Spine title: SAS/ENGLISH software. Includes bibliographical references and index.

SAS:1993:SESc

[SAS93-34] SAS Institute. SAS/ENGLISH software: user’s guide, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-570-5. xiv + 151 pp. LCCN ????

SAS:1993:SESd

[SAS93-35] SAS Institute. SAS/ETS software: applications guide 2: econometric modeling, simulation, and forecasting: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-549-7. xxii + 429 pp. LCCN ????

SAS:1993:SEU

[SAS93-36] SAS Institute. SAS/ETS user’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1993. ISBN 1-55544-554-3. vi + 1022 pp. LCCN HA30.3.S28 1993. On t.p. the patent registration symbol “R” is superscript following “SAS/ ETS” in the title. Includes bibliographical references and indexes.

SAS:1993:SGSa

[SAS93-37] SAS Institute. SAS/GRAPH software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-500-4 (paperback). xvi + 295 pp. LCCN ????

SAS:1993:SGSb

[SAS93-38] SAS Institute. SAS/GRAPH software: using graphics devices in the VMS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 6 edition, 1993. ISBN 1-55544-558-6. xvi + 179 pp. LCCN ????

SAS:1993:SIU

[SAS93-39] SAS Institute. SAS/INSIGHT user’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1993. ISBN 1-55544-526-8. vi + 490 pp. LCCN QA276.4.S228 1993.

Includes bibliographical references and index.

SAS:1993:SSPa

[SAS93-40] SAS Institute. SAS/OR software: project management examples, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-544-6. v + 219 pp. LCCN ????

SAS:1993:SSPb

[SAS93-41] SAS Institute. SAS/OR software: the PROJMAN menu system, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-543-8. v + 203 pp. LCCN ????

SAS:1993:SOU

[SAS93-42] SAS Institute. SAS/OR user’s guide: project management, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-528-4. xvi + 597 pp. LCCN QA276.4.S23485 1993. “56137”–Cover. Includes bibliographical references and index.

SAS:1993:SUG

[SAS93-43] SAS Institute. SAS/OR user’s guide: project management, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-528-4. xvi + 597 pp. LCCN ????

SAS:1993:SPS

[SAS93-44] SAS Institute. SAS/PH-Clinical software: administrator’s reference, changes, and enhancements: releases 6.08 and 6.09. SAS technical report; P-259. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-573-X. xiii + 178 pp. LCCN ????

SAS:1993:SSSa

[SAS93-45] SAS Institute. SAS/SHARE software for the CMS environment: release 6.08. SAS technical report; P-261. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-578-0. xiv + 34 pp. LCCN ????

SAS:1993:SSSb

[SAS93-46] SAS Institute. SAS/SHARE software for the MVS environment: release 6.08. SAS technical report; P-260. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-577-2. xii + 36 pp. LCCN ????

SAS:1993:SSSc

[SAS93-47] SAS Institute. SAS/STAT software: Syntax, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-545-4. viii + 151 pp. LCCN ????

SAS:1993:SSSd

[SAS93-48] SAS Institute. SAS/STAT software: the GENMOD procedure: release 6.09. SAS technical report; P-243. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-547-0. 88 pp. LCCN QA276.4.S23397 1993. Includes bibliographical references and index.

SAS:1993:SSSe

[SAS93-49] SAS Institute. SAS/STAT software: the MODECLUS procedure: release 6.09. SAS technical report; P-256. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-506-3. 66 pp. LCCN ????

SAS:1993:UMO

[SAS93-50] SAS Institute. User-written methods and objects in SAS/ EIS software: release 6.08. SAS technical report; P-253. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-559-4. vi + 70 pp. LCCN ????

SAS:1993:URP

[SAS93-51] SAS Institute. Using the RE-PORT procedure in a nonwindowing environment: release 6.07. SAS technical report; P-258. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-572-1. xviii + 276 pp. LCCN ????

SAS:1993:USQ

[SAS93-52] SAS Institute. Using the SQL query window: release 6.08. SAS technical report; P-254. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-563-2. xvi + 134 pp. LCCN ????

SAS:1994:ESE

[SAS94a] SAS Institute. Extending SAS/ EIS software capabilities: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-635-3. vi + 136 pp. LCCN ????

SAS:1994:GSS

[SAS94b] SAS Institute. Getting started with SAS/ACCESS software: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-601-9. 79 pp. LCCN ????

SAS:1994:HSS

[SAS94c] SAS Institute. Highlighting SAS/SPECTRAVIEW software: Version 6 First edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-607-8. 74 pp. LCCN ????

SAS:1994:IES

[SAS94d]  SAS Institute.  Image ex-

tensions to SAS/GRAPH software. SAS technical report; P-263. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-629-9. xii + 134 pp. LCCN ????

SAS:1994:IMR

[SAS94e] SAS Institute. Introduction to market research using the SAS System. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-622-1. xii + 180 pp. LCCN ????

SAS:1994:JIG

[SAS94f] SAS Institute. JMP introductory guide: version 3; statistical software for the Macintosh from SAS Institute Inc. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, January 1994. ISBN 1-55544-588-8. v + 147 pp. LCCN ????

SAS:1994:JSG

[SAS94g] SAS Institute. JMP statistics and graphics guide: version 3; statistical software for the Macintosh from SAS Institute Inc. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-590-X. ix + 580 pp. LCCN ????

SAS:1994:MWE

[SAS94h] SAS Institute. Microsoft Windows environment: changes and enhancements to the SAS system: release 6.10. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-632-9. xiv + 76 pp. LCCN ????

SAS:1994:OEC

[SAS94i] SAS Institute. OS/2 environment: changes and enhancements to the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-633-7. xii + 62 pp. LCCN ????

SAS:1994:RFS

[SAS94j] SAS Institute. Reporting from the field: SAS software experts present real-world reportwriting applications. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-644-2. v + 329 pp. LCCN ????

SAS:1994:SCG

[SAS94k] SAS Institute. SAS consultant’s guide: supporting the SAS System. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-592-6. x + 72 pp. LCCN ????

SAS:1994:SGR

[SAS94l] SAS Institute. SAS guide to report writing: examples, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-637-X. viii + 208 pp. LCCN ????

SAS:1994:SMF

[SAS94m] SAS Institute. SAS macro facility tips and techniques: Version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-605-1. xii + 320 pp. LCCN ????

SAS:1994:SSCa

[SAS94n] SAS Institute. SAS Screen Control Language, reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1994. ISBN 1-55544-659-0. xiv + 648 pp. LCCN ????

SAS:1994:SSA

[SAS94o] SAS Institute. SAS software:

abridged reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-582-9. xvi + 607 pp. LCCN ????

SAS:1994:SSCb

[SAS94p]  SAS Institute.  SAS soft-

ware: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-638-8. xvi + 160 pp. LCCN ????

SAS:1994:SAIa

[SAS94q] SAS Institute. SAS/ACCESS interface to Informix technical report: SQL procedure passthrough facility: release 6.09. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-626-4. 31 pp. LCCN ????

SAS:1994:SAIb

[SAS94r] SAS Institute. SAS/ACCESS interface to PC file formats: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-639-6. x + 95 pp. LCCN ????

SAS:1994:SASa

[SAS94s] SAS Institute. SAS/ACCESS software changes and enhancements: SQL procedure passthrough facility: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-660-4. viii + 44 pp. LCCN ????

SAS:1994:SASb

[SAS94t] SAS Institute. SAS/ACCESS software for relational databases: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-631-0. various pp. LCCN ????

SAS:1994:SASc

[SAS94u] SAS Institute. SAS/ASSIST software: query and reporting from DB2. SAS technical report; P-264. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-652-3. xii + 138 pp. LCCN ????

SAS:1994:SCT

[SAS94v] SAS Institute. SAS/C technical report: SAS/C cross-platform compiler, usage, and reference: release 5.50. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-624-8. x + 144 pp. LCCN ????

SAS:1994:SCSa

[SAS94w] SAS Institute. SAS/CONNECT software, usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1994. ISBN 1-55544-625-6. xii + 368 pp. LCCN ????

SAS:1994:SCSb

[SAS94x] SAS Institute. SAS/CPE software for the MVS environment: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-591-8. xviii + 331 pp. LCCN ????

SAS:1994:SES

[SAS94y] SAS Institute. SAS/ETS software: syntax: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-628-0. iv + 105 pp. LCCN ????

SAS:1994:SGS

[SAS94z] SAS Institute. SAS/GRAPH software. The Encore series Encore series (Cary, N.C.). SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-653-1. vi + 159 pp. LCCN ????

SAS:1994:SIS

[SAS94-27] SAS Institute. SAS/INSIGHT software: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-636-1. 204 pp. LCCN ????

SAS:1994:SQS

[SAS94-28] SAS Institute. SAS/QC software: ADX menu system for design of experiments, version 6. Getting started series Getting started series (Cary, N.C.). SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-443-1. xii + 149 pp. LCCN ????

SAS:1994:SSSa

[SAS94-29] SAS Institute. SAS/SHARE software for the OpenVMS environment. SAS technical report; P-265. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-651-5. x + 60 pp. LCCN ????

SAS:1994:SSTa

[SAS94-30] SAS Institute. SAS/SHARE technical report for the Microsoft Windows environment: release 6.10. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-650-7. 55 pp. LCCN ????

SAS:1994:SSTb

[SAS94-31] SAS Institute. SAS/SHARE technical report for the OS/2 environment: Release 6.08. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-610-8. 58 pp. LCCN ????

SAS:1994:SSTc

[SAS94-32] SAS Institute. SAS/SHARE technical report for UNIX environments: release 6.09. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-609-4. 46 pp. LCCN ????

SAS:1994:SSSb

[SAS94-33] SAS Institute. SAS/STAT software: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-627-2. vi + 121 pp. LCCN ????

SAS:1994:STF

[SAS94-34] SAS Institute. SAS/TUTOR:

fundamentals of the SAS system: course guide: release 6.08: OS/2, Windows; release 6.09: OpenVMS for AXP, Windows NT, UNIX Environments; release 6.10: Microsoft Windows Environments, OS/2, UNIX environments. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. xvi + 131 + 1 pp.

SAS:1994:SMA

[SAS94-35] SAS Institute. Stock market analysis using the SAS System: portfolio selection and evaluation. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-623-X. x + 243 pp. LCCN ????

SAS:1994:BSU

[SAS94-36] SAS Institute Inc. Basic statistics using SAS/STAT software. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-602-7. xii + 720 pp. LCCN ????

SAS:1995:BYC

[SAS95a] SAS Institute. Building your client/server solution with the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995.

ISBN 1-55544-217-X. viii + 263 pp. LCCN ????

SAS:1995:CSC

[SAS95b] SAS Institute. Client/server computing with the SAS system: tips and techniques. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-218-8. x + 348 pp. LCCN ????

SAS:1995:CMS

[SAS95c] SAS Institute. Combining and modifying SAS data sets: examples: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-220-X. x + 197 pp. LCCN ????

SAS:1995:DIA

[SAS95d]  SAS Institute.  Developing

ISPF applications with the SAS system: release 6.08. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-223-4. 39 pp. LCCN ????

SAS:1995:GSS

[SAS95e] SAS Institute. Getting started with the SAS System in the MVS environment. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, version 6 edition, 1995. ISBN 1-55544-224-2. vi + 65 pp. LCCN ????

SAS:1995:LRE

[SAS95f] SAS Institute. Logistic regression examples using the SAS system: Version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-674-4. viii + 163 pp. LCCN QA276.4.L64 1995.

SAS:1995:MWE

[SAS95g] SAS Institute. Microsoft Windows environment: changes and enhancements to the SAS system: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-271-4. xiv + 108 pp. LCCN ????

SAS:1995:OEC

[SAS95h] SAS Institute. OS/2 environment: changes and enhancements to the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-270-6. x + 49 pp. LCCN ????

SAS:1995:SGR

[SAS95i] SAS Institute. SAS guide to the REPORT procedure: reference. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-265-X. vi + 123 pp. LCCN ????

SAS:1995:SGSa

[SAS95j] SAS Institute. SAS guide to the SQL Query Window: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-266-8. iv + 167 pp. LCCN ????

SAS:1995:SSC

[SAS95k]  SAS Institute.  SAS soft-

ware: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-253-6. xviii + 288 pp. LCCN ????

SAS:1995:SAIa

[SAS95l] SAS Institute. SAS/ACCESS interface to CA-IDMS: reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-661-2. iv + 82 pp. LCCN ????

SAS:1995:SAIb

[SAS95m] SAS Institute. SAS/ACCESS interface to IMS-DL/I: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1995. ISBN 1-55544-227-7. viii + 365 pp. LCCN ????

SAS:1995:SASa

[SAS95n] SAS Institute. SAS/ACCESS software for PC file formats: reference. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-212-9. various pp. LCCN ????

SAS:1995:SASb

[SAS95o] SAS Institute. SAS/AF software: FRAME application development concepts, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-243-9. xii + 200 pp. LCCN ????

SAS:1995:SASc

[SAS95p] SAS Institute. SAS/AF software: FRAME class dictionary, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-206-4. various pp. LCCN ????

SAS:1995:SASd

[SAS95q] SAS Institute. SAS/ASSIST software: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-216-1. v + 322 pp. LCCN ????

SAS:1995:SESa

[SAS95r] SAS Institute. SAS/EIS software: reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1995. ISBN 1-55544-205-6. various pp. LCCN ????

SAS:1995:SET

[SAS95s] SAS Institute. SAS/EIS technical report: control chart object. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-263-3. vi + 12 pp. LCCN ????

SAS:1995:SESb

[SAS95t] SAS Institute. SAS/ETS software: changes and enhancements: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-273-0. 113 pp. LCCN ????

SAS:1995:SESc

[SAS95u] SAS Institute. SAS/ETS software: time series forecasting system: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-275-7 (paperback). vii + 254 pp. LCCN ????

SAS:1995:SGSb

[SAS95v] SAS Institute. SAS/GIS software: usage and references: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-561-6. x + 261 pp. LCCN ????

SAS:1995:SIS

[SAS95w] SAS Institute. SAS/IML software: changes and enhancements, release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-281-1. 302 pp. LCCN ????

SAS:1995:SIU

[SAS95x] SAS Institute. SAS/INSIGHT user’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, third edition, 1995. ISBN 1-55544-751-1. 576 pp. LCCN ????

SAS:1995:SQSb

[SAS95y] SAS Institute. SAS/QC software: SQC Menu System for quality improvement. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1995. viii + 188 pp.

SAS:1995:SQSa

[SAS95z] SAS Institute. SAS/QC software: usage and reference: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-658-2 (set paperback). xxiv + 1697 pp. LCCN ???? Two volumes.

SAS:1995:SSS

[SAS95-27] SAS Institute. SAS/SHARE software for the VSE environment: release 6.08. SAS technical report; P-267. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-228-5. x + 60 pp. LCCN ????

SAS:1995:SSTa

[SAS95-28] SAS Institute. SAS/SHARE technical report for the Microsoft Windows environment: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-260-9. x + 75 pp. LCCN ????

SAS:1995:SSTb

[SAS95-29] SAS Institute. SAS/SHARE technical report for the OS/2 environment: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-262-5. viii + 71 pp. LCCN ????

SAS:1995:SSTc

[SAS95-30] SAS Institute. SAS/SHARE technical report for UNIX environments: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-261-7. x + 73 pp. LCCN ????

SAS:1995:STS

[SAS95-31] SAS Institute. SAS/TRADER software, product administrator’s guide: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-221-8. vi + 104 pp. LCCN ????

SAS:1995:SMA

[SAS95-32] SAS Institute. Stock market analysis using the SAS system: technical analysis. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-222-6. vi + 187 pp. LCCN ????

SAS:1995:SCM

[SAS95-33] SAS Institute, Inc. SAS companion for the Macintosh, Version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. iv + 109 pp.

SAS:1996:ESI

[SAS96a] Execsolutions: SAS Institute’s magazine for business decision makers, page various, 1996. ISSN 1086-167X. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA.

SAS:1996:GSS

[SAS96b] SAS Institute. Getting started with the SAS System using SAS/ASSIST software: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1996. ISBN 1-55544-756-2. vi + 175 pp. LCCN ????

SAS:1996:STR

[SAS96c] SAS Institute. SAS technical report P-269, SAS/ACCESS interface to CA-IDMS: DATA step interface: 6.09 enhanced release. SAS technical report: P-269. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN 1-55544-776-7. viii + 82 pp. LCCN ????

SAS:1996:SSS

[SAS96d] SAS Institute. SAS/STAT software: changes and enhancements. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1996. ISBN 1-55544-274-9. viii + 1094 pp. LCCN ????

Saur:2004:RSM

[Sau04a] Joe Saur. Review of SASÒ for Monte Carlo studies: a guide for quantitative researchers by Xitao Fan, Ákos Felsvlyi, Stephen A. Sivo, and Sean C. Keenan. SAS Institute, Inc. 2002. ACM SIGSOFT Software Engineering Notes, 29 (5):39, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Saur:2004:RVC

[Sau04b] Joe Saur. Review of Visualizing categorical data by Michael Friendly. SAS Institute, Inc. 2000. ACM SIGSOFT Software Engineering Notes, 29(5): 39–40, September 2004. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Sauter:2011:BRB

[Sau11] Roger M. Sauter. Book review: R for SAS and SPSS Users (1st ed) by Robert A. Muenchen. Technometrics, 53 (1):101, February 2011. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/40997301.

Scheuchenpflug:1996:CCU

[SB96] Thomas Scheuchenpflug and Maria Blettner. Coding confusion using PROC LOGIS-TIC in SAS^{(}R). Computational Statistics & Data Analysis, 21(1):111–115, January 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0167947396900055.

Somerville:2001:FSI

[SB01] Paul N. Somerville and Frank Bretz. FORTRAN 90 and SAS-IML programs for computation of critical values for multiple testing and simultaneous confidence intervals. Journal of Statistical Software, 6(5):1– 14, 2001. CODEN JSSOBK. ISSN ???? URL http://

www.jstatsoft.org/v06/i05;
 http://www.jstatsoft.org/
v06/i05/EXAMPLE/;   http:
//www.jstatsoft.org/v06/
i05/FORTRAN/;   http://www.
jstatsoft.org/v06/i05/jss2r.
pdf;   http://www.jstatsoft.
org/v06/i05/SAS/;   http:
//www.jstatsoft.org/v06/
i05/updates.

Smith:1985:DSM

[SBJ85] L. D. Smith, John Blodgett, and Marius Janson. Decision support for marketing research and corporate planning. Inf. Manage., 8(3):133–145, March 1985. From Computing Reviews: “The authors describe, in a well-written “here is how we did it” report, their experience in using a Statistical Analysis System (SAS) to construct a decision support system for use by three people in an insurance company.”.

Shoukri:2007:ACD

[SC07] M. M. (Mohamed M.) Shoukri and Mohammad A. Chaudhary. Analysis of correlated data with SAS and R. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2007. ISBN 1-58488-619-6. 295 pp. LCCN RA652.2.M3 S53

2007. URL http://www.loc.

                                        gov/catdir/enhancements/
                                        fy0733/2007000593-d.htm;
                                        http://www.loc.gov/catdir/
                                        toc/ecip078/2007000593.htm.

Schill:2006:ALL

[Sch06a]  Walter Schill.  Advanced

log–linear models using SAS. Biometrics, 63(2):949–950, September 2006. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2006:BRA

[Sch06b] Walter Schill. Book review: Advanced Log–Linear Models Using SAS, by D. Zelterman, SAS Institute, Cary, North Carolina, 2002. vii + 189 pp . US$34.95, ISBN 1-59047-080-X. Biometrics, 63(2):949– 950, September 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schill:2006:BRC

[Sch06c] Walter Schill. Book review: Categorical Data Analysis Using the SAS System, by M. E. Stokes, C. S. Davis and G. G. Koch. Biometrics, 62(1): 302–303, March 2006. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stokes:1995:CDA

[SDK95] Maura Ellen Stokes, Charles S. Davis, and Gary G. (Gary Grove) Koch. Categorical data analysis using the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-219-6. iv + 499 pp. LCCN ????

Searle:1980:ACOe

[Sea80] Shayle R. Searle. Annotated computer output for variance components (ACO [sigma]2): SAS HARVEY. BU; 723-M. Biometrics Unit, Cornell University, Ithaca, NY, USA, rev edition, 1980. ix + 35 pp.

Seaver:2011:BRBa

[Sea11] William Seaver. Book review: A SAS/IML Companion for Linear Models by Jamis J. Perret. Technometrics, 53(1): 107, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/40997311.

Seaver:2012:BRB

[Sea12] W. L. Seaver. Book review:

Statistical Programming with SAS/IML Software by Rick Wicklin. Technometrics, 54(2): 201–202, May 2012. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/41714874.

Schill:2014:SPL

[SED14] Walter Schill, Dirk Enders, and Karsten Drescher. A SAS package for logistic two-phase studies. Journal of Statistical Software, 57(9):??, April 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v57/i09.

Selukar:2011:SSM

[Sel11] Rajesh Selukar. State space modeling using SAS. Journal of Statistical Software, 41(12):??, May 2011. CODEN JSSOBK. ISSN 1548-7660. URL http://

                                        www.jstatsoft.org/v41/i12.

SCH:1993:IAL

[Ser93a] Services conseils Hardy. Introduction au logiciel SAS: session de formation. Les Services, Sainte-Foy, PQ, Canada, 1993. ISBN 2-921623-02-1. 199 pp. LCCN ????

SCH:1993:DMS

[Ser93b] Services conseils Hardy. Le Display Manager de SAS en version 6: session de formation. Les Services, Sainte-Foy, PQ, Canada, 1993. ISBN 2-921623-03-X. 63 pp. LCCN ????

SCH:1993:MVV

[Ser93c] Services conseils Hardy. Migration vers la version 6.07 de SAS: session de formation. Les Services, Sainte-Foy, PQ, Canada, 1993. ISBN 2-921623-01-3. 207 pp. LCCN ????

SCH:1993:SAP

[Ser93d] Services conseils Hardy. SAS/ ACCESS pour DB2: session de formation. Les Services, Sainte-Foy, PQ, Canada, 1993. ISBN 2-921623-00-5. 87 pp. LCCN ????

Searle:1980:ACOa

[SG80] Shayle R. Searle and B. A. Grimes. Annotated computer output for variance components (ACO [sigma]2): SAS VAR-COMP. BU; 708-M. Biometrics Unit, Cornell University, Ithaca, NY, USA, rev edition, 1980. xxi + 41 pp.

Searle:1980:ACOc

[SH80a] Shayle R. Searle and H. V.

Henderson. Annotated computer output for analyses of unbalanced data: SAS GLM. Paper … in the Biometrics Unit; no. BU-641-M. Biometrics Unit, Cornell University, Ithaca, NY, USA, rev edition, 1980. 67 pp.

Searle:1980:ACOd

[SH80b] Shayle R. Searle and H. V.

Henderson. Annotated computer output for analyses of unbalanced data: SAS HARVEY. Paper … in the Biometrics Unit mimeo series; no. BU-659-M. Biometrics Unit, Cornell University, Ithaca, NY, USA, rev edition, 1980. 59 pp.

Searle:1980:ACOb

[SH80c] Shayle R. Searle and G. F. S.

Hudson. Annotated computer output for variance components (ACO [sigma]2): SAS RAN-DOM (and SAS NESTED). Paper … in the Biometrics Unit; no. BU-722-M. Biometrics Unit, Cornell University, Ithaca, NY, USA, rev edition, 1980. xvii + 30 pp.

Siddique:2009:MSM

[SH09] Juned Siddique and Ofer Harel.

MIDAS: a SAS macro for multiple imputation using distanceaided selection of donors. Journal of Statistical Software, 29 (9):??, February 2009. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v29/i09.

Shanmugam:2006:BRS

[Sha06] Ramalingam Shanmugam. Book review selective books on software SAS, SPSS, and Internet outsourcing. Journal of Statistical Computation and Simulation, 76(12):1107–1113, 2006. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2009:BSS

[Sha09] Ramalingam Shanmugam. Books on SAS, SPSS, and computing are reviewed. Journal of Statistical Computation and Simulation, 79(5):747–750, 2009. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2015:BRSa

[Sha15] Ramalingam Shanmugam. Book reviews: SAS books are reviewed. Journal of Statistical Computation and Simulation, 85(3):638–640, 2015. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2020:BRE

[Sha20] Ramalingam Shanmugam. Book review: Essentials of Excel, Excel VBA, SAS, and Minitab for statistical and financial analysis, by Cheng-Few Lee, John Lee, John-Ran Chang and Tzu Tai, Chan, Springer Verlag Press, 2016, 1041 + xx pp., $109.00 (paperback), ISBN: 978-3-319-388865-6. Journal of Statistical Computation and Simulation, 90(1):179–180, 2020. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Shanmugam:2022:BRB

[Sha22] Ramalingam Shanmugam. Book review of three excellent books on R language for statistics:

1. Stochastic processes with R: an introduction, by Olga Korosteleva, 2022, Boca Raton, FL, CRC Press, 190 + ix pp., $93.22 (hardcover), ISBN: 978-1-138-59152-3; 2 Statistical methods for mediation, confounding and moderation analysis using R and SAS, by Qingzhao Yu and Bin Li, 2022, Boca Raton, FL, CRC Press, 277 + xv pp., $144.49 (hardcover), ISBN: 978-0-367-36547-9; 3 Spatial predictive modeling with R, by Jin Li, 2022, Boca Raton, FL, CRC Press, 383 + xix pp., $118.10 (hardcover), ISBN: 978-0-367-55054-7. Journal of Statistical Computation and Simulation, 92(12):2642– 2643, 2022. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Sarwar:1988:IFL

[SHD88] S. M. Sarwar, S. J. Hahn, and J. A. Davis. Implementing functional languages on a combinator-based reduction machine. ACM SIGPLAN Notices, 23(4):65–70, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shin:1996:SGV

[Shi96] Kilman Shin. SAS guide: for DOS version 5.0 and Windows version 6.10. The Irwin statistical software series. Irwin, Chicago, IL, USA, second edition, 1996. ISBN 0-256-20650-3 (Student’s ed.), 0-256-20651-1 (Instructor’s ed.). xi + 406 pp. LCCN ????

Shoesmith:1990:BRB

[Sho90] Eddie Shoesmith. Book review: Mastering the SAS System, by J. A. Jaffe. Journal of the Royal Statistical Society. Series A (Statistics in Society), 153 (2):258–259, ???? 1990. CO-DEN JSSAEF. ISSN 0964-1998 (print), 1467-985X (electronic).

                                        URL http://www.jstor.org/
                                        stable/2982811.

Stribling:1996:RWC

[SHO^{+}96] D. Stribling, T. Hunter, L. Olszewski, A. Corrigan, R. Mullis, and L. Allen. A real world conversion to SGML. In ACM [ACM96], pages 75–86. ISBN 0-89791-799-5. LCCN QA 76.9 D6 I56 1996.

Shoukri:2018:ACD

[Sho18] M. M. (Mohamed M.) Shoukri. Analysis of Correlated Data with SAS and R. Chapman and Hall/CRC, Boca Raton, FL, USA, fourth edition edition, 2018. ISBN 1-138-19745-9 (hardcover), 1-315-27770-0 (Mobipocket e-book), 1-315-27771-9 (ePub e-book), 1-315-27772-7 (PDF e-book), 1-315-27773-5 (e-book). xv + 497 pp. LCCN RA652.2.M3 .S53 2018.

Silk:1996:SEC

[Sil96] Julian Silk. Systems estimation: A comparison of SAS, SHAZAM and TSP. Journal of Applied Econometrics, 11 (4):437–450, July 1996. CO-DEN JAECET. ISSN 0883-7252 (print), 1099-1255 (electronic).

Singh:2008:BRB

[Sin08] Ashok K. Singh. Book review:

Introduction to Data Mining Using SAS Enterprise Miner by Patricia B. Cerrito. Technometrics, 50(4):545, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/25471539.

Schlotzhauer:1987:SSE

[SL87] Sandra D. Schlotzhauer and Ramon C. Littell. SAS system for elementary statistical analysis. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1987. ISBN 1-55544-076-2. xxii + 416 + 6 pp. LCCN QA 276.4 S35 1987. By Sandra D. Schlotzhauer and Ramon C. Littell. Includes bibliographical references and index.

Stewart:1986:SML

[SLMN86] Larry Stewart, Timothy Lehman, Jerry Mock, and Rebeccah K. Neff. SAS macro language. Version 5th edition. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 1-55544-002-9. xiv + 457 pp. LCCN QA276.4 .S73 1986.

Stammer:1982:EAS

[SM82] Robert E. Stammer and Howard A. McKamey. Effective applications of SAS procedures to facilitate transportation planning. Oak Ridge, TN, USA, 1982. 11 pp. Microfiche. Springfield, Va.: National Technical Information Service.

Stroup:2018:SMM

[SMCW18] Walter Stroup, George Milliken, Elizabeth Claassen, and Russell Wolfinger. SAS for Mixed Models. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2018. ISBN 1-63526-135-X, 1-63526-152-X (e-book), 1-63526-154-6. 608 pp. LCCN QA276.4 .L58 2018.

Sauerbrei:2006:MRM

[SMHBR06] W. Sauerbrei, C. Meier-Hirmer, A. Benner, and P. Royston. Multivariable regression model building by using fractional polynomials: Description of SAS, STATA and R programs. Computational Statistics & Data Analysis, 50(12):3464–3485, August 2006. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947305001623.

SCLIHEAP:1994:LHE

[Sou94] South Carolina.Low-income Home Energy Assistance Program. Low-income Home Energy Assistance Program (LI-HEAP): TPFY 1994 emergency cooling assistance program statistical reports. South Carolina Office of the Governor, Division of Economic Opportunity, LIHEAP?, Columbia, SC, 1994. 2 + 47 pp.

Sparapani:2006:BRB

[Spa06] Rodney Sparapani. Book review: Statistical Analysis of Medical Data Using SAS. Journal of Statistical Software, 16 (BR-5):1–5, July 2006. CO-DEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/v16/b05.

Spector:1993:SPR

[Spe93] Paul E. Spector. SAS programming for researchers and social scientists. Sage Publications, Newbury Park, 1993. ISBN 0-8039-4984-7 (cloth), 0-8039-4985-5 (paperback). xii + 167 pp. LCCN QA276.4 .S645 1993. Includes bibliographical references (p. 163) and index.

Spencer:2004:SPO

[Spe04] Neil Spencer. SAS programming: the one-day course. Chapman and Hall/CRC, Boca Raton, FL, USA, 2004. ISBN 1-58488-409-6. 193 pp. LCCN QA276.4 .S6455 2004. US$.

Speed:2008:SPE

[Spe08] F. Michael Speed. SAS programming for enterprise guide users. The American Statistician, 62(4):362, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Spicer:1996:QEA

[Spi96] J. Spicer. Quick and easy AF applications using SAS/ ASSIST software, SAS data views, block menus and HTML. In Anonymous [Ano96a], pages 764–769 (vol. 1). Two volumes.

Smith:2007:WSM

[SR07] Michael K. Smith and Helen Richardson. WinBUGSio: a SAS macro for the remote execution of WinBUGS. Journal of Statistical Software, 23(9):1– 10, December 2007. CODEN JSSOBK. ISSN 1548-7660.

                                        URL http://www.jstatsoft.
                                        org/v23/i09.

Silva:1998:NTG

[SS98] A. Pedro Duarte Silva and Antonie Stam. Nonparametric two-group classification: Concepts and a SAS-based software package. The American Statistician, 52(2):185– 197, May 1998. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://www.jstor.org/
stable/2685479.

Stanish:1986:CDA

[Sta86] William M. Stanish. Categorical data analysis strategies using SAS software. In Allen [All86], pages 239– 256. ISBN 0-444-70018-8 (paperback). LCCN QA276.4 .S95 1985. URL http:

//catalog.hathitrust.org/
api/volumes/oclc/13455449.
html;   http://www.dtic.
mil/dtic/tr/fulltext/u2/
a166019.pdf;   http://zbmath.
org/?q=an:0647.62008.

Stanley:1994:BOS

[Sta94] Don Stanley. Beyond the obvious with SAS Screen Control Language. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-600-0. vi + 347 pp. LCCN ????

Starkings:2007:SBRd

[Sta07a] Susan Starkings. Short book review: Data Preparation for Analytics Using SAS(R) by Gerhard Svolba. International Statistical Review = Revue Internationale de Statistique, 75 (3):429, December 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/
stable/41509899;   http://
www.jstor.org/stable/41509900.

Starkings:2007:SBRg

[Sta07b] Susan Starkings. Short book review: Learning SAS(R) by Example: A Programmer’s Guide by Ron Cody. International Statistical Review = Revue Internationale de Statistique, 75 (3):430, December 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2007:SBRe

[Sta07c] Susan Starkings. Short book review: Pharmaceutical Statistics Using SAS(R): A Practical Guide by Alex Dmitrienko, Christy Chuang-Stein, Ralph D’Agostino. International Statistical Review = Revue Internationale de Statistique, 75(3): 429, December 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Starkings:2007:SBRc

[Sta07d] Susan Starkings. Short book review: Statistics Using SAS(R) Enterprise Guide(R) by James B. Davis. International Statistical Review = Revue Internationale de Statistique, 75(3): 428–429, December 2007. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

                                        URL http://www.jstor.org/
                                        stable/41509898.

Stadtlander:2021:BRA

[Sta21] Christian T. K.-H. Stadtländer. Book review: Analysis of correlated data with SAS and R (4th ed.), by Mohamed M. Shoukri, Boca Raton, FL, CRC Press, 2018, 497 + xv pp., $139.95 (hardback), ISBN: 978-1-1381-9745-9. Journal of Statistical Computation and Simulation, 91(7):1478–1479, 2021. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Stewart:1983:SVS

[Ste83] J. Larry Stewart. SAS views, SAS basics. SAS training series. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1983. ISBN 0-917382-40-4. 754 pp. LCCN QA 276.4 S238 1983. Edited by J. Larry Stewart. Includes index.

Stewart:1986:SB

[Ste86] Larry Stewart. SAS basics. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1986. ISBN 1-55544-004-5. xii + 664 pp. LCCN QA276.4 .S728 1986.

Stewart:1999:BRB

[Ste99] Paul Stewart. Book review:

Linear Mixed Models in Practice: A SAS-Oriented Approach (Lecture Notes in Statistics 126) by Geert Verbeke; Geert Molenberghs. Technometrics, 41(1):79–80, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271007.

Steinley:2006:BRB

[Ste06] Douglas Steinley. Book review: Categorical Data Analysis with SAS and SPSS Applications, by Bayo Lawal; Categorical Data Analysis Using the SAS System, by Maura E. Stokes; Charles S. Davis; Gary G. Koch. Journal of the American Statistical Association, 101 (473):400, March 2006. CO-DEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://www.jstor.org/
stable/30047484.

Steibel:2011:BRS

[Ste11] Juan P. Steibel. Book review: SAS and R Data Management, Statistical Analysis, and Graphics, by Kleinman, K. and Horton, N. Biometrics, 68(3):1179– 1180, September 2011. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stone:1982:SID

[Sto82] Alan A. Stone. SAS/IMS-DL/I version 1 release 2 enhancements for SAS 82.2. SAS technical report; P-122. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. 7 + [7] pp.

SAS:1982:PSA

[SUG82] Proceedings of the Seventh Annual SUGI-SAS Users Group International Conference: San Francisco, California, February 14–17, 1982. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1982. LCCN QA276.4 .S236 1982. Includes bibliographies and indexes.

SAS:1983:PEA

[SUG83] Proceedings of the eighth annual SAS Users Group International (SUGI) Conference, New Orleans, Louisiana, January 16– 19, 1983. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1983. LCCN QA 276 A1 S2 1983. Includes bibliographies and indexes.

SAS:1984:PNA

[SUG84] Proceedings of the ninth annual SAS Users Group International (SUGI) Conference, Hollywood Beach, Florida, March 18–21, 1984. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1984. LCCN QA 276 A1 S2 1984. Includes bibliographies and indexes.

SAS:1985:SPT

[SUG85] SUGI 10: proceedings of the Tenth Annual SAS Users Group International Conference, Reno, Nevada, March 10– 13, 1985. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1985. LCCN QA276.4 .S236 1985.

SAS:1987:SUG

[SUG87] Proceedings of the Asia-Pacific SAS Users Group Conferences: SUNZ ’87, SAS Users of New Zealand, Wellington, New Zealand, October 15–16, 1987: SUGA ’87 SAS Users Group of Australia, Canberra, ACT October 7-9, 1987. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1987. LCCN QA276.4 .S236 1987.

SAS:1989:PFA

[SUG89] Proceedings of the Fourteenth Annual SAS Users Group International Conference, San Francisco, California, April 9– 12, 1989. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. LCCN QA276.4 .S236 1989. Includes bibliographical references.

SAS:1991:SUG

[SUG91] SAS Users Group International. Conference (16th: 1991: New Orleans, LA) Proceedings of the Sixteenth Annual SAS Users Group International Conference, New Orleans, Louisiana, February 17– 20, 1991/SUGI 16; conference chair, Martin J. Rosenberg. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-438-5. LCCN QA276.4 .S236 1991.

Sun:1998:UHF

[Sun98] J. Sun. Update HTML formatted SAS output within Web browser. In Anonymous [Ano98], pages 899–?? ISBN 1-58025-149-8. LCCN ????

Senn:2011:CSM

[SWH^{+}11] Stephen Senn, James Weir, Tsushung A. Hua, Conny Berlin, Michael Branson, and Ekkehard Glimm. Creating a suite of macros for meta-analysis in SAS(R): a case study in collaboration. Statistics & Probability Letters, 81(7):842–851, July 2011. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167715211000484.

Taeger:2014:SHT

[Tae14] Dirk Taeger. Statistical hypothesis testing with SAS and R. John Wiley and Sons, Inc., New York, NY, USA, 2014. ISBN 1-119-95021-X (hardcover), 1-118-76260-6 (PDF), 1-118-76258-4, 1-118-76261-4 (ePub), 1-306-32264-2. 308 (est.) pp. LCCN QA277.

Tam:1994:SMS

[Tam94] Katy K. (Katy King Chi) Tam.

SAS macros for sampling methods. Thesis (m.s.), University of Georgia, Athens, GA, USA, 1994. v + 32 pp. Directed by Nancy I. Lyons.

Teichroew:1985:SDM

[TD85] Daniel Teichroew and Gabor David, editors. System description methodologies: proceedings of the IFIP TC2 Conference on System Description Methodologies, Kecskemet, Hungary, 23– 27 May, 1983. Elsevier Science Publishers, Amsterdam, The Netherlands, 1985. ISBN 0-444-87731-2. LCCN QA76.9.S88 I327 1983.

Terplan:1981:NPR

[Ter81] K. Terplan. Network performance reporting. ACM SIG-METRICS Performance Evaluation Review, 11(1):156–170, Spring 1981. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Thomson:1992:BRB

[Tho92] Steve Thomson. Book review: How SAS Works by Paul Herzberg. Technometrics, 34 (1):117, February 1992. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/1269582.

Tang:2012:ACC

[THT12] Wan. Tang, Hua. He, and Xin M. Tu. Applied categorial and count data analysis. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-0624-1. xx + 363 pp. LCCN ????

Tierie:1988:SPS

[Tie88] Ingrid Tierie. Seugi ’88: proceedings of the SAS European Users Group International Conference, Copenhagen, April 26–29 1988. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. x + 450 pp. Edited and compiled by Ingrid Tierie. Includes bibliographical references.

Tierie:1989:SPS

[Tie89] Ingrid Tierie. SEUGI ’89:

proceedings of the SAS European Users Group International Conference, Cologne, May 9– 12, 1989. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1989. xi + 595 pp. Edited and compiled by Ingrid Tierie. Includes bibliographical references.

Tilanus:1994:WSS

[Til94] Erik W. Tilanus. Working with the SAS system. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-645-0 (paperback). xii + 572 pp. LCCN ????

Tilke:1996:BRBb

[Til96]  Clemens Tilke.  Book re-

view: SAS programming for researchers and social scientists: Paul E. Spector: Sage Publications (1993), 167 p. ISBN 0-8039-4984-7 (cloth), ISBN 0-8039-4985-5 (paperback). Computational Statistics & Data Analysis, 21(5):606, May 1996. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947396901413.

Tilahun:2007:FSM

[TPAM07] Abel Tilahun, Assam Pryseley, Ariel Alonso, and Geert Molenberghs. Flexible surrogate marker evaluation from several randomized clinical trials with continuous endpoints, using R and SAS. Computational Statistics & Data Analysis, 51(9):4152–4163, May 15, 2007. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167947307000679.

Thomas:2014:CST

[TR14] Laine Thomas and Eric M. Reyes. Code snippet: Tutorial: Survival estimation for Cox regression models with timevarying coefficients using SAS and R. Journal of Statistical Software, 61(CS-1):??, October 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v61/c01.

Truppi:1988:CDB

[Tru88] Lawrence E. Truppi. A computer data base from EPA’s complex terrain field studies: description of a SAS computer data base from tracer field studies at Cinder Cone Butte, ID, Hogback Ridge, NM, and the Tracy Power Plant, NV. Atmospheric Sciences Research Laboratory, 1988. 3 pp. Caption title. At head of title: Project summary. Distributed to depository libraries in microfiche.

Turner:2012:BRS

[Tur12] Robin Turner. Book review: SAS and R: Data management, statistical analysis, and graphics. Statistics in Medicine, 31 (13):1405–1406, June 15, 2012. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uyar:1990:RPS

[UE90] Bulent Uyar and Orhan Erdem. Regression procedures in SAS: Problems? The American Statistician, 44(4):296–??, November 1990. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Uranga:2014:LCM

[UM14] R. Uranga and G. Molenberghs.

Longitudinal conditional models with intermittent missingness: SAS code and applications. Journal of Statistical Computation and Simulation, 84(4):753–780, 2014. CO-DEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

USGAOOP:1992:PPD

[Uni92] United States. General Accounting Office. Office of Policy, P.O. Box 6015, Gaithersburg 20884-6015, MD, USA. Planning, preparing, documenting, and referencing SAS products, 1992. 49 pp.

USGAOOP:1993:PPD

[Uni93] United States.General Accounting Office.Office of Policy. Planning, preparing, documenting, and referencing SAS products. ???? GAO/IMTEC-11.1.2, U.S. General Accounting Office, Washington, DC, USA, 1993. 49 pp.

USNHTSA:1994:TSDa

[Uni94a] United States.National Highway Traffic Safety Administration. Traffic safety data, 1988– 1993 a compilation of traffic safety reports and data files, 1994. 1 computer laser optical disc (12 cm.).

USNHTSA:1994:TSDb

[Uni94b] United States.National Highway Traffic Safety Administration. Traffic safety data, 1988– 1993 a compilation of traffic safety reports and data files, 1994. 1 computer laser optical disc (12 cm.).

Utlaut:2005:BRB

[Utl05] Theresa L. Utlaut. Book review: Statistical Analysis and Data Display: An Intermediate Course with Examples in S-PLUS. R, and SAS by Richard M. Heiberger; Burt Holland. Technometrics, 47(3): 371, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/25471029.

VanHentenryck:1997:SAI

[Van97] Pascal Van Hentenryck, editor. Static analysis: 4th International Symposium, SAS’97, Paris, France, September 1997: proceedings, volume 1302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1997. CODEN LNCSD9. ISBN 3-540-63468-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6.I5733 1997. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1302.htm;   http:
//www.springerlink.com/content/
978-3-540-63468-3;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1302.

VanBrackle:2004:BRB

[Van04]  Lewis VanBrackle.  Book

review: Step-by-Step Basic Statistics Using SAS(R): Student Guide by Larry Hatcher. Technometrics, 46(4):494–495, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/25470913.

Vock:2014:CSS

[VDT14] David Vock, Marie Davidian, and Anastasios Tsiatis. Code snippet: SNP NLMM: A SAS macro to implement a flexible random effects density for generalized linear and nonlinear mixed models. Journal of Statistical Software, 56(CS-2): ??, January 2014. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v56/c02.

Vencatasawmy:2001:BRB

[Ven01] Coomaren P. Vencatasawmy.

Book review: Introduction to Time Series Analysis and Forecasting with Applications of SAS and SPSS, by R. Yaffee; M. McGee. Journal of the Royal Statistical Society. Series D (The Statistician), 50 (3):362–363, ???? 2001. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2680962.

Voronca:2018:MSM

[VGDM^{+}18] Delia C. Voronca, Mulugeta Gebregziabher, Valerie Durkalski-Mauldin, Lei Liu, and Leonard E. Egede. MTPmle: A SAS macro and Stata programs for marginalized inference in semi-continuous data. Journal of Statistical Software, 87 (??):??, ???? 2018. CO-DEN JSSOBK. ISSN 1548-

                                        7660.  URL https://www.
                                        jstatsoft.org/index.php/
                                        jss/article/view/v087i06;
                                         https://www.jstatsoft.org/
                                        index.php/jss/article/view/
                                        v087i06/v87i06.pdf.

Voelkel:2009:SBR

[Voe09]  Joseph G. Voelkel.  Short

book review: R for SAS and SPSS Users by Robert A. Muenchen. International Statistical Review = Revue Internationale de Statistique, 77(3): 465–466, December 2009. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

                                        URL http://www.jstor.org/
                                        stable/27919771.

Walsh:1990:SSS

[Wal90] Anthony Walsh. Statistics for the social sciences: with computer applications. Harper and Row, New York, NY, USA, 1990. ISBN 0-06-046894-7. xvii + 373 pp. LCCN HA29 .W33574 1990.

Walker:2002:CSM

[Wal02] Glenn A. Walker. Common Statistical Methods for Clinical Research with SAS Examples. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 2002. ISBN 1-59047-040-0. xiv + 464 pp. LCCN W 20.5 W179t 2002.

Wang:2019:BRB

[Wan19] Xin Wang. Book review: Business Survival Analysis Using SAS: An Introduction to Lifetime Probabilities. The American Statistician, 73(2):208– 209, 2019. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http:

//amstat.tandfonline.com/
doi/full/10.1080/00031305.
2018.1538851.

Warner:2005:BRD

[War05] John Warner. Book review:

Data Mining Using SAS Applications by George Fernandes Chapman & Hall/CRC, with sep. CD-ROM, 2002. Computational Statistics, 20 (3):535–537, September 2005. CODEN CSTAEB. ISSN 0943-4062 (print), 1613-9658 (electronic). URL http://

link.springer.com/article/
10.1007/BF02741313.

Wellman:1985:USS

[WB85] Barry Wellman and Susan Gonzalez Baker. Using SAS software to link personal network, tie and individual data: a novice’s guide. Research paper; no. 157. Centre for Urban and Community Studies, University of Toronto, 1985. ISBN 0-7727-1320-0. 19 pp. LCCN HM131 .W42 1985.

Webel:2012:BRP

[Web12]  Karsten Webel.  Book re-

view: Panik, M. (2009): Regression Modeling — Methods, Theory, and Computation with SAS. Statistical Papers, 53(3): 803–804, August 2012. CO-DEN STPAE4. ISSN 0932-5026 (print), 1613-9798 (electronic).

                                        URL http://link.springer.
                                        com/accesspage/article/10.
                                        1007/s00362-011-0368-3.

Weinreich:1982:SIR

[Wei82] M. Weinreich. An SAS implementation of the Rasch item analysis with two categories. In Caussinus et al. [CEM82], pages 271–272. ISBN 3-7051-0002-5 (vol. 1), 3-7051-0001-7 (vol. 2). LCCN QA276.4 .C194 1982. Two volumes.

Weitkunat:1994:DSS

[Wei94] Rolf Weitkunat. Deskriptive Statistik in SAS: Einführung mit Beispielen aus der medizinischen Biometrie und Epidemiologie. Gustav Fischer Verlag, Stuttgart, Germany, 1994.

ISBN 3-437-50371-5. xii + 250 pp. LCCN ????

Wei:2012:CSP

[Wei12]  Xin Wei.  Code snippet:

%PROC R: a SAS macro that enables native R programming in the base SAS environment. Journal of Statistical Software, 46(CS-2):??, January 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v46/c02.

White:1988:ITP

[WHG88] Kenneth J. White, Shirley A.

Haun, and David J. Gow. Introduction to the theory and practice of econometrics — a computer handbook using SHAZAM and SAS: for use with Judge-Hill-Griffiths-Lutkepohl-Lee, “Introduction to the theory and practice of econometrics, second edition”. John Wiley and Sons, Inc., New York, NY, USA, 1988. ISBN 0-471-85946-X. 530 pp. LCCN HB 139 W523 1988. Includes index.

White:1994:CVN

[Whi94] A. P. White. Cross validation of nearest neighbour discriminant analysis — a warning to SAS users. Journal of Statistical Computation and Simulation, 49(3–4):129–140, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Wiegand:2006:SAM

[Wie06] Ryan E. Wiegand. Statistical analysis of medical data using SAS. The American Statistician, 60(4):345–346, November 2006. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wiegand:2008:DPA

[Wie08] Ryan E. Wiegand. Data preparation for analytics using SAS. The American Statistician, 62 (4):356–357, November 2008. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Willis:1989:RSC

[Wil89] Janet Marie Willis. Resources for statistical computation using SAS. Master of arts, University of California–Santa Cruz, 1989. viii + 191 + [52] pp. Typescript.

Wilson:2009:BRB

[Wil09] Brigid Wilson. Book review: A Handbook of Statistical Analyses Using SAS (Third Edition). Journal of Statistical Software, 30(BR-2):??, April 2009. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

                                        jstatsoft.org/v30/b02.

Wittkowski:1992:SAU

[Wit92] Knut M. Wittkowski. Statistical analysis of unbalanced and incomplete designs — experiences with BMDP and SAS. Computational Statistics & Data Analysis, 14(1):119–124, June 1992. CODEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
016794739290086U.

Wang:2019:EDM

[WJ19] Jianjun Wang and Dallas E.

Johnson. An examination of discrepancies in multiple imputation procedures between SAS^{T M} and SPSS^{T M}. The American Statistician, 73(1):80–88, 2019. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://amstat.
tandfonline.com/doi/full/
10.1080/00031305.2018.1437078.

Woodward:2005:BBC

[Woo05] Phil Woodward. BugsXLA:

Bayes for the Common Man. Journal of Statistical Software, 14(5):1–18, January 31, 2005. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/counter.php?
id=130&url=v14/i05&ct=2;
http://www.jstatsoft.org/
counter.php?id=130&url=v14/
i05/v14i05.pdf&ct=1.

Woods:2011:BRD

[Woo11] David Woods. Book reviews:

Design and analysis of experiments with SAS. John Lawson, Chapman and Hall/CRC, Boca Raton, 2010. No. of pages: 582. Price: $99.95. ISBN 10: 1-4200-6060-0, ISBN 13: 978-1-4200-6060-7. Statistics in Medicine, 30(10):1180– 1181, May 10, 2011. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wulff:2007:SMM

[Wul07] Shaun S. Wulff. SAS for mixed models. The American Statistician, 61(2):184–185, May 2007. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Wang:2009:CSC

[WWG09] Bing-Shun Wang, Xiao-Jin Wang, and Li-Kun Gong. Code snippet: The construction of a Williams design and randomization in cross-over clinical trials using SAS. Journal of Statistical Software, 29(CS-1):??, February 2009. CODEN JS-SOBK. ISSN 1548-7660. URL

                                        http://www.jstatsoft.org/
                                        v29/c01.

Wang:2024:NNL

[WWM^{+}24] Guoqiao Wang, Whedy Wang, Brian Mangal, Yijie Liao, Lon Schneider, Yan Li, Chengjie Xiong, Eric McDade, Richard Kennedy, Randall Bateman, and Gary Cutter. Novel nonlinear models for clinical trial analysis with longitudinal data: a tutorial using SAS for both frequentist and Bayesian methods. Statistics in Medicine, 43 (15):2987–3004, July 10, 2024. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:2010:SPE

[XB10] Xian-Jin Xie and Aihua Bian.

A SAS package for evaluating logistic and proportional odds model fit. Advances and Applications in Statistics, 14(1): 57–64, January 2010. CO-DEN ???? ISSN 0972-3617.

URL http://www.pphmj.com/
abstract/4520.htm.

Xie:2014:PHH

[XM14] Liang Xie and Laurence V.

Madden. %HPGLIMMIX: A highperformance SAS macro for GLMM estimation. Journal of Statistical Software, 58(8):??, June 2014. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v58/i08.

Yahav:2011:SAI

[Yah11] Eran Yahav, editor. Static Analysis: 18th International Symposium, SAS 2011, Venice, Italy, September 14–16, 2011. Proceedings, volume 6887 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2011. CO-DEN LNCSD9. ISBN 3-642-23701-0 (print), 3-642-23702-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-642-23702-7.

Yi:2006:SAI

[Yi06] Kwangkeun Yi, editor. Static Analysis: 13th International Symposium, SAS 2006, Seoul, Korea, August 29–31, 2006. Proceedings, volume 4134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2006. CO-DEN LNCSD9. ISBN 3-540-37756-5 (print), 3-540-37758-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.
springerlink.com/content/
978-3-540-37758-0.

York:1994:ARP

[Yor94] David York. Audit reports: a practical guide to preparation under SAS 600. Kogan Page, London, UK, 1994. ISBN 0-7494-0829-4. iv + 376 pp. LCCN ????

Yuan:2011:MIU

[Yua11] Yang Yuan. Multiple imputation using SAS software. Journal of Statistical Software, 45 (6):??, December 2011. CO-DEN JSSOBK. ISSN 1548-

7660.  URL http://www.
jstatsoft.org/v45/i06.

Zhang:2016:JSM

[ZCI^{+}16] Danjie Zhang, Ming-Hui Chen, Joseph G. Ibrahim, Mark E. Boye, and Wei Shen. JMFit: A SAS macro for joint models of longitudinal and survival data. Journal of Statistical Software, 71(??):??, ???? 2016. CODEN JSSOBK. ISSN 1548-7660. URL https://www.

jstatsoft.org/index.php/
jss/article/view/v071i03;

 https://www.jstatsoft.org/
index.php/jss/article/view/
v071i03/v71i03.pdf.

Zelterman:2002:ALL

[Zel02] Daniel Zelterman. Advanced Log–Linear Models Using SAS. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 2002. ISBN 1-59047-080-X. vii + 189 pp. LCCN QA278 .Z44 2002. URL http://

catdir.loc.gov/catdir/toc/
fy0701/2004304992.html.

Zhao:2019:BRS

[Zha19] Lili Zhao. Book review: SAS for Mixed Models: Introduction and Basic Applications, Walter W. Stroup, George A. Milliken, Elizabeth A. Claassen and Russell D. Wolfinger, Chapman and Hall/CRC, 2018, 446 pages, ebook, ISBN: 978-1-63526-152-3. International Statistical Review = Revue Internationale de Statistique, 87(2):449–450, August 2019. CODEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

Ziegel:1989:BRBr

[Zie89] Eric Ziegel. Book reviews: Annotated Computer Output for Analysis of Variance of Unbalanced Data: SAS GLM by S. R. Searle; R. P. Yerex; Annotated Computer Output for Analysis of Variance of Unbalanced Data: SPSSX ANOVA by S. R. Searle; R. P. Yerex. Technometrics, 31(3):397–398, August 1989. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/3556172.

Ziegel:1992:BRBu

[Zie92a] Eric R. Ziegel. Book review: SAS System for Linear Models by R. Littell; R. Freund; P. Spector. Technometrics, 34(4): 500, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1268965.

Ziegel:1992:BRBv

[Zie92b]  Eric R. Ziegel.  Book re-

view: SAS System for Regression by R. Freund; R. Littell. Technometrics, 34(4):500– 501, November 1992. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1268966.

Ziegel:1993:BRBj

[Zie93] Eric R. Ziegel. Book review: SAS System for Statistical Graphics by M. Friendly. Technometrics, 35(1):101, February 1993. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

                                        jstor.org/stable/1269319.

Ziegel:1995:BRBav

[Zie95a] Eric R. Ziegel. Book review: A Step-by-Step Approach to Using the SAS System for Univariate and Multivariate Statistics by L. Hatcher; E. Stepanski. Technometrics, 37(4):

471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1269762.

Ziegel:1995:BRBad

[Zie95b] Eric R. Ziegel. Book review:

Mastering the SAS System by Jay Jaffe. Technometrics, 37 (3):354–355, August 1995. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1269935.

Ziegel:1995:BRBau

[Zie95c] Eric R. Ziegel. Book review:

SAS Software Roadmaps by L. Burch; S. King. Technometrics, 37(4):470–471, November 1995. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1269761.

Ziegel:1996:BRBai

[Zie96a] Eric R. Ziegel. Book review:

Applied Multivariate for Statistics with SAS(R) Software by R. Khattree; D. Naik. Technometrics, 38(3):300, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270637.

Ziegel:1996:BRBav

[Zie96b]  Eric R. Ziegel.  Book re-

view: Categorical Data Analysis Using the SAS System by M. Stokes; C. Davis; G. Koch. Technometrics, 38(4):

413, November 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271334.

Ziegel:1996:BRBak

[Zie96c] Eric R. Ziegel. Book review: Introduction to Market Research Using the SAS(R) System by SAS Staff. Technometrics, 38 (3):300–301, August 1996. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1270639.

Ziegel:1996:BRBb

[Zie96d]  Eric R. Ziegel.  Book re-

view: Logistic Regression Examples Using the SAS System by SAS Institute Staff. Technometrics, 38(1):86–87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1268922.

Ziegel:1996:BRBc

[Zie96e] Eric R. Ziegel. Book review: Quick Results with SAS/Graph Software by A. Carpenter; C. Shipp. Technometrics, 38(1): 87, February 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1268923.

Ziegel:1996:BRBx

[Zie96f] Eric R. Ziegel. Book review: SAS Programming by Example by R. Cody; R. Pass. Technometrics, 38(2):194–195, May 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270440.

Ziegel:1996:BRBaj

[Zie96g] Eric R. Ziegel. Book review:

The How-To Book for SAS/ Graph Software by Thomas Miron. Technometrics, 38(3): 300, August 1996. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270638.

Ziegel:1997:BRBaf

[Zie97a]  Eric R. Ziegel.  Book re-

view: Forecasting Examples for Business and Economics Using the SAS(R) System by SAS Staff. Technometrics, 39(3): 345, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271162.

Ziegel:1997:BRBaq

[Zie97b] Eric R. Ziegel. Book review:

Handbook of Statistical Analysis Using SAS by B. Everitt; G. Der. Technometrics, 39(4): 434, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271530.

Ziegel:1997:BRBap

[Zie97c] Eric R. Ziegel. Book review:

Quick Start to Data Analysis with SAS by F. Di Iorio; K.

Hardy. Technometrics, 39(4): 434, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271529.

Ziegel:1997:BRBad

[Zie97d] Eric R. Ziegel. Book review: SAS(R) System for Mixed Models by R. Littell; G. Milliken; W. Stroup; R. Wolfinger. Technometrics, 39(3):344, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271160.

Ziegel:1997:BRBac

[Zie97e] Eric R. Ziegel. Book review: Statistical Quality Control Using the SAS(R) System by Dennis King. Technometrics, 39(3): 343–344, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271159.

Ziegel:1997:BRBar

[Zie97f]  Eric R. Ziegel.  Book re-

view: Stock Market Analysis Using the SAS System: Technical Analysis by SAS Staff. Technometrics, 39(4):434–435, November 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271531.

Ziegel:1997:BRBae

[Zie97g]  Eric R. Ziegel.  Book re-

view: Survival Analysis Using the SAS System by Paul D. Allison. Technometrics, 39(3):344– 345, August 1997. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271161.

Ziegel:1998:BRBae

[Zie98a]  Eric R. Ziegel.  Book re-

view: Applied Statistics and the SAS Programming Language by Ronald P. Cody; Jeffrey K. Smith. Technometrics, 40(3): 266–267, August 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271202.

Ziegel:1998:BRBr

[Zie98b]  Eric R. Ziegel.  Book re-

view: Extending SAS(R) Survival Analysis Techniques for Medical Research by Alan Cantor. Technometrics, 40(2): 164, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270668.

Ziegel:1999:BRBs

[Zie99a]  Eric R. Ziegel.  Book re-

view: Common Statistical Methods for Clinical Research with SAS Examples by Glenn A. Walker. Technometrics, 41 (2):179–180, May 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1270753.

Ziegel:1999:BRBae

[Zie99b]  Eric R. Ziegel.  Book re-

view: Programming Techniques for Object-Based Statistical Analysis with SAS Software by Tanya Kolosova; Samuel Berestizhevsky. Technometrics, 41(3):272, August 1999. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

                                        tronic). URL http://www.
                                        jstor.org/stable/1270590.

Ziegel:1999:BRBj

[Zie99c] Eric R. Ziegel. Book review: SAS(R) System for Elementary Statistical Analysis by Sandra D. Schlotzhauer; Ramon C. Littell. Technometrics, 41(1): 84, February 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1271019.

Ziegel:1999:BRBaf

[Zie99d]  Eric R. Ziegel.  Book re-

view: Strategic Data Warehousing Principles Using SAS Software by Peter R. Welbrock. Technometrics, 41(3): 272, August 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1270591.

Ziegel:1999:BRBt

[Zie99e]  Eric R. Ziegel.  Book re-

view: Survey Research Using the SAS System by Archer Gravely. Technometrics, 41 (2):180, May 1999. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1270754.

Ziegel:2000:BRBx

[Zie00a] Eric R. Ziegel. Book review:

Applied Multivariate Statistics with SAS(R) Software by Ravindra Khattree; Dayanand N. Naik. Technometrics, 42 (3):324, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271113.

Ziegel:2000:BRBw

[Zie00b] Eric R. Ziegel. Book review:

Logistic Regression Using the SAS(R) System by Paul D. Allison. Technometrics, 42(3):323– 324, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271112.

Ziegel:2000:BRBj

[Zie00c]  Eric R. Ziegel.  Book re-

view: SAS(R) Macro Programming Made Easy by Michele M. Burlew. Technometrics, 42(2): 221–222, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271485.

Ziegel:2001:BRBab

[Zie01a]  Eric R. Ziegel.  Book re-

view: Categorical Data Analysis Using the SAS(R) System by Maura E. Stokes; Charles S. Davis; Gary G. Koch. Technometrics, 43(2):249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1271072.

Ziegel:2001:BRBaa

[Zie01b]  Eric R. Ziegel.  Book re-

view: Multivariate Data Reduction and Discrimination with SAS(R) Software by Ravindra Khattree; Dayanand N. Naik. Technometrics, 43(2): 248–249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/
stable/1271071.

Ziegel:2001:BRBau

[Zie01c] Eric R. Ziegel. Book review: SAS(R) System for Regression by Rudolf J. Freund; Ramon C. Littell. Technometrics, 43 (4):498, November 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.
jstor.org/stable/1270841.

Ziegel:2002:BRBah

[Zie02] Eric R. Ziegel. Book review: A Handbook of Statistical Analyses Using SAS by Geoff Der; Brian S. Everitt. Technometrics, 44(3):300–301, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270518.

Zimmerman:2007:SPD

[Zim07] M. Bridget Zimmerman. SAS programming and data analysis: a theory and programdriven approach. The American Statistician, 61(3):277–278, August 2007. CODEN AS-TAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

Zink:2012:NSI

[ZK12] Richard C. Zink and Gary G.

Koch. NParCov3: a SAS/ IML macro for nonparametric randomization-based analysis of covariance. Journal of Statistical Software, 50(3):??, July 2012. CODEN JSSOBK. ISSN 1548-7660. URL http://

www.jstatsoft.org/v50/i03.

Zamar:2007:ESI

[ZMG07] David Zamar, Brad McNeney, and Jinko Graham. elrm: Software implementing exact-like inference for logistic regression models. Journal of Statistical Software, 21(3):1–18, October 2007. CODEN JSSOBK. ISSN 1548-7660. URL http://www.

jstatsoft.org/v21/i03.